### v1

# Mathematical Language Models: A Survey

Abstract
In recent years, there has been remarkable progress in leveraging Language Models (LMs), encompassing Pre-trained Language Models (PLMs) and Large-scale Language Models (LLMs), within the domain of mathematics. This paper conducts a comprehensive survey of mathematical LMs, systematically categorizing pivotal research endeavors from two distinct perspectives: tasks and methodologies. The landscape reveals a large number of proposed mathematical LLMs, which are further delineated into instruction learning, tool-based methods, fundamental CoT techniques, and advanced CoT methodologies. In addition, our survey entails the compilation of over 60 mathematical datasets, including training datasets, benchmark datasets, and augmented datasets. Addressing the primary challenges and delineating future trajectories within the field of mathematical LMs, this survey is positioned as a valuable resource, poised to facilitate and inspire future innovation among researchers invested in advancing this domain.
keywords:
Mathematics, Language Models, Pre-trained, LLMs, Survey
[…]
1 Introduction
[…]
Mathematics is the queen of sciences.
– Carl Friedrich Gauss
Mathematics stands as a foundational skill integral to human intelligence, wielding significance across diverse fields such as natural sciences, engineering, medicine, finance, computer science, and social sciences. Within the field of natural language processing (NLP), the development of computer models aimed at autonomously resolving mathematical word problems has captivated researchers since as early as 1963 [feigenbaum1963computers, bobrow1964natural, briars1984integrated, fletcher1985understanding]. We believe that addressing this problem is a potential way toward illuminating pathways for general reasoning mechanisms, consequently advancing the pursuit of general artificial intelligence (AI).
One main traditional solution to math word problems is statistical learning-based methodologies [hosseini-etal-2014-learning, kushman-etal-2014-learning, zhou2015learn, mitra2016learning]. Notably, machine learning techniques [kushman-etal-2014-learning, roy2015reasoning, roy2015solving] alongside semantic parsing methods [shi2015automatically, koncel2015parsing] have been employed to address this challenge, showcasing promising outcomes on certain datasets. The evolution of deep learning has prompted considerable interest in crafting neural networks capable of resolving mathematical problems [wang2017deep, couperus2023large]. Notably, recent years have witnessed significant advancements in the domain of mathematical AI [wei2022chain, matzakos2023learning], propelled by the emergence of powerful language models (LMs) [kenton2019bert, touvron2023llama]. These language models, comprising pre-trained language models (PLMs) and large-scale language models (LLMs), have assumed a central role in reshaping the landscape of mathematical exploration and practical applications. The paramount focus of this comprehensive survey lies in assessing the impact of these models on the field of mathematics. The survey endeavors to provide a thorough overview of extant research and its consequential implications.
PLMs such as BERT [devlinBERTPretrainingDeep2019], RoBERTa [liuRoBERTaRobustlyOptimized2019], BART [lewis2020bart], GPT-1 [radford_gpt-1_2018] and GPT-2 [radford_gpt-2_2019] undergo pre-training on extensive textual corpora to assimilate worldly knowledge. To enhance mathematical performance, certain endeavors focus on either pre-training or fine-tuning PLMs using mathematical datasets [clarkRegentsScienceExams2021, gevaInjectingNumericalReasoning2020, fengInjectingNumericalReasoning2022]. For instance, methodologies like GenBERT [gevaInjectingNumericalReasoning2020], NF-NSM [fengInjectingNumericalReasoning2022], MathBERT[pengMathBERTPretrainedModel2021] and LISA [jiang2021lisa] incorporate numerical data or mathematical formulas into PLMs to augment their capabilities. Moreover, specialized modules or tailored loss functions are devised, leveraging existing PLMs to learn mathematical reasoning and operations [zhu2021tat, zhao2022multihiertt, jieLearningReasonDeductively2022, liSeekingPatternsNot2022].
The recent advent of LLMs, exemplified by OpenAI’s GPT-4 [openai2023gpt4], has catalyzed an unforeseen surge in innovation, underscoring the multifaceted potential of AI within the domain of mathematics. These models ([touvron2023llama, touvron2023llama2]) have demonstrated remarkable success across diverse Natural Language Processing (NLP) tasks, leveraging in-context learning [min2022metaicl, brown2020language, chen2022meta] and instruction learning [yang2023gpt, liu_goat_2023]. Recent studies by Wang et al. [wang2022self] indicate that LLMs featuring over 100 billion parameters (e.g., GPT-3 [brown2020language] with 175 billion, PaLM [chowdhery2022palm] with 540 billion) exhibit the capability to address intricate tasks by employing a chain-of-thought (CoT) mechanism [wei2022chain] when furnished with a limited set of reasoning examples as demonstrations. Advancements in CoT frameworks [chen2022program, lu2022dynamic, zhang2023automatic, fu2022complexity] have been tailored to enhance mathematical performance, incorporating tools and programs [gao2023pal, drori2022neural, he2023solving]. An exemplar in this domain is LLEMMA [azerbayev8llemma], an openly accessible LLM specifically designed for mathematical tasks. LLEMMA showcases the capacity to generate self-contained textual solutions for mathematical problems and formulate formal proofs.
We mainly review the most related surveys about mathematical reasoning and LLMs. Lu et al. [lu-etal-2023-survey] delineate the landscape of deep learning applications specifically pertaining to mathematical reasoning. Qiao et al. [qiao-etal-2023-reasoning] predominantly delve into reasoning mechanisms centered around language models, encompassing arithmetic, commonsense, logical, symbolic, and multi-modal reasoning. Chu et al. [chuCoTReasoningSurvey2023] offer an exploration of studies focused on chain-of-thought (CoT) reasoning. Additionally, Qiu el al. [qiu2020pre] and Naveed et al. [naveed2023comprehensive] contribute comprehensive reviews elucidating the realm of existing PLMs and LLMs, respectively.
Diverging from these perspectives, our work delves into the burgeoning domain of mathematical language models, systematically summarizing the diverse range of studies and innovations in this field. In light of the demonstrated proficiency of LMs in comprehending, generating, and manipulating mathematical expressions, a multitude of advanced methodologies have emerged within the field of mathematics leveraging LMs. By shedding light on the current landscape, tasks, datasets, challenges, and future prospects, we aim to inspire and inform those who aspire to harness the power of language models to revolutionize mathematics and its myriad applications.
Specifically, our categorization of mathematical tasks (§2) contain two primary domains: mathematical calculation (§2.1), consisting of arithmetic representation and arithmetic calculation, and mathematical reasoning (§2.2), consisting of problem-solving and theorem proving. Furthermore, our taxonomy of existing algorithms segregates them into PLMs-based approaches (§3), including autoregression (§3.1) and non-autoregression (§3.2) LMs, and LLMs-based methodologies (§4), which include instruction learning (§4.1), tool-based strategies (§4.2), fundamental CoT techniques (§4.3), and advanced CoT methodologies (§4.4). Moreover, we provide a comprehensive compilation of over 60 mathematical datasets (§5), systematically categorized into training (§5.1), benchmark (§5.2), and augmented datasets (§5.3). This classification aims to facilitate research by delineating the utility and applicability of these datasets within distinct research contexts. Lastly, our survey meticulously addresses the primary challenges and explores the future directions of this domain (§6), including faithfulness, multi-modal, uncertainty, evaluation, creation and application. In conclusion (§7), we believe that this comprehensive survey will serve as a valuable asset for researchers and practitioners alike, seeking to push the boundaries of innovation in the field of mathematical language models.
2 Mathematical Tasks
In this section, we summarize the existing mathematical tasks into mathematical calculation and mathematical reasoning (Figure 1).
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=6pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.6em,font=,, where level=2text width=9.4em,font=,, where level=3text width=9.5em,font=,, where level=4text width=12em,font=,, [ Mathematical Tasks, ver [ Mathematical
Calculation  (§2.1) [ Arithmetic
Representation [ DigitRNN  [spithourakis2018numeracy], DigitCNN  [wallace2019nlp], GenBERT  [gevaInjectingNumericalReasoning2020], NumBERT  [zhangLanguageEmbeddingsCapture2020] , leaf, text width=44.6em ] ] [ Arithmetic
Calculation [ ScratchpadGPT  [nye_show_2021], Goat  [liu_goat_2023], MathGLM  [yang2023gpt], , leaf, text width=44.6em ] ] ] [ Mathematical
Reasoning  (§2.2) [ Math Word Problem
Solving [ MathPrompter  [imani2023mathprompter], MetaMath  [yu2023metamath], WizardMath  [luo2023wizardmath], MathAttack  [zhou2023mathattack], LLEMMA  [azerbayev8llemma] , leaf, text width=44.6em ] ] [ Math Question
Answering [ GEOS  [seo-etal-2015-solving], DROP  [dua2019drop], Mathematics  [saxton2018analysing], Lila  [mishra2022lila] , leaf, text width=44.6em ] ] [ Theorem
Proving [ DeepMath  [irving2016deepmath], ASTactic  [yang2019learning], NaturalProofs  [welleck2021naturalproofs], INT  [wu2020int] , leaf, text width=44.6em ] ] ] ]
Figure 1: Taxonomy of mathematical tasks.
2.1 Mathematical Calculation
The advent of Language Models (LMs) has ushered in a new era of exploration into their computational capabilities, particularly in arithmetic. Initially, these models demonstrated basic computational abilities by representing numbers in textual formats. With the increasing capabilities of LMs, it has been observed that they can acquire arithmetic skills through methods such as fine-tuning, even without carefully crafted numeric representations.
2.1.1 Arithmetic Representation
In the early stages of the research, numerical values were either omitted or oversimplified, treated as ordinary text, or categorized as “unknown" (UNK). This approach, however, proved inadequate for tasks. For instance, BERT performs five times worse when the answer is a number rather than a textual span over the DROP question-answering benchmark [dua2019drop].
Recent literature proposes various methods for numerical representation. Geva et al. [gevaInjectingNumericalReasoning2020] introduce GenBERT, which tokenizes numbers at the digit level and undergoes fine-tuning with arithmetic word problems and simple arithmetic tasks. Zhang et al. [zhang2020language] experiment with a retrained BERT model by converting numbers into scientific notation (e.g., 314.1 as 3141[EXP]2). Spithourakis and Riedel [spithourakis2018numeracy], along with Wallace et al. [wallace2019nlp], explore the integration of digit embeddings into a singular embedding that represents the entire number. Berg-Kirkpatrick and Spokoyny [berg-kirkpatrickEmpiricalInvestigationContextualized2020] propose a method with digit-RNN and exponent embeddings. This method specifically emphasizes the exponent while disregarding the mantissa. Goat [liu_goat_2023] introduces Consistent Tokenization, enhancing the relationships between similar numerical values.
2.1.2 Arithmetic Calculation
There has been significant research into the arithmetic capabilities of LLMs. Nogueira et al. [nogueira2021investigating] and Wang et al. [wang2021exploring] assessed addition and subtraction tasks. Muffo et al. [muffo_evaluating_nodate] evaluated two-digit multiplication. Yuan et al. [yuan2023well] evaluated the arithmetic operation capabilities of different models, including GPT-4 [openai2023gpt4], Galactica [taylor_galactica_2022], and LLaMA [touvron2023llama].
Traditionally, it was presumed that LLMs could not accurately perform complex arithmetic, especially multiplication involving more than eight digits. However, recent approaches challenge this assumption. Zhou et al. [zhouTeachingAlgorithmicReasoning2022] apply specialized prompt engineering to enhance addition capabilities but note limitations in multiplication beyond seven digits. Jelassi et al. [jelassiLengthGeneralizationArithmetic2023] investigate length generalization in basic arithmetic tasks using techniques like relative position embeddings and training set priming. ScratchpadGPT [nye_show_2021] demonstrates the effectiveness of pre-generating a Chain of Thought (CoT) before producing an answer in 8-digit addition tasks. Goat [liu_goat_2023] utilizes supervised instruction for fine-tuning elementary arithmetic operations with large integers. MathGLM [yang2023gpt] excels in intricate arithmetic tasks by pre-training on datasets that decompose complex arithmetic expressions into simpler steps. Through this process, it gradually generates answers and learns the rules of computation.
2.2 Mathematical Reasoning
Mathematical reasoning is a pivotal aspect of artificial intelligence that facilitates the understanding and solving of complex mathematical problems. The integration of LLMs in this domain has been significant, owing to their ability to interpret, process, and generate complex natural text. This section delves into the current research and developments within this field, focusing on two main areas: math problem solving and theorem proving.
2.2.1 Math Problem Solving
In the field of math problem solving in artificial intelligence, math problem solving refers to the process of using algorithms, computational models, and increasingly LLMs to understand, explain, and solve mathematical problems. This approach to solving problems spans all levels from basic arithmetic to advanced mathematics, including but not limited to algebra, geometry, statistics, and calculus. In this subsection, we condense it into Math Word Problems (MWPs) and Math Question Answering (MQA).
Math Word Problem Solving.
Math Word Problems (MWPs) refer to problems that present mathematical concepts and calculations in the form of written descriptions. Such questions usually include one or more story situations describing a series of events or conditions from which the solver needs to extract relevant mathematical information and apply appropriate mathematical principles to solve the problem. Over the years, using various efficient and intelligent algorithms to solve MWPs has also been a research focus in the fields of computer science and artificial intelligence [bobrow1964natural].
Recent research has highlighted the growing capabilities of LLMs in the field of mathematical word problem solving, emphasizing the trend toward more nuanced and sophisticated AI-driven mathematical analysis. MathPrompter [imani2023mathprompter] uses GPT-3 Davinci [brown2020language] to solve MWPs with excellent results, demonstrating the potential of LLMs to not only explain but also generate complex mathematical reasoning, reflecting a human-like understanding of complex problem sets.
Yuan et al. [yuan2023scaling] investigated the interaction between various factors such as pre-training loss, amount of supervised data, and augmented data to improve the mathematical inference performance of LLMs. They found that pre-training loss is a more reliable performance metric and empirically found a log-linear relation between data amount and model performance, but it is worth mentioning that this improvement diminishes with better pre-training models. At the same time, they proposed Rejection sampling Fine-Tuning (RFT) to enhance performance by using supervised models to generate and collect correct reasoning paths, thereby achieving better inference generalization and more significant improvements. Meanwhile, the proficiency with which LLMs perform basic arithmetic tasks is critical to establishing the baseline capabilities of these models [yuan2023well]. This assessment is integral to understanding the limitations and potential of LLMs in future work when faced with mathematical challenges that require precise and logical calculations.
MetaMath [yu2023metamath] further extends the practicality of the LLMs, proposing a paradigm whereby the LLMs generate their mathematical problems, thereby creating a self-sustaining learning environment that encourages continuous improvement of the model’s problem-solving acuity. Similarly, WizardMath [luo2023wizardmath] explores new ways to enhance LLMs’ mathematical reasoning capabilities by reinforcing evolutionary instructions, indicating an important step towards autonomous self-improvement of artificial intelligence models. Instead, MathAttack [zhou2023mathattack] introduces a critical perspective by exploring the sensitivity of LLMs to specialized adversarial input designed to test their mathematical problem-solving abilities. Such investigations are crucial to developing more resilient and reliable LLMs that can withstand and adapt to a variety of problem-solving situations. LLEMMA [azerbayev8llemma], an open language model dedicated to mathematics, outlines ongoing efforts to make LLMs accessible and adaptable to a wide range of mathematical inquiries, from basic problem solving to advanced theorem proving.
The present findings confirm that these studies comprehensively depict the current research hotspots and future development potential directions of LLMs in the MWPs field. They emphasize that LLMs can have a transformative impact on the field of mathematics, providing innovative tools for educators, students, and researchers. As these models continue to develop, they are expected to revolutionize the way we approach and solve mathematical problems, ushering in a new era of mathematical reasoning and assisted education through artificial intelligence.
Math Question Answering.
Math Question Answering (MQA) refers to the computational task where a system is required to understand and automatically solve math-related questions presented in natural language. These questions can range from simple arithmetic problems to complex high school or college-level mathematics, including algebra, calculus, and geometry. The challenge for MQA systems is to accurately interpret the text, convert it into an appropriate mathematical representation, perform the necessary computations, and generate a correct and often step-by-step solution, mimicking the problem-solving process of a human.
Sachan et al. [saxton2018analysing] introduce a framework for evaluating neural architectures through a suite of math problems, highlighting the importance of algebraic generalization. This is complemented by GEOS [seo-etal-2015-solving], an innovative system that solves SAT geometry problems by integrating text understanding with diagram interpretation. The introduction of datasets and systems like Inter-GPS [lu2021inter], IconQA [lu2021iconqa], and PGDP5K [Hao2022PGDP5KAD, Zhang2022, Zhang2023PGPS] further advances the field, offering new benchmarks for abstract diagram understanding and visual language reasoning. Additionally, Scharpf et al. [scharpf2022mining] indicate that unsupervised methods for formula labeling contribute to the automated understanding of mathematical documents. Collectively, these works demonstrate the evolving landscape of MQA, where the synergy between natural language processing, computer vision, and symbolic reasoning opens new avenues for educational technology and AI research.
2.2.2 Theorem Proving
Theorem proving (TP) refers to the process of demonstrating that a statement is correct based on existing axioms and facts, which is a long-term challenge in AI [10.1145/1455567.1455605].In the context of LLMs, people began to try to use the huge knowledge base of LLMs and some manual guidance to solve this task. LLMs like GPT-3 [brown2020language] can process natural text to understand the premises of the theorem, and apply certain logical reasoning rules to complete the proof requirements. This can then be done by calling relevant tools or using manual inspection to check the correctness of the process. The advanced approach aims to leverage LLM’s superior computing power and vast knowledge base relative to humans to automate the theorem-proving process, a task that previously required professional mathematicians.
In the field of AI, research on theorem proving has progressed from data sets containing large numbers of human-written proofs (such as CoqGym [yang2019learning]) to complex models that can autonomously generate proof strategies (such as ASTactic [yang2019learning]). This improvement is reflected in the application of some language models that are based on Transformer [vaswani2017attention] to TP tasks, especially in recent research such as GPT-f [polu2020generative]. Most notably, some of the proofs generated by these models have been formally recognized by the mathematical community. Jiang et al. [jiang2022draft] further refined this process by using informal proofs to guide TP models. NaturalProofs [welleck2021naturalproofs] extends these capabilities by leveraging the language of natural mathematics to create a rich corpus of model training and evaluation. At the same time, DeepMath [irving2016deepmath] and INT [wu2020int] have promoted the development of this field by demonstrating the effectiveness of neural sequence models for premise selection and evaluating the LLMs’ generalization ability in theorem proving.
Ongoing research is delving into the integration of language models and interactive proof assistants within the realm of mathematics. This involves generating proofs through tactic prediction [DBLP:journals/corr/abs-2009-03393, han2021proof, lample2022hypertree, jiang2022thor], automating formalization processes [DBLP:conf/nips/WuJLRSJS22, jiang2022draft], and developing unified tools [llmstep]. Due to the substantial computational demands of exploration, language models applied in this domain have traditionally been limited in size. However, recent advancements have showcased potential in employing larger models [first2023baldur, jiang2022draft]. The work of Azerbayev et al. (LLEMMA) [azerbayev8llemma] demonstrates few-shot proof autoformalization and tactic prediction, offering a vast dataset of formal mathematics and an openly accessible model for further investigation in these directions. These endeavors establish crucial theoretical and practical frameworks for automating and enhancing tactic prediction.
The above series of progressive research work shows that utilizing the powerful computing power and extensive knowledge base of LLMs is the main development direction for future TP tasks. The more excellent models (such as GPT-4 [openai2023gpt4]) obviously have better reasoning capabilities without any human intervention, but we also need to be clear about one thing, that is the inevitable hallucination problem in neural text generation models [maynez-etal-2020-faithfulness]. The hallucination problem refers to the generated content that is nonsensical or unfaithful to the provided source content [ji2023survey]. The occurrence of hallucinations at any step in the reasoning process will lead to errors in the final reasoning results. Therefore, solving the hallucination problem is also particularly important in future work related to TP.
3 PLMs-based Methods
Based on the transformer architecture [vaswani2017attention] with self-attention mechanisms, the “pre-training and fine-tuning" paradigm has revolutionized the field of natural language processing. Within the context of Pre-trained Language Models (PLMs), numerous approaches have been proposed to tackle text generation problems [noorbakhsh2021pretrained]. Autoregressive LMs (ALMs, e.g., GPT-1 [radford_gpt-1_2018] and T5 [rael_exploring_nodate]) and Non-Autoregressive LMs (NALMs, e.g., BERT [kenton2019bert] and Roberta [liuRoBERTaRobustlyOptimized2019]) are the two primary strategies for mathematics (Figure 2).
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.4em,font=,, where level=2text width=8.4em,font=,, where level=3text width=8.0em,font=,, where level=4text width=12em,font=,, [ Language Models, ver [ Pre-trained
Language
Models  (§3) [ Autoregression
LMs  (§3.1) [ GPT-$f$ [polu2020generative], EPT [kim_point_2020] , Generate &\&& Rank [shen_generate_2021] , Thor [jiang2022thor] , HTPS [lample_hypertree_nodate] , Galactica [taylor_galactica_2022] , MATH-PLM [hendrycks2021measuring] , LISA [jiang2021lisa]
PACT [han2021proof] , Minerva [lewkowycz_solving_nodate] , LIME [wu_lime_nodate] , leaf, text width=54.2em ] ] [ Non-Autoregression
LMs  (§3.2) [ Aristo [clarkRegentsScienceExams2021], GenBERT [gevaInjectingNumericalReasoning2020] , NF-NSM [fengInjectingNumericalReasoning2022] , MWP-BERT [liangMWPBERTNumeracyaugmentedPretraining2022] , MathBERT[pengMathBERTPretrainedModel2021] , TAGOP [zhu2021tat] , MT2Net [zhao2022multihiertt]
DeductReasoner [jieLearningReasonDeductively2022] , BERT-TD [liSeekingPatternsNot2022] , leaf, text width=54.2em ] ] ] [ Large
Language
Models  (§4) [ Instruction
Learning  (§4.1) [ Instruction
Building [ Auto-explanation [nakamoto2023enhancing], ProofNet [azerbayev2023proofnet] , WizardLM [xu_wizardlm_2023] , Wizardmath [luo2023wizardmath] , leaf, text width=44.6em ] ] [ Instruct
Tuning [ MathGLM [yang2023gpt], Goat [liu_goat_2023], Calculon [muffo_evaluating_nodate], PaLM 2-L-Math [liu_improving_2023], LLEMMA [azerbayev8llemma] , leaf, text width=44.6em ] ] [ In-context
Learning [ ScratchpadGPT [nye_show_2021] , Codex-math [drori2022neural] , In-Context Sampling [liu2022makes] , ComplexPrompt [fu2022complexity]
ART [zhang2023automatic] , leaf, text width=44.6em ] ] ] [ Tool-based
Methods  (§4.2) [ Single-tool
Methods [ SymbLLM [he2023solving], PoT [chen2022program], Codex-math [drori2022neural] , MathPrompter [imani2023mathprompter] , PAL [gao2023pal] , Code-scratchpad [upadhyay-etal-2023-improving] , leaf, text width=44.6em ] ] [ Multi-tool
Methods [ Toolformer [schick2023toolformer], Chameleon [lu2023chameleon] , ART [paranjape2023art] , ToolkenGPT [hao2023toolkengpt] , CRITIC [gou2023critic] , TALM [parisi2022talm]
Documentation [hsieh2023tool] , leaf, text width=44.6em ] ] ] [ Fundamental CoT
Methods  (§4.3) [ Foundation of
CoT [ Answer Rationales [ling2017program] , MathQA [amini2019mathqa], CoT [wei2022chain], Complexity-based CoT [fu2022complexity] , LVSS [lightman2023lets], RFT [yuan2023scaling]
Thought Propagation [yu2023thought] , AFT [wang2023making] , PoT [chen2022program] , AOT [sel2023algorithm] , MAmmoTH [yue2023mammoth] , leaf, text width=44.6em ] ] [ Construction of
CoT [ Zero-shot CoT [kojima_large_nodate] , Auto-CoT [zhang2023automatic] , Complexity-based CoT [fu2022complexity] , PromptPG-CoT [lu2022dynamic]
AutoMate CoT [shum_automatic_2023] , BoostedPrompt [pitis_boosted_2023] , leaf, text width=44.6em ] ] ] [ Advanced CoT
Methods  (§4.4) [ Verify-based
Methods [ Code-based self-verification [zhou2023solving] , VerifyCoT [ling2023deductive] , DIVERSE [li2023making] , Verify-and-Edit [zhao2023verify]
Retrieval-CoT [he_rethinking_2022] , SCREWS [shridhar2023screws] , leaf, text width=44.6em ] ] [ Ensemble-based
Methods [ Self-Consistency [wang2022self] , Diversity-of-Thought [naik2023diversity] , Complexity-based CoT [fu2022complexity], DIVERSE [li2023making]
Self-check [miao2023selfcheck], MCR [yoran2023answering], Rank-verifier [cobbe2021training], GRACE [khalifa2023discriminator] , leaf, text width=44.6em ] ] [ Planing-based
Methods [ ToT [yao2023tree] , TouT [mo2023tree] , ToT-system [long2023large] , GoT [yao2023beyond, besta_graph_2023, lei2023boosting] , RESPROMPT [jiang2023resprompt]
RAP [hao2023reasoning] , LATS [zhou2023language] , LLM+P [liu2023llm+] , LLM+DP [dagan2023dynamic] , Self-refine [madaan2023self, sun2023adaplanner] , ISR-LLM [zhou2023isr]
Reflexion [shinn2023reflexion] , leaf, text width=44.6em ] ] [ Socratic Teaching
Methods [ ScoREval [he2023socreval], Socratic Models [zeng2022socratic], SOCRATIC QUESTIONING [qi2023art] , leaf, text width=44.6em ] ] ] ] ]
Figure 2: Taxonomy of language models for mathematic.
Figure 3: The distinctions between Autoregression LMs and Non-Autoregression LMs.
3.1 Autoregression LMs
The distinctions between Autoregression LMs(ALMs) and Non-Autoregressive LMs(NALMs) in sequence generation tasks are depicted in Figure 3. Furthermore, ALMs can be categorized into two architectures: causal decoder (e.g., GPT-1 [radford_gpt-1_2018] and GPT-2 [radford_gpt-2_2019]) and encoder-decoder (e.g., T5 [rael_exploring_nodate]). The generation mechanisms of these two structures for sequences are depicted on the left and in the middle of Figure 3, respectively. These two architectures have not only significantly propelled the advancement of PLMs but have also become the predominant foundational architectures for LLMs.
To enhance the performance of PLMs in mathematical problems, many improved methods based on ALMs have been proposed. GPT-$f$ [polu2020generative] presents an automated prover and proof assistant to explore the performance of LMs to automated theorem proving. To resolve challenges related to expression fragmentation and operand-context separation, Kim et al. [kim_point_2020] introduce the Expression-Pointer Transformer (EPT), a pure neural model that employs both an ‘Expression’ token and operand-context pointers during the generation of solution equations. Generate &\&& Rank [shen_generate_2021] introduces a novel ranking task for MWP, a multi-task framework built upon a generative PLM, effectively addressing the challenge of minor mistakes in mathematical expressions. Thor [jiang2022thor] introduces a framework that integrates language models and automated TP. The latter is employed to selectively choose relevant premises from an extensive library to assist the language model in TP. HTPS [lample_hypertree_nodate] presents HyperTree Proof Search, a novel algorithm leveraging online training from past proof searches, enhancing model performance and surpassing the previous SOTA GPT-$f$. Galactica [taylor_galactica_2022] proposes a working memory token approach that can achieve strong performance over existing methods on mathematical MMLU [hendryckstest2021] and MATH [hendrycks2021measuring] benchmarks.
Furthermore, MATH-PLM [hendrycks2021measuring] introduces a new dataset named MATH, consisting of challenging competition mathematics problems. It is found that relying solely on increasing budgets and model parameter counts would be impractical for achieving robust mathematical reasoning. LISA [jiang2021lisa] gathers an extensive collection of lemmas and theorems extracted from the Isabelle standard library and the Archive of Formal Proofs (AFP). Using this vast corpus, it builds LMs that prove theorems effectively within the AFP. PACT [han2021proof] extracts abundant self-supervised data from kernel-level proof terms, enabling joint training alongside the standard tactic prediction objective. This effectively enhances the model’s ability for TP. Minerva [lewkowycz_solving_nodate] trains a language model, initially pre-trained on general natural language data, through additional training on technical content to address Quantitative Reasoning Problems. It achieves a SOTA performance on technical benchmarks including MATH [hendrycks2021measuring]. LIME [wu_lime_nodate] introduces a novel pre-training methodology that specifically learns inductive bias for mathematical reasoning. Observations indicate that models trained using LIME exhibit superior performance compared to standard transformers.
3.2 Non-Autoregression LMs
Unlike ALMs, NALMs enable the model to generate all parts of a sequence simultaneously, without depending on previously generated content, as depicted on the right side of Figure 3. Notably, models such as BERT [kenton2019bert], Roberta [liuRoBERTaRobustlyOptimized2019], and related architectures utilize masked word representations, leveraging pre-trained context-aware embeddings as comprehensive semantic features. This significantly elevates the performance benchmarks of NLP tasks.
In the field of mathematics, researchers have proposed various methods for designing models to address mathematical reasoning and computational problems. For example, Aristo [clarkRegentsScienceExams2021] fine-tunes BERT using scientific curriculum data, yielding promising results on science exams. GenBERT [gevaInjectingNumericalReasoning2020] and NF-NSM [fengInjectingNumericalReasoning2022] enhance the numerical reasoning capabilities of models by incorporating numerical data into the training process of PLMs. MWP-BERT [liangMWPBERTNumeracyaugmentedPretraining2022] further enhances the model’s capacity to represent and calculate numerical values by incorporating numeric attributes into symbol placeholders. MathBERT[pengMathBERTPretrainedModel2021] employs additional joint training of text and formulas to effectively capture the semantic-level structural information of formulas. TAGOP [zhu2021tat], MT2Net [zhao2022multihiertt], and DeductReasoner [jieLearningReasonDeductively2022] utilize BERT or RoBERTa to extract the fundamental arithmetic relationships between quantities, enabling mathematical reasoning and operations. BERT-TD [liSeekingPatternsNot2022] utilizes semantic encoding and contrastive learning to cluster problems with similar prototype equations, thereby enhancing the understanding of MWP patterns.
4 LLMs-based Methods
Large-scale language models (LLMs) are designed for processing and generating text akin to human communication [touvron2023llama, openai2023gpt4]. Mathematics is also a form of language, that communicates complex concepts and relationships through a structured system of symbols and notation, akin to the rules of a spoken language. Thus, a language model that grasps these mathematical rules can “speak" the language of mathematics, proving to be a valuable asset for mathematicians. In numerous ways, a sophisticated language model like GPT-4 [openai2023gpt4] becomes an invaluable tool in the field of mathematics [azerbayev8llemma, luo2023wizardmath]. We classify the existing studies into four parts: instruction learning, tool-based methods, fundamental CoT methods, and advanced CoT methods (Figure 2).
4.1 Instruction Learning
Numerous approaches have been proposed to enhance the mathematical performance of models by instruction learning. These approaches are categorically delineated as instruction building, instruction tuning, and in-context learning, based on their distinctive characteristics.
Instruction Building.
A semi-supervised approach, as presented in [nakamoto2023enhancing], leverages LLMs to create datasets for automating the scoring of mathematical self-explanations in the context of mathematics education. A challenging benchmark ProofNet [azerbayev2023proofnet] is introduced to build a system of automated theorem proving. Meanwhile, prompt retrieval and distilled back translation methods are introduced for statement autoformalization. WizardLM [xu_wizardlm_2023] presents a groundbreaking method called Evol-Instruct, designed to autonomously generate high-quality instructions by LLMs themselves. The conceptual flow is depicted in Figure 4. The instructional process commences with a straightforward “initial instruct" and progresses into diverse forms, deepening, increase reasoning, add constraints, concretizing, complicate Input (formula), in-Breadth evolving, etc. Furthermore, Wizardmath [luo2023wizardmath] proposes a reinforced Evol-instruct method (RLEIF) to build more complexity of instruct dataset. This method effectively enhances the mathematical reasoning capabilities of Llama-2 by supervised fine-tuning and PPO training [schulman2017proximal].
Figure 4: Example of Evol-Instruct taken from WizardLM [xu_wizardlm_2023].
Instruction Tuning.
Instruction tuning stands as a potent method to elevate the prowess of large models, aiding in steering the model towards generating outputs that align with human intent [zhang_instruction_2023]. Compared to the pre-training phase of LLMs, instruction tuning demands notably less instruction data and computational resources, making it a prevalent strategy for enhancing a model’s domain-specific capabilities. For instance, MathGLM [yang2023gpt] showcases that even a language model with 2 billion parameters can adeptly execute multi-digit arithmetic operations with limited training data. Goat [liu_goat_2023] discerns patterns in the tokenization of different numbers and proposes a consistent approach for their tokenization. Utilizing the LLaMA model, Goat ensures consistent tokenization for various numbers, illustrated in Figure 5. Similarly, Calculon [muffo_evaluating_nodate] refines the model’s arithmetic skills by employing a digit decomposition technique to construct a fine-tuning dataset, as depicted in Figure 6.
Figure 5: Consistent tokenization for similar number in Goat [liu_goat_2023].
Figure 6: Number decomposition taken from Calculon [muffo_evaluating_nodate].
In PaLM 2-L-Math [liu_improving_2023], three fine-tuning strategies were explored, revealing key insights: 1) The quality and style of step-by-step solutions significantly influence model performance; 2) Combining solution re-ranking and majority voting yields a more substantial performance enhancement compared to using them individually. Additionally, multi-task fine-tuning, separating solution generation and evaluation tasks, surpasses the baseline achieved by solution fine-tuning alone. Moving beyond the enhancement of mathematical capabilities through fine-tuning, LLEMMA [azerbayev8llemma] introduces the Proof-Pile-2 dataset, an amalgamation of mathematical texts and code. Continual pre-training with Code Llama empowers the model to leverage Python interpreters and formal theorem provers, demonstrating impressive performance on the MATH benchmark.
In-context Learning.
In-context Learning (ICL) [brown2020language, kaplan2020scaling] empowers LLMs to execute target tasks by presenting specific task examples as conditions during inference, without updating model parameters. Inspired by this, ScratchpadGPT [nye_show_2021] enhances its proficiency in multi-step computations by mandating the model to output intermediate computation steps into a designated “scratchpad." Codex-math [drori2022neural] refines Codex by generating programs through code and integrates few-shot learning to automatically create programs for solving mathematical problems. This method has substantially improved the previous state-of-the-art accuracy in automatic solutions by 73.1% across various benchmarks like MIT’s mathematics courses, Columbia University’s Computational Linear Algebra, and the MATH benchmark [hendrycks2021measuring]. Notably, few-shot learning contributed to a 10% accuracy boost. Recent studies also highlight the variability in in-context learning performance across different chosen examples [liu2022makes]. Fu et al. [fu2022complexity] and Zhang et al. [zhang2023automatic] selected intricate and diverse examples to enhance reasoning performance.
Figure 7: Use a “scratchpad" to improve arithmetic calculation in ScratchpadGPT [nye_show_2021].
4.2 Tool-based Methods
LLMs are designed to use tools, such as codes and calculators, to enhance their problem-solving abilities [schick2023toolformer, parisi2022talm].
Single-tool Methods.
To improve the performance of mathematical reasoning, math-specific tools such as symbolic solvers and programs are utilized for LLMs [he2023solving, chen2022program]. For example, SymbLLM [he2023solving] solved math word problems by combining language models with symbolic solvers. They focus on automatically generating high-quality, step-by-step solutions to mathematical word problems, especially those encountered in mathematical applications.
Furthermore, previous studies have explored the process of converting mathematical problems into code [chen2022program, drori2022neural]. PoT [chen2022program] proposes a fusion of CoT with programs, while Drori et al. [drori2022neural] showcased the effectiveness of a neural network pre-trained on text and fine-tuned on code, particularly leveraging OpenAI’s Codex transformer. This approach successfully solves, explains and generates math problems at the university level. MathPrompter [imani2023mathprompter] employs a zero-shot chain-of-thought prompting technique to generate multiple algebraic expressions or Python functions in varied ways to solve the same math problem, enhancing reliance on output results. PAL [gao2023pal] introduces an innovative approach to bolster the performance of pre-trained language models (PLMs) in mathematical problem-solving. This involves utilizing LLMs to comprehend natural language problems and generate programs as intermediate reasoning steps.
In addition to solving math problems, programs can also play a role in tutoring math. For instance, Upadhyay explores the role of large language models in tutoring systems [upadhyay-etal-2023-improving], introducing a “code scratchpad" alongside the traditional “language scratchpad" to enhance the model’s performance in tutoring steps, particularly using a grade school mathematics dataset.
Multi-tool Methods.
To facilitate the seamless integration of LLMs with diverse tools, several multi-tool approaches have been proposed to enable LLMs to learn how to use multiple tools simultaneously. Toolformer [schick2023toolformer] adopts a self-supervised training approach to enable the utilization of different tools such as search engines, calculators and translation systems via simple API calls. This is accomplished by fine-tuning on a vast collection of sampled API calls and filtering based on their ability to reduce perplexity on subsequent tokens. Chameleon [lu2023chameleon], on the other hand, enhances LLMs with plug-and-play modules designed for compositional reasoning. It creates programs by combining various tools, including LLMs, off-the-shelf vision models, web search engines, Python functions, and heuristic-based modules, to tackle complex reasoning tasks. The Automatic Reasoning and Tool-use (ART) framework [paranjape2023art] leverages frozen LLMs to automatically generate intermediate reasoning steps as a program. This seamlessly incorporates external tools to support computations that surpass the core capabilities of LLMs. Meanwhile, ToolkenGPT [hao2023toolkengpt] adopts a strategy of learning “toolken" embeddings to represent each tool as a token. This empowers LLMs to effortlessly utilize tools similar to generating word tokens. It caters to a broader range of tools and utilizes extensive demonstration data to learn toolkit embeddings.
Furthermore, CRITIC [gou2023critic] significantly enhances the outputs of LLMs by enabling them to verify and self-correct through interactions with external tools. Inspired by human cognition and critical thinking, CRITIC continuously refines text generated by LLMs through iterative interactions with tools like search engines and code interpreters. Tool Augmented Language Models (TALM) [parisi2022talm] seamlessly integrate language models with non-differentiable tools. They employ an iterative “self-play" method to bootstrap performance based on a limited number of tool demonstrations. In contrast, Hsieh et al. [hsieh2023tool] opted for tools based on documentation rather than relying on demonstrations. Current practices involve teaching LLMs through a few-shot demonstration of tool usage, a process prone to bias and overfitting. The alternative proposed by Hsieh et al., utilizing tool documentation that provides descriptions of individual tool usage, is argued to be a more effective approach. ToRA [gou2023tora] employs a set of Tool-integrated Reasoning Agents, which integrate computational libraries and symbolic solvers, among other mathematical tools, to effectively solve complex mathematical problems.
4.3 Fundamental CoT Methods
A multitude of fundamental Chain-of-Thought (CoT) methods integrated with LLMs have been proposed to enhance the mathematical reasoning abilities of LLMs.
Foundation of CoT.
In the initial stages of research, a limited body of work has leveraged the principles of chain-of-thought (CoT) to enhance the mathematical capabilities of language models. Notably, Ling et al [ling2017program] propose a methodology involving the generation of a series of concise steps, termed "Answer Rationales," to guide the resolution of algebraic word problems. MathQA [amini2019mathqa] suggests decomposing Math Word Problems into multiple steps corresponding to programmatic operations for resolution.
Utilizing the in-context learning of LLMs, CoT [wei2022chain] explicitly introduces the concept of the chain of thought (CoT) for the first time, and substantiates its efficacy in enhancing reasoning abilities. A CoT example is shown in Figure 8. This validation is also corroborated in Complexity-based CoT [fu2022complexity]. Following a similar line of thought, LVSS [lightman2023lets] presents a Process Reward Model (PRM) and compares it with the Outcome Reward Model (ORM). Empirical findings affirm that process supervision outperforms outcome supervision, leading to notable performance improvements on the MATH [hendrycks2021measuring] benchmarks. Consequently, several methods [chen2022program, yuan2023scaling, yu2023thought, wang2023making, sel2023algorithm, yue2023mammoth] leverage CoT methodologies to enhance the mathematical performance of models.
Figure 8: Example of CoT in Wei et al.[wei2022chain].
RFT [yuan2023scaling] proposes the application of rejection sampling finetuning (RFT) to gather more optimal reasoning paths, thereby enhancing mathematical reasoning performance. Thought Propagation [yu2023thought] considers gaining insights from analogous problems to assist in addressing the current issue. Wang et al. [wang2023making] proposed a Fine-Tuning (AFT) paradigm that aligns the model to prioritize the generation of CoT with superior performance. Additionally, PoT [chen2022program] introduces a Program of Thoughts, combining CoT and programming, and AOT [sel2023algorithm] enhances reasoning abilities through algorithmic-style demonstrations of CoT. Furthermore, MAmmoTH [yue2023mammoth] integrates CoT and PoT [chen2022program] rationales to instruct large language models in utilizing code tools for solving mathematical problems.
Construction of CoT.
To streamline the process of CoT creation, various approaches [lu2022dynamic, zhang2023automatic, fu2022complexity, wang2023making, kojima_large_nodate, shum_automatic_2023, pitis_boosted_2023, sel2023algorithm, yu2023thought] have been introduced for its automatic generation.
Zero-shot CoT [kojima_large_nodate] introduces a simple CoT example, “Let’s think step by step," which effectively enhances the model’s reasoning capabilities. Auto-CoT [zhang2023automatic] selects representative questions through clustering and answers them using “Let’s think step by step," concatenating the question-answer pairs one by one to automatically generate CoT. In the case of Complexity-based CoT [fu2022complexity], the demonstration is chosen by simply selecting the reasoning chain with the highest number of inference steps sampled from the model. Using reinforcement learning, PromptPG-CoT [lu2022dynamic] trains a model with policy gradient to assist GPT-3 in selecting suitable CoT demonstrations. Similarly, AutoMate CoT [shum_automatic_2023] employs a variance-reduced policy gradient strategy to estimate the significance of each example in a black box language model, thereby selecting more effective examples. BoostedPrompt [pitis_boosted_2023] presents an iterative prompt ensemble method that enhances prompts when the current demonstration faces challenges in handling specific problems.
4.4 Advanced CoT Methods
To further enhance the capability of CoT in LLMs, advanced CoT methods have been proposed, including Verify-based Methods, Ensemble-based Methods, Planning-based Methods, and Socratic Teaching Methods.
Verify-based Methods.
LLMs often produce incorrect reasoning steps, which can lead to a series of cascading errors in their solutions. Implementing verification and refining the reasoning process based on feedback can significantly reduce these errors. This approach is akin to human reflection and involves critically evaluating each step of the reasoning process. Various methods [zhou2023solving, ling2023deductive, li2023making, zhao2023verify, he_rethinking_2022, shridhar2023screws] have been proposed to address these issues.
Zhou et al. [zhou2023solving] propose a code-based self-verification approach, utilizing a zero-shot prompt to encourage the GPT-4 Code Interpreter to employ code for self-verifying its responses, enhancing its mathematical reasoning capabilities. To mitigate the challenge of validating the entire deductive reasoning process, VerifyCoT [ling2023deductive] introduces a deductive reasoning form, ensuring that each reasoning step strictly relies on the preceding steps. Furthermore, DIVERSE [li2023making] independently verifies each reasoning step and a voting mechanism to eliminate incorrect answers. Both Verify-and-Edit [zhao2023verify] and Retrieval-CoT [he_rethinking_2022] utilizes external retrieval tools to support the model in validating reasoning rationales. The key difference is the former edits rationales using retrieved information, while the latter helps the model self-rethink to improve performance on complex reasoning tasks. Both methods effectively reduce factual mistakes during the reasoning process. Shridhar et al. [shridhar2023screws] summarize the use of the revisions approach in the SCREWS framework, which includes a selection module to choose between the original and modified reasoning steps.
Ensemble-based Methods.
Due to the inherent stochastic nature of language models, which output probability distributions over predicted words, they may randomly generate incorrect reasoning steps and outcomes. To tackle this challenge, some methods [wang2022self, naik2023diversity, fu2022complexity, li2023making, miao2023selfcheck, yoran2023answering, cobbe2021training, khalifa2023discriminator] leverage the concept of ensemble learning, employing techniques such as voting and ranking to eliminate uncertainties in the reasoning process.
Self-Consistency [wang2022self] employs multiple reasoning paths and selects the final response through a simple majority vote. Similarly, Diversity-of-Thought [naik2023diversity] generates diverse reasoning paths by altering prompts and aggregates the responses via the majority vote. Complexity-based CoT[fu2022complexity] favors answers derived from more intricate reasoning paths. DIVERSE [li2023making] uses a weighted voting mechanism to filter out incorrect answers. Nevertheless, these voting-based methods often overlook the potentially useful information within unsuccessful CoT reasoning and lack an efficient integration of multiple reasoning chains to improve performance. Self-check [miao2023selfcheck] addresses this by incorporating reasoning steps into the voting mechanism, ensuring both consistent answers and reliable reasoning. MCR [yoran2023answering] takes a step further by consolidating information across various reasoning chains, acquiring the most relevant facts, facilitating a more comprehensive analysis to make successful reasoning.
Additionally, there are a few methods based on ranking. Rank-verifier [cobbe2021training] proposes using a ranking system to judge the correctness of model completions. Furthermore, GRACE [khalifa2023discriminator] leverages a discriminator trained via contrastive learning to rank each reasoning step.
Planing-based Methods.
The original structure of the Chain of Thought (CoT) is sequential, facing limitations when handling highly complex problems and lacking the ability for retrospective correction. To make the CoT chain structure more systematic and intricate, certain planning-based methods [yao2023tree, mo2023tree, long2023large, yao2023beyond, besta_graph_2023, lei2023boosting, jiang2023resprompt, hao2023reasoning, zhou2023language, liu2023llm+, dagan2023dynamic, madaan2023self, sun2023adaplanner, zhou2023isr, shinn2023reflexion] have been proposed. These methods achieve this by altering the organizational structure of reasoning steps or incorporating mechanisms for refinement and reflection.
The Tree-of-Thought (ToT) [yao2023tree] proposes organizing reasoning paths into a tree structure, where each node represents a reasoning step, and edges denote dependencies between nodes. During the inference process, ToT can use self-assessment to determine future actions. This structure facilitates both forward and backward exploration, employing either deep-first or breadth-first search techniques. TouT [mo2023tree] effectively employs Monte Carlo Dropout to assess uncertainty scores associated with diverse local responses of language models at intermediate steps, enhancing response precision through integration with global search algorithms. Long et al. [long2023large] introduced a multi-module ToT system, where the state of the ToT thought chain process can be stored and retraced through its storage module. Furthermore, Yao et al. [yao2023beyond], Besta et al. [besta_graph_2023], and Lei et al. [lei2023boosting] introduced a novel thought structure called the Graph of Thought (GoT). Similar to ToT, each graph node represents a distinct thought step or state, with edges indicating dependencies between nodes. These methods can effectively integrate diverse reasoning thoughts, fostering collaborative outcomes and improving reasoning by incorporating feedback loops. This contributes to an overall enhancement in the inference capacity of the thought network. Additionally, RESPROMPT [jiang2023resprompt] introduces Residual Connection Prompting, which converts the input prompt into complex reasoning graphs.
Inspired by Monte Carlo tree search, RAP [hao2023reasoning] utilizes the Monte Carlo tree search algorithm to traverse tree structures during the reasoning process. This approach effectively balances global and local searches to obtain a high-reward reasoning path. On the other hand, LATS [zhou2023language] employs LLMs as agents, value functions, and optimizers, repurposing their inherent strengths to enhance decision-making. Liu et al. [liu2023llm+] and Dagan et al. [dagan2023dynamic] introduced LLM+P and LLM+DP, incorporating the generation of Planning Domain Definition Language (PDDL) to break down complex problems and execute plans using specialized models.
Furthermore, the self-refine approach, as proposed by Madaan et al. [madaan2023self] and Sun et al. [sun2023adaplanner], focuses on error correction and summarizing past experiences. Specifically, Self-Refine performs pre-execution validation, providing feedback to the model for refinement in case of errors. Zhou et al. [zhou2023isr] then took a step further by merging the iterative self-refine with PDDL in ISR-LLM. Additionally, Shinn et al. [shinn2023reflexion] introduced re-flexion to rectify errors stemming from previous actions.
Socratic Teaching Methods.
The Socratic teaching method [nelson1980socratic], rooted in the philosophies of ancient Greek thinker Socrates, offers a distinctive perspective when considering the integration of large-scale language models (LLMs) in mathematics education [davies2021advancing]. This approach, emphasizing questioning and dialogue, aims to lead learners toward deep thinking [brickhouse2009socratic]. In field of mathematics education, employing this method can aid students in grasping concepts with clarity, fostering more rigorous logical application, and promoting reflection and self-discovery.
When applied to LLMs like GPT-3, utilizing the Socratic teaching method initiates a guided dialogue between the LLMs and the user [davis1997dialogue]. According to Chang et al. [chang2023prompting], employing techniques such as definition, dialectics, cross-examination, induction, and counterfactual reasoning enables LLMs to gain a deeper understanding of user intentions, thereby assisting users more effectively in addressing queries. These methods not only enhance the model’s comprehension but also, when combined with traditional inductive, deductive, and analogical reasoning techniques, establish a comprehensive analytical framework for tackling mathematical problems [chang2023prompting, alhossami2023language].
The modular framework of Socratic Models (SMs) [zeng2022socratic] significantly broadens the scope of interactive capabilities. SMs facilitate the amalgamation of multiple LLMs to tackle novel multi-modal tasks using structured Socratic dialogues. This approach proves highly effective in integrating diverse inputs seamlessly without necessitating additional training. Expanding on this Socratic teaching, the SocREval method [he2023socreval] focuses on evaluating the performance of LLMs in complex mathematical reasoning tasks. By automatically evaluating the quality of reasoning chains, this method offers a novel reference-free approach to reasoning assessment. This not only reduces time and costs but also enhances the evaluation’s precision.
Insights in this domain are further illuminated by studies like Zeng et al. [zeng2022socratic] and Shridhar et al. [shridhar2022automatic]. Qi et al. [qi2023art] and its counterpart that explores zero-shot multi-modal reasoning. These works delve into how recursive thinking and self-questioning can augment the problem-solving capabilities of LLMs. Moreover, contributions from Ang et al. [ang2023socratic] and Al et al. [al2023socratic] provide invaluable datasets and evaluation methodologies for Socratic questioning. These contributions underscore the practical applications and potential impact of Socratic methods in counseling and coaching, showcasing their diverse applicability.
Overall, the integration of Socratic teaching techniques in LLMs offers novel perspectives and tools for mathematics education. Unlike CoT [wei2022chain], Socratic questioning explicitly guides the thinking process, stimulating effective recursive thinking and displaying greater resilience to errors in the thinking process [qi2023art]. Ranging from basic conversational guidance to intricate problem-solving, these methods demonstrate the effective fusion of modern technology with classical ideas to bolster the learning and comprehension abilities of LLMs.
[…]
VERBPHYSICS [forbes2017verb] 733 1,096 1,828 3,657 EN Calculation Training Formula
Clinical [spithourakis2016numerically, spithourakis2018numeracy] 11,170 1,625 3,220 16,015 EN Calculation Training Formula
Scientific [spithourakis2018numeracy] 14,694 2,037 4,231 20,962 EN Calculation Training Formula
DoQ [elazar2019large] 587 5,418 6,007 12,012 EN Calculation Training Formula
DROP [dua2019drop] 77,409 9,536 9,622 96,567 EN Calculation Training Text
AddSub [hosseini-etal-2014-learning] - - - 395 EN MWP Training Formula
SingleOp [roy2015reasoning] 265 107 159 531 EN MWP Training Formula
SingleEq [koncel2015parsing] - - - 508 EN MWP Training Formula
MultiArith [roy2015solving] 420 - 180 600 EN MWP Training Formula
Alg514 [kushman-etal-2014-learning] - - - 514 EN MWP Training Formula
Math23k [wang2017deep] 22,162 - 1000 23,162 CH MWP Training Formula
AQuA [ling2017program] 97,467 254 254 97,975 EN MWP Training Text
MathQA [amini2019mathqa] 29,837 4,475 28,985 37,297 EN MWP Training Formula
GSM8K [cobbe2021training] 7,473 1,319 8,792 EN MWP Training Text
SVAMP [patel2021nlp] 700 300 1,000 EN MWP Training Formula
DRAW [upadhyay2015draw] - - - 1,000 EN MWP Training Formula
Dolphin1878 [shi2015automatically] - 374 1,504 1,878 EN MWP Training Formula
AllArith [roy2017unit] - - - 831 EN MWP Training Formula
DRAW-1K [upadhyay2017annotating] - - - 1,000 EN MWP Training Formula
HMWP [qin2020semantically] - - - 5,470 CH MWP Training Formula
ArMATH [alghamdi2022armath] - - - 6,000 Arabic MWP Training Formula
TabMWP [lu2022dynamic] - - - 38,431 EN MWP Training Text
Ape210K [zhao2020ape210k] - - - 210,488 CH MWP Training Formula
TAL-SCQ5K-CH  [^1] https://github.com/math-eval/TAL-SCQ5K 3,000 - 2,000 5,000 CH MWP Training Text
TAL-SCQ5K-EN  [^2] https://github.com/math-eval/TAL-SCQ5K 3,000 - 2,000 5,000 EN MWP Training Text
FinQA [chen2021finqa] 6,251 883 1,147 8,281 EN MWP Training Formula
REALFP [kalyan2021much] 185 185 558 928 EN MWP Training Formula
SYNTHFP [kalyan2021much] 10,000 1,000 1,000 12,000 EN MWP Training Formula
TAT-QA [zhu2021tat] - - - 16,552 EN MWP Training Text
MultiHiertt [zhao2022multihiertt] 7,830 1,044 1,566 10,440 EN MWP Training Formula
MML [grabowski2015four] - - - 57,882 EN TP Training Formula
HolStep [kaliszyk2016holstep] 2,013,046 - 196,030 2,209,076 EN TP Training Formula
Feit-Thompson [huang2018gamepad] - - - 83,478 EN TP Training Formula
CoqGym [yang2019learning] - - - 71,000 EN TP Training Formula
HOList [bansal2019holist] - - - 29,462 EN TP Training Formula
IsarStep [li2020isarstep] 820,000 5,000 5,000 830,000 EN TP Training Formula
LISA [jiang2021lisa] - - - 183,000 EN TP Training Formula
NaturalProofs [welleck2021naturalproofs] 32,000 EN TP Training Text
LeanStep [han2021proof] - - - 21,606,000 EN TP Training Formula
NumGLUE [mishra2022numglue] - - - 101,835 EN Calculation Benchmark Text
Dophin18k [huang2016well] - - - 18,460 EN MWP Benchmark Text
MAWPS [koncel2016mawps] - - - 3,320 EN MWP Benchmark Formula
ASDiv [miao2020diverse] - - - 2,305 EN MWP Benchmark Formula
MATH [hendrycks2021measuring] 7,500 5,000 12,500 EN MWP Benchmark Text
MGSM [shi2022language] - - - - Multilingual MWP Benchmark Text
Mathematics [saxton2018analysing] 2,000,000 100,000 2,100,000 EN MWP Benchmark Formula
MMLU-Math [hendryckstest2021] - - - 906 EN WMP Benchmark Formula
AGIEval [zhong2023agieval] - - - 469/220 CH/EN MWP Benchmark Formula
MATH 401 [yuan2023well] - - - 401 EN MWP Benchmark Formula
INT [wu2020int] - - - - EN TP Benchmark Formula
miniF2F [zheng2021minif2f] - 244 244 488 EN TP Benchmark Formula
Aggregate [roy2018mapping] - - - 1,492 EN MWP Augmented Formula
MathQA-Python [austin2021program] 19,209 2,822 2,822 23,914 EN MWP Augmented Code
Math50k [li2023camel] - - - 50,000 EN WMP Augmented Text
PRM800K [lightman2023lets] - - - 2,868 EN WMP Augmented Text
MetaMathQA [yu2023metamath] - - - 395,000 EN MWP Augmented Text
Lila [mishra2022lila] - - - 134,000 EN MWP Augmented Code
PEN [kim2022ept] - - - 3,581 EN MWP Augmented Formula
miniF2F+informal [jiang2022draft] - 244 244 488 EN TP Augmented Formula
NaturalProofs-Gen [welleck2022naturalprover] 12,500 1,000 1,000 14,500 EN TP Augmented Text
[…]
Table 1: The statistics information of mathematical datasets. Solution means the format of the output, such as text, formula and code.
5 Datasets
To train and evaluate the arithmetic and math ability of language models, various math word problems datasets [saxton2018analysing, hendrycks2021measuring, cobbe2021training, kushman-etal-2014-learning] are designed. In this paper, we organize the frequently used datasets for mathematical language models by dividing them into training, benchmark and data augmentation datasets (Table 1), and also sort out the common evaluation metrics.
5.1 Training Datasets
5.1.1 Mathematical Calculation
Several studies have introduced datasets aiming to identify numerical information within text [spithourakis2016numerically, spithourakis2018numeracy, elazar2019large], specifically targeting the prediction of attribute values. The Clinical Data [spithourakis2016numerically, spithourakis2018numeracy] consists of clinical records sourced from the London Chest Hospital. Each patient record includes a text report and structured KB tuples detailing 20 potential numeric attributes like age and gender. Scientific Data [spithourakis2018numeracy] encompasses paragraphs extracted from Cornell’s ARXIV repository, featuring over half a million converted papers across 37 scientific sub-fields. Additionally, Elazar et al. [elazar2019large] proposed the Distributions over Quantities (DoQ) dataset, comprising empirical counts of scalar attribute values linked to more than 350K nouns, adjectives, and verbs across 10 different attributes.
Furthermore, certain studies consider the relational knowledge between various numbers. For instance, the VERBPHYSICS [forbes2017verb] dataset aggregates crowdsourced information about actions and objects, encompassing the relative knowledge of grounded object pairs and implications of actions associated with those objects. More intricately, the DROP [dua2019drop] dataset necessitates Discrete Reasoning Over the content of Paragraphs. In this benchmark of 55,000 adversarially-created questions, a system is tasked with resolving references within a question, potentially spanning multiple input positions, and performing discrete operations like addition, counting, or sorting.
5.1.2 Math Word Problems
A large number of datasets are proposed for MWPs. Hosseini et al. [hosseini-etal-2014-learning] curated the AddSub dataset, primarily focusing on addition and subtraction problems (Figure 9). This dataset identifies relevant variables and their respective values to translate sentences into problem statements, represented as equations. Another dataset, SingleOp [roy2015reasoning], encompasses a wider range of mathematical operations including multiplication and division. Its purpose is to facilitate reasoning about quantities articulated in natural language. DRAW [upadhyay2015draw] comprises 1000 algebra word problems semi-automatically annotated for evaluating automatic solvers. It features gold coefficient alignments crucial for uniquely identifying equation system derivations. The Alg514 [kushman-etal-2014-learning] dataset comprises 514 linear algebra problems structured around 28 equation templates. Meanwhile, MultiArith [roy2015solving] is designed to handle arithmetic problems involving multiple steps and operations, without predefined templates.
Several studies have introduced methods focused on predicting equations through semantic parsing. The SingleEq [koncel2015parsing] dataset revolves around grade-school algebra word problems. It comprises samples that can be correlated to a single equation involving multiple mathematical operations on one variable, structured as a parsing tree (Figure 10). In contrast, the Dolphin1878 [shi2015automatically] dataset contains 1,878 word problems in mathematics, wherein each sample might encompass multiple equations and variables depicted through multiple trees. AllArith [roy2017unit] offers a concise representation of dependencies among number units mentioned in a given problem, referred to as Unit Dependency Graphs (UDGs). To augment dataset size, Math23k [wang2017deep] gathers 23,161 problems labeled with structured equations and corresponding answers. GSM8K [cobbe2021training] comprises 8.5K high-quality grade school math problems crafted by human problem writers. These problems typically involve 2 to 8 steps for resolution, predominantly requiring a sequence of elementary calculations using basic arithmetic operations to arrive at the final solution.
Figure 9: An example of AddSub, taken from [hosseini-etal-2014-learning].
Figure 10: An example of SingleEQ, taken from [roy2015reasoning].
HMWP [qin2020semantically] aims to enhance dataset diversity by extracting math word problems from a Chinese K12 problem bank. This effort sought to validate the universality of math word problem solvers and expand research in MWPs to better align with real-world scenarios. The dataset encompasses three types of MWPs: arithmetic word problems, equation set problems, and non-linear equation problems. In total, it comprises 5,491 MWPs, including 2,955 single-variable linear MWPs, 1,636 two-variable linear MWPs, and 900 single-variable nonlinear MWPs. Additionally, MathQA [amini2019mathqa] presents a diverse collection of 37,000 English multiple-choice math word problems spanning various mathematical domains (Figure 11).
Figure 11: An example of MathQA, taken from [amini2019mathqa].
Existing research endeavors to enhance the comprehensibility and precision of intricate reasoning by providing fine-grained annotations of reasoning processes [upadhyay2017annotating, ling2017program, chen2021finqa, lightman2023lets, zhao2022multihiertt]. DRAW-1K [upadhyay2017annotating] introduces a novel dataset featuring 1,000 general algebra word problems, each annotated with derivations. AQUA [ling2017program] structures each question into four components: problem description, multiple-choice answer options, rationale, and correct option label (Figure 12). FinQA [chen2021finqa] presents an expert-annotated dataset comprising 8,281 financial QA pairs, meticulously annotated with numerical reasoning processes to ensure comprehensive explication. PRM800K [lightman2023lets] assigns labels (positive, negative, or neutral) to each step in solving MATH problems sampled by a large-scale generator. The training set encompasses 800K step-level labels across 75K solutions to 12K problems.
Figure 12: An example of AQUA, taken from [ling2017program].
Additionally, Patel et al. [patel2021nlp] introduced the SVAMP challenge set to establish a more robust evaluation framework for methods designed to tackle elementary-level math word problems. Alghamdi et al. [alghamdi2022armath] contributed the inaugural large-scale dataset, ArMATH, for Arabic MWPs. This dataset comprises 6,000 samples of primary-school math problems written in Modern Standard Arabic. Kalyan et al. [kalyan2021much] proposed two datasets, REALFP and SYNTHFP, for a novel task called Fermi Problems (FPs). FPs entail questions whose answers can only be approximately estimated due to the impracticality or impossibility of precise computation. This task necessitates the amalgamation of various reasoning abilities, including suitable abstractions, commonsense knowledge, and the creative synthesis of problem-solving strategies.
Recently, mathematical reasoning has incorporated multiple input modalities such as textual and tabular data [lu2022dynamic, zhu2021tat, zhao2022multihiertt]. For instance, TabMWP [lu2022dynamic] presents a new dataset comprising 38,431 open-domain grade-level problems that necessitate mathematical reasoning across both textual and tabular data. Zhu et al. [zhu2021tat] introduced a large-scale QA dataset, TAT-QA, encompassing both tabular and textual data. This dataset often requires numerical reasoning—such as addition, subtraction, multiplication, division, counting, comparison/sorting, and their combinations—based on real financial reports to infer answers. Additionally, MultiHiertt [zhao2022multihiertt] is a proposed dataset consisting of QA pairs derived from Multi Hierarchical Tabular and Textual data (Figure 13). Each document in this dataset contains multiple tables and longer unstructured texts, accompanied by fine-grained annotations of reasoning processes and supporting facts to elucidate complex numerical reasoning.
Figure 13: An example of MultiHiertt, taken from [zhao2022multihiertt].
5.1.3 Theorem Proving
Current research delves into exploring mathematical theorem datasets (e.g., INT [wu2020int], Feit-Thompson [huang2018gamepad], and IsarStep [li2020isarstep]) to develop novel machine learning-based theorem-proving strategies. For instance, the format of IsarStep is shown in 14. Huang et al. [huang2018gamepad] introduced the Feit-Thompson dataset, encompassing 1,602 lemmas, expanding into 83,478 proof states for the Feit-Thompson theorem [gonthier2013machine]. IsarStep [li2020isarstep] constitutes a non-synthetic dataset extracted from the largest repository of proofs handwritten by human experts in a theorem prover. Furthermore, NaturalProofs [welleck2021naturalproofs] comprises 32,000 theorem statements and proofs, 14,000 definitions, and 2,000 other types of pages (e.g., axioms, corollaries). This dataset draws from three domains: broad-coverage data sourced from ProofWiki, deep-coverage data from the Stacks project, and low-resource data extracted from mathematics textbooks.
Figure 14: An example of IsarStep, taken from [li2020isarstep].
Several formal mathematical libraries [megill2019metamath, rudnicki1992overview, grabowski2015four] encompass a range of formal languages tailored for theorem proving (TP). Examples include MML [grabowski2015four], Coq [bertot2013interactive], Lean [han2021proof], Isabelle [wenzel2008isabelle], and Hol light [slind2008brief]. For instance, MML [grabowski2015four], also known as the Mizar Mathematical Library, was established to systematically construct a centralized, reusable knowledge base reflecting the standard foundations of mathematics, namely classical first-order logic and set theory. Yang et al. [yang2019learning] compiled the CoqGym dataset, comprising 71,000 human-written proofs from 123 projects developed using the Coq proof assistant. LeanStep [han2021proof] serves as the tactic proof dataset for the Lean theorem prover, featuring high-level human-written tactics alongside kernel-level proof terms. LISA [jiang2021lisa], one of the largest proof corpora for interactive theorem provers, contains 183,000 theorems and 2.16 million proof steps. This dataset is extracted from the Archive of Formal Proofs and the standard library of Isabelle. Kaliszyk et al. [kaliszyk2016holstep] introduced the HolStep dataset rooted in Higher-Order Logic (HOL) proofs. HOList [bansal2019holist] includes almost 30,000 theorems and proofs across three corpora: core, complex, and flyspeck. The core corpus contains fundamental theorems necessary for defining tactics, while the complex corpus consists of theorems related to complex calculus. Flyspeck contains the majority of the lemmas and theorems related to the Kepler conjecture [hales2017formal].
5.2 Benchmark Datasets
Various benchmark datasets have been introduced to assess the performance of different mathematical algorithms, such as MAWPS [koncel2016mawps]. Huang et al. [huang2016well] developed the Dolphin18K dataset, a large-scale and diverse collection containing over 18,000 annotated math word problems. The Mathematics dataset [saxton2018analysing] encompasses problems spanning arithmetic, algebra, probability, and calculus. These problems involve sequential questions and answers presented in a free-form textual input/output format. For assessing arithmetic understanding, Mishra et al. [mishra2022numglue] introduced NumGLUE, a multi-task benchmark that evaluates AI systems across eight distinct tasks. miniF2F [zheng2021minif2f] is a dataset designed to create a unified cross-system benchmark for neural theorem proving. It comprises 488 formal Olympiad-level mathematics problem statements, encompassing Metamath, Lean, Isabelle, and HOL Light.
Several datasets consider varying difficulty levels across educational stages, ranging from elementary school [miao2020diverse, shi2022language] to high school [hendrycks2021measuring]. Miao et al. [miao2020diverse] introduced the ASDiv corpus, comprising 2,305 math word problems (MWPs) categorized by problem type and elementary school grade level. The Multilingual Grade School Math Benchmark (MGSM) [shi2022language] leverages problems from GSM8K, translating them into 10 languages with the assistance of human annotators. The MATH dataset [hendrycks2021measuring] contains 12,500 problems sourced from high school math competitions. Each problem within MATH is equipped with a step-by-step solution, enabling models to learn to generate answer derivations and explanations.
Furthermore, several datasets serve to evaluate foundation models [zhong2023agieval, yuan2023well]. AGIEval [zhong2023agieval] functions as a human-centric benchmark tailored to assess foundation models’ general abilities in tasks related to human cognition and problem-solving. This dataset is collected from standardized human-centric exams like college entrance exams, law school admission tests, math competitions, and lawyer qualification tests. Yuan et al. [yuan2023well] introduced the arithmetic dataset MATH 401, designed to test the latest large language models such as GPT-4, ChatGPT, InstrctGPT, Galactica, and LLaMA. It features various arithmetic expressions and offers a comprehensive analysis of the capabilities of large language models.
5.3 Augmented Datasets
Augmented datasets serve to enhance existing datasets by incorporating additional samples or information. Roy et al. [roy2018mapping] introduced Aggregate by augmenting the AllArith dataset with 661 word problems from Perturb. Perturb contains problems that were manually pruned either for not yielding the desired solution $a$ or being too dissimilar from the input problem $p$. Large-scale language models like ChatGPT and GPT-4 have been utilized to expand datasets. For instance, MetaMathQA [yu2023metamath] rephrases questions from multiple perspectives without introducing extra knowledge to bootstrap mathematical questions. The Math5K dataset [li2023camel] comprises 50,000 problem-solution pairs generated using GPT-4. These datasets leverage advanced language models to enrich the existing collection of problems and solutions.
Several studies have expanded existing datasets by including supplementary information, such as code programs [austin2021program, mishra2022lila] and explanations [kim2022ept]. MathQA-Python [austin2021program] is a Python adaptation of the MathQA benchmark, comprising 23,914 problems. This dataset aims to evaluate models’ capabilities in synthesizing code from intricate textual descriptions. Lila [mishra2022lila] is a comprehensive mathematical reasoning benchmark composed of 23 diverse tasks. It extends 20 existing datasets by gathering task instructions and solutions as Python programs, offering not only correct answers but also explainable solutions. PEN [kim2022ept] focuses on providing plausible and accurate explanations for solving algebraic word problems present in three benchmark datasets: ALG514, DRAW-1K, and MAWPS. Moreover, some studies have contributed statements and proofs based on existing datasets for theorem proving [jiang2022draft, welleck2022naturalprover]. For example, miniF2F+informal [jiang2022draft] is a dataset consisting of manually curated informal statements and proofs aligned with formal statements from the miniF2F dataset. NaturalProofs-Gen [welleck2022naturalprover] adapts data from NATURALPROOFS, including theorem statements, proofs, definitions, and additional pages sourced from ProofWiki. Each proof follows a multi-step structure and references a variable number of sources.
6 Challenges and Further Directions
Faithfulness.
Mathematical LLMs reside in the phenomena of hallucination and faithfulness, where the generated output may lack factual accuracy or logical grounding [rawte2023survey, ji2023survey]. This phenomenon leads to the production of erroneous or misleading mathematical results, undermining the reliability of the model’s outputs. Some studies explored to address this problem by integrating extra knowledge [he_rethinking_2022], reinforcement learning from human feedback [lightman2023lets], tools [schick2023toolformer, parisi2022talm]. However, the improvement of the hallucination phenomena is limited, which influences the trustworthiness and utility of mathematical language models in practical applications and scholarly pursuits.
Multi-modal.
In particular, math problems (e.g., geometry problems [gelernter1960empirical]) involve not only textual information but also various modalities such as diagrams, graphs, or mathematical symbols [lu2021iconqa]. While existing models excel in processing textual information, interpreting and reasoning across multiple modalities concurrently remains a formidable task. Few multi-modal mathematical datasets and methods are proposed for this task [zhao2022multihiertt, lu2021iconqa]. Bridging the gap between textual representations and visual/mathematical elements necessitates robust mechanisms that can effectively capture and synthesize information from disparate sources, ultimately leading to comprehensive and accurate problem-solving capabilities. In fact, the multi-modal information in mathematics is much more complex than general multi-modal tasks like vision question answering [antol2015vqa] and image captioning [hossain2019comprehensive]. Achieving proficiency in handling multi-modal mathematical problems stands as a pivotal yet intricate objective in advancing the competency of mathematical LLMs.
Uncertainty.
The uncertainty of LLMs [gawlikowski2023survey, duan2023shifting] leads to the ambiguity and variability problem of mathematical problems. While these models excel in deterministic tasks, their handling of uncertainty, such as probabilistic reasoning or dealing with incomplete or vague information, poses a significant challenge. Mathematical problems often entail nuanced interpretations, fuzzy constraints, or scenarios where a single precise solution might not exist. Several studies investigated this problem via controlled generation technologies [DBLP:conf/icml/ZhouJWCS23]. However, ensuring that LLMs can navigate and appropriately account for these uncertainties while providing accurate and contextually relevant solutions remains a complex task.
Evaluation.
It is still a challenge to evaluate mathematical LMs with robust and comprehensive evaluation metrics that adequately capture the models’ performance across various mathematical tasks. Traditional evaluation metrics in natural language processing might not effectively capture the intricacies of mathematical reasoning and problem-solving. Designing evaluation benchmarks [saxton2018analysing, hendrycks2021measuring, cobbe2021training, kushman-etal-2014-learning] and metrics [wang2021gpt, chung2022scaling, thoppilan2022lamda, yuan2023well] that encompass a wide spectrum of mathematical tasks, spanning from basic arithmetic to complex theorem proving, while accounting for linguistic fluency and mathematical accuracy, remains a significant challenge. Addressing these challenges is crucial to ascertain the reliability, efficacy, and generalizability of mathematical LMs, fostering advancements in this burgeoning field.
Creation.
While previous models exhibit remarkable capabilities in understanding and manipulating existing mathematical concepts, their ability to autonomously devise and rigorously prove entirely new theorems presents a formidable hurdle. The development of novel mathematical theorems demands not only profound mathematical reasoning but also creative and insightful problem-solving abilities, aspects that necessitate a deeper understanding of abstract mathematical concepts beyond what the models have been trained on. Davies et al. [davies2021advancing] applied machine learning techniques to discover potential patterns and relationships among mathematical entities. Their approach involved employing attribution techniques to comprehend these connections, leveraging these insights to foster intuition and put forward conjectures. Bridging the gap between the models’ proficiency in assimilating existing mathematical knowledge and their capability to generate novel, verifiable, and impactful theorems represents a significant frontier in leveraging LLMs for the advancement of mathematical knowledge and discovery.
Application.
While mathematical LMs exhibit promising potential in autonomously solving math problems, their deployment in educational settings as teaching aids or tutors necessitates addressing several pivotal challenges. Tailoring these models to serve as effective educational tools demands not only mathematical proficiency but also adeptness in pedagogy and instructional methodologies. Few studies applied Socratic questioning for mathematical teaching [zeng2022socratic, shridhar2022automatic]. Customizing LLMs to cater to diverse learning styles, adapting explanations to suit different proficiency levels, and fostering an interactive and engaging learning environment is interesting.
7 Conclusions
The survey elucidates the pivotal role of mathematical LMs in reshaping the landscape of mathematical problem-solving, leveraging a spectrum of models, from pre-trained language models (PLMs) to large-scale language models (LLMs), to address diverse mathematical tasks. Our taxonomical delineation of mathematical tasks and algorithms provides a systematic framework for comprehending the intricacies of LMs-based methodologies, distinguishing between arithmetic calculation, mathematical reasoning, and various algorithmic approaches employed in these models. The compilation of over 60 diverse mathematical datasets, categorized meticulously into training, benchmark, and augmented datasets, underscores the pivotal role of data in advancing mathematical research, facilitating informed research endeavors within distinct mathematical contexts. Moreover, by critically addressing challenges such as faithfulness, multi-modality, uncertainty, evaluation, theorem creation, and application, this survey paves the way for future investigations aimed at refining and advancing mathematical LMs capabilities. By shedding light on the current state-of-the-art, challenges, and avenues for future exploration, we envision this comprehensive overview to be a cornerstone in driving innovation and shaping the trajectory of mathematical LMs research, ultimately contributing to the evolving landscape of mathematics and artificial intelligence.
[…]