### v4

# Mathematical Language Models: A Survey

Abstract.
In recent years, there has been remarkable progress in leveraging Language Models (LMs), encompassing Pre-trained Language Models (PLMs) and Large-scale Language Models (LLMs), within the domain of mathematics. This paper conducts a comprehensive survey of mathematical LMs, systematically categorizing pivotal research endeavors from two distinct perspectives: tasks and methodologies. The landscape reveals a large number of proposed mathematical LLMs, which are further delineated into instruction learning, tool-based methods, fundamental CoT techniques, advanced CoT methodologies and multi-modal methods. To comprehend the benefits of mathematical LMs more thoroughly, we carry out an in-depth contrast of their characteristics and performance. In addition, our survey entails the compilation of over 60 mathematical datasets, including training datasets, benchmark datasets, and augmented datasets. Addressing the primary challenges and delineating future trajectories within the field of mathematical LMs, this survey is poised to facilitate and inspire future innovation among researchers invested in advancing this domain.
Mathematics, Language Models, Pre-trained, LLMs, Survey
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Natural language generationccs: Computing methodologies Scene understandingccs: Computing methodologies Cognitive roboticsccs: Computing methodologies Cognitive scienceccs: Computing methodologies Intelligent agents
1. Introduction
Mathematics is the queen of sciences.
– Carl Friedrich Gauss
Mathematics stands as a foundational skill integral to human intelligence, wielding significance across diverse fields such as natural sciences, engineering, medicine, finance, computer science, and social sciences. Within the field of natural language processing (NLP), the development of computer models aimed at autonomously resolving mathematical word problems has captivated researchers since as early as 1963 (Feigenbaum et al., 1963; Bobrow et al., 1964; Briars and Larkin, 1984; Fletcher, 1985). We believe that addressing this problem is a potential way toward illuminating pathways for general reasoning mechanisms, consequently advancing the pursuit of artificial general intelligence (AGI).
One main traditional solution to math word problems is statistical learning-based methodologies (Hosseini et al., 2014; Kushman et al., 2014; Zhou et al., 2015; Mitra and Baral, 2016). Notably, machine learning techniques (Kushman et al., 2014; Roy et al., 2015; Roy and Roth, 2015) alongside semantic parsing methods (Shi et al., 2015; Koncel-Kedziorski et al., 2015) have been employed to address this challenge, showcasing promising outcomes on certain datasets. The evolution of deep learning has prompted considerable interest in crafting neural networks capable of resolving mathematical problems (Wang et al., 2017; Couperus, 2023). Notably, recent years have witnessed significant advancements in the domain of mathematical AI (Wei et al., 2022; Matzakos et al., 2023), propelled by the emergence of powerful language models (LMs) (Kenton and Toutanova, 2019; Touvron et al., 2023a). These language models, comprising pre-trained language models (PLMs) and large-scale language models (LLMs), have assumed a central role in reshaping the landscape of mathematical exploration and practical applications. The paramount focus of this comprehensive survey lies in assessing the impact of these models on the field of mathematics. The survey endeavors to provide a thorough overview of extant research and its consequential implications.
PLMs such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020), GPT-1 (Radford et al., 2018) and GPT-2 (Radford et al., 2019) undergo pre-training on extensive textual corpora to assimilate worldly knowledge. To enhance mathematical performance, certain endeavors focus on either pre-training or fine-tuning PLMs using mathematical datasets (Clark et al., 2021; Geva et al., 2020; Feng et al., 2022). For instance, methodologies like GenBERT (Geva et al., 2020), NF-NSM (Feng et al., 2022), MathBERT(Peng et al., 2021) and LISA (Jiang et al., 2021) incorporate numerical data or mathematical formulas into PLMs to augment their capabilities. Moreover, specialized modules or tailored loss functions are devised, leveraging existing PLMs to learn mathematical reasoning and operations (Zhu et al., 2021; Zhao et al., 2022; Jie et al., 2022; Li et al., 2022).
The recent advent of LLMs, exemplified by OpenAI’s GPT-4 (OpenAI, 2023), has catalyzed an unforeseen surge in innovation, underscoring the multifaceted potential of AI within the domain of mathematics. These models ((Touvron et al., 2023a, b)) have demonstrated remarkable success across diverse Natural Language Processing (NLP) tasks, leveraging in-context learning (Min et al., 2022; Brown et al., 2020; Chen et al., 2022b) and instruction learning (Yang et al., 2023; Liu and Low, 2023). Recent studies by Wang et al. (Wang et al., 2023d) indicate that LLMs featuring over 100 billion parameters (e.g., GPT-3 (Brown et al., 2020) with 175 billion, PaLM (Chowdhery et al., 2022) with 540 billion) exhibit the capability to address intricate tasks by employing a chain-of-thought (CoT) mechanism (Wei et al., 2022) when furnished with a limited set of reasoning examples as demonstrations. Advancements in CoT frameworks (Chen et al., 2022a; Lu et al., 2023c; Zhang et al., 2023e; Fu et al., 2023) have been tailored to enhance mathematical performance, incorporating tools and programs (Gao et al., 2023; Drori et al., 2022; He-Yueya et al., 2023). Moreover, openly accessible LLMs are specifically designed for mathematical tasks, such as LLEMMA (Azerbayev et al., [n. d.]), Qwen-Math (Yang et al., 2024) and InternLM-Math (Ying et al., 2024). Recently, o1 (OpenAI, 2024) and o3 obtained state-of-the-art performance on mathematical reasoning by integrating reinforcement learning with the Monte Carlo tree.
We mainly review the most relevant surveys about mathematical reasoning and LMs (Min et al., 2023). For surveys focused on mathematical reasoning, Lu et al. (Lu et al., 2023d) delineate the landscape of deep learning applications specifically pertaining to mathematical reasoning (DL4Math). Qiao et al. (Qiao et al., 2023) predominantly delve into general reasoning mechanisms centered around language models, encompassing arithmetic, commonsense, logical, symbolic, and multi-modal reasoning (LM4Reasoning). Additionally, Chu et al. (Chu et al., 2023) offer an exploration of studies focused on CoT reasoning. For surveys related to LMs, Qiu et al. (Qiu et al., 2020) and Zhao et al. (Zhao et al., 2023b) contribute comprehensive reviews elucidating the realm of existing PLMs and LLMs, respectively. Yan et al. (Yan et al., 2024) summary the related studies about multimodal LLMs-based mathematical reasoning (MMLM4Math).
In contrast to these existing works, our study shifts focus to the emerging field of mathematical language models (MLMs), presenting a systematic and in-depth examination of the novel methods, tasks, and datasets that have emerged specifically within this domain. While previous surveys primarily address general reasoning capabilities of LMs or broader trends in LMs, our work distinguishes itself by offering a detailed exploration of how LMs are being uniquely applied to advanced mathematical tasks. We not only discuss the successes and challenges of existing approaches but also identify critical gaps and emerging trends that are shaping the future of mathematical reasoning. By providing a comprehensive overview of mathematical language models, including their applications, challenges, and future directions, our work aims to inspire further research and innovation in this rapidly developing area. We believe that highlighting the intersections between mathematics and LMs will encourage new approaches for solving complex mathematical problems, and ultimately, contribute to the broader goal of leveraging language models to revolutionize mathematics and its real-world applications.
Specifically, our categorization of mathematical tasks (§2) contain two primary domains: mathematical calculation (§2.1), consisting of arithmetic representation and arithmetic calculation, and mathematical reasoning (§2.2), consisting of math problem-solving and theorem proving. Furthermore, our taxonomy of existing algorithms segregates them into PLMs-based approaches (§3), including autoregression (§3.1) and non-autoregression (§3.2) LMs, and LLMs-based methodologies (§4), which include instruction learning (§4.1), tool-based strategies (§4.2), fundamental CoT techniques (§4.3), advanced CoT methodologies (§4.4) and multi-modal methods (§4.5). Moreover, we provide a comprehensive compilation of over 60 mathematical datasets (§5), systematically categorized into training (§5.1), benchmark (§5.2), and augmented datasets (§5.3). This classification aims to facilitate research by delineating the utility and applicability of these datasets within distinct research contexts. Also, we analyze and discuss the characteristics and performance of various mathematical LMs (§6). Lastly, our survey meticulously addresses the primary challenges and explores the future directions of this domain (§7), including faithfulness, multi-modal, uncertainty, evaluation, creation, application, and data scarcity. In conclusion (§8), we believe that this comprehensive survey will serve as a valuable asset for researchers and practitioners alike, seeking to push the boundaries of innovation in the field of mathematical language models.
2. Mathematical Tasks
In this section, we summarize the existing mathematical tasks into mathematical calculation and mathematical reasoning (Figure 1, 2).
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=6pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.6em,font=,, where level=2text width=9.4em,font=,, where level=3text width=9.5em,font=,, where level=4text width=12em,font=,, [ Mathematical Tasks, ver [ Mathematical
Calculation  (§2.1) [ Arithmetic
Representation [ DigitRNN  (Spithourakis and Riedel, 2018), DigitCNN  (Wallace et al., 2019), GenBERT  (Geva et al., 2020), NumBERT  (Zhang et al., 2020a) , leaf, text width=44.6em ] ] [ Arithmetic
Calculation [ ScratchpadGPT  (Nye et al., 2021), Goat  (Liu and Low, 2023), MathGLM  (Yang et al., 2023), , leaf, text width=44.6em ] ] ] [ Mathematical
Reasoning  (§2.2) [ Math Problem
Solving [ MathPrompter  (Imani et al., 2023), MetaMath  (Yu et al., 2023b), WizardMath  (Luo et al., 2023), MathAttack  (Zhou et al., 2023c), LLEMMA  (Azerbayev et al., [n. d.]), GEOS  (Seo et al., 2015),
DROP  (Dua et al., 2019), Mathematics  (Saxton et al., 2019), Lila  (Mishra et al., 2022a)
, leaf, text width=44.6em ] ] [ Theorem
Proving [ DeepMath  (Irving et al., 2016), ASTactic  (Yang and Deng, 2019), NaturalProofs  (Welleck et al., 2021), INT  (Wu et al., 2021) , leaf, text width=44.6em ] ] ] ]
Figure 1. Taxonomy of mathematical tasks.
Figure 2. Examples of mathematical tasks, where (a) illustrates responses with different representation types.
2.1. Mathematical Calculation
The advent of LMs has ushered in a new era of exploration into their computational capabilities, particularly in arithmetic. Initially, these models demonstrated basic computational abilities by representing numbers in textual formats. With the increasing capabilities of LMs, it has been observed that they can acquire arithmetic skills through methods such as fine-tuning, even without carefully crafted numeric representations.
2.1.1. Arithmetic Representation
In the early stages of the research, numerical values were either omitted or oversimplified, treated as ordinary text, or categorized as “unknown”. These approaches, however, proved inadequate for mathematical tasks. For instance, BERT performs five times worse when the answer is a number rather than a textual span on the DROP benchmark (Dua et al., 2019).
Recent literature proposes various methods for numerical representation. Geva et al. (Geva et al., 2020) introduce GenBERT, which tokenizes numbers at the digit level and undergoes fine-tuning with arithmetic word problems and simple arithmetic tasks. Zhang et al. (Zhang et al., 2020b) experiment with a retrained BERT model by converting numbers into scientific notation (e.g., 314.1 as 3141[EXP]2). Spithourakis and Riedel (Spithourakis and Riedel, 2018), along with Wallace et al. (Wallace et al., 2019), explore the integration of digit embeddings into a singular embedding that represents the entire number. Berg-Kirkpatrick and Spokoyny (Berg-Kirkpatrick and Spokoyny, 2020) propose a method with digit-RNN and exponent embeddings. This method specifically emphasizes the exponent while disregarding the mantissa. Goat (Liu and Low, 2023) introduces Consistent Tokenization, enhancing the relationships between similar numerical values.
2.1.2. Arithmetic Calculation
There has been significant research into the arithmetic capabilities of LLMs. Nogueira et al. (Nogueira et al., 2021) and Wang et al. (Wang et al., 2021) assessed addition and subtraction tasks. Muffo et al. (Muffo et al., [n. d.]) evaluated two-digit multiplication. Yuan et al. (Yuan et al., 2023b) evaluated the arithmetic operation capabilities of different models, including GPT-4 (OpenAI, 2023), Galactica (Taylor et al., 2022), and LLaMA (Touvron et al., 2023a).
Traditionally, it was presumed that LLMs could not accurately perform complex arithmetic, especially multiplication involving more than eight digits. However, recent approaches challenge this assumption. Zhou et al. (Zhou et al., 2022) apply specialized prompt engineering to enhance addition capabilities but note limitations in multiplication beyond seven digits. Jelassi et al. (Jelassi et al., 2023) investigate length generalization in basic arithmetic tasks using techniques like relative position embeddings and training set priming. ScratchpadGPT (Nye et al., 2021) demonstrates the effectiveness of pre-generating a Chain of Thought (CoT) before producing an answer in 8-digit addition tasks. Goat (Liu and Low, 2023) utilizes supervised instruction for fine-tuning elementary arithmetic operations with large integers. MathGLM (Yang et al., 2023) excels in intricate arithmetic tasks by pre-training on datasets that decompose complex arithmetic expressions into simpler steps.
2.2. Mathematical Reasoning
Mathematical reasoning is a pivotal aspect of artificial intelligence that facilitates the understanding and solving of complex mathematical problems. The integration of LLMs in this domain has been significant, owing to their ability to interpret, process, and generate complex natural text. This section delves into the current research and developments within this field, focusing on two main areas: math problem solving and theorem proving.
2.2.1. Math Problem Solving
Math problem solving refers to the process of using algorithms, computational models, and increasingly LLMs to understand, explain, and solve mathematical problems. These approaches to solving problems spans all levels from basic arithmetic to advanced mathematics, including but not limited to algebra, geometry, statistics, and calculus. Generally, mathematical problem solving can be categorized into two primary forms: the first involves contextualized problems, where the solver must comprehend the given scenario, translate it into a purely mathematical problem, and then proceed to solve it with solutions; the second form presents the solver with a direct mathematical problem, requiring only the provision of a specific answer.
Math Word Problems (MWPs) refer to problems that present mathematical concepts and calculations in the form of written descriptions. Such questions usually include one or more story situations that describe a series of events or conditions from which the solver needs to extract relevant mathematical information and apply appropriate mathematical principles to solve the problem. Over the years, the use of various efficient and intelligent algorithms to solve MWPs has also been a research focus in the fields of computer science and artificial intelligence (Bobrow et al., 1964). Recent research has highlighted the growing capabilities of LLMs in MWPs task, emphasizing the trend toward more nuanced and sophisticated AI-driven mathematical analysis. MathPrompter (Imani et al., 2023) uses a GPT3 DaVinci to solve MWPs with excellent results, demonstrating the potential of LLMs to not only explain but also generate complex mathematical reasoning. Yuan et al. (Yuan et al., 2023a) investigated the interaction between various factors such as pre-training loss, amount of supervised data, and augmented data to improve the mathematical inference performance of LLMs. They proposed Rejection sampling Fine-Tuning (RFT) to enhance performance by using supervised models to generate and collect correct reasoning paths, thereby achieving better inference generalization.
MetaMath (Yu et al., 2023b) further extends the practicality of LLMs, proposing a paradigm by which LLMs generate their mathematical problems, thus creating a self-sustaining learning environment that encourages continuous improvement of the model’s problem-solving acuity. Similarly, WizardMath (Luo et al., 2023) explores new ways to enhance LLMs’ mathematical reasoning capabilities by reinforcing evolutionary instructions, indicating an important step towards the autonomous self-improvement of models. Instead, MathAttack (Zhou et al., 2023c) introduces a critical perspective by exploring the sensitivity of LLMs to specialized adversarial input designed to test their mathematical problem-solving abilities. Such investigations are crucial to developing more resilient and reliable LLMs that can withstand and adapt to a variety of problem-solving situations. Furthermore, LLEMMA (Azerbayev et al., [n. d.]), an open-source LM dedicated to mathematics, outlines ongoing efforts to make LLMs accessible to a wide range of mathematical inquiries, from basic problem solving to advanced theorem proving.
Correspondingly, if the problems are non-contextualized and purely mathematical, the focus shifts more towards the Math Question-Answer (MQA) system’s ability to understand and automatically solve the related computational tasks. These questions can range from simple arithmetic problems to complex high-school or college-level mathematics, including algebra, calculus, and geometry. The challenge for MQA systems is to accurately interpret the text, convert it into an appropriate mathematical representation, perform the necessary computations, and generate a correct and often step-by-step solution, mimicking the problem-solving process of a human.
Sachan et al. (Saxton et al., 2019) introduce a framework for evaluating neural architectures through a suite of math problems, highlighting the importance of algebraic generalization. This is complemented by GEOS (Seo et al., 2015), an innovative system that solves SAT geometry problems by integrating text understanding with diagram interpretation. The introduction of datasets and systems like Inter-GPS (Lu et al., 2021a), IconQA (Lu et al., 2021b), and PGDP5K (Hao et al., 2022; Zhang et al., 2022, 2023d) further advances the field, offering new benchmarks for abstract diagram understanding and visual language reasoning. Additionally, Scharpf et al. (Scharpf et al., 2022) indicate that unsupervised methods for formula labeling contribute to the automated understanding of mathematical documents. Collectively, these works demonstrate the evolving landscape of MQA, where the synergy between natural language processing, computer vision, and symbolic reasoning opens new avenues for educational technology and AI research.
2.2.2. Theorem Proving
Theorem proving (TP) refers to the process of demonstrating that a statement is correct based on existing axioms and facts, which is a long-term challenge in AI (Newell et al., 1957). In the context of LLMs, people began to try to use the huge knowledge base of LLMs and some manual guidance to solve this task. LLMs like GPT-3 (Brown et al., 2020) can process natural text to understand the premises of the theorem, and apply certain logical reasoning rules to complete the proof requirements. This can then be done by calling relevant tools or using manual inspection to check the correctness of the process. The advanced approach aims to leverage LLM’s superior computing power and vast knowledge base relative to humans to automate the theorem-proving process, a task that previously required professional mathematicians.
In the field of AI, research on theorem proving has progressed from data sets containing large numbers of human-written proofs (such as CoqGym (Yang and Deng, 2019)) to complex models that can autonomously generate proof strategies (such as ASTactic (Yang and Deng, 2019)). This improvement is reflected in the application of some language models that are based on Transformer (Vaswani et al., 2017) to TP tasks, especially in recent research such as GPT-f (Polu and Sutskever, 2020a). Most notably, some of the proofs generated by these models have been formally recognized by the mathematical community. Jiang et al. (Jiang et al., 2023b) further refined this process by using informal proofs to guide TP models. NaturalProofs (Welleck et al., 2021) extends these capabilities by leveraging the language of natural mathematics to create a rich corpus of model training and evaluation. At the same time, DeepMath (Irving et al., 2016) and INT (Wu et al., 2021) have promoted the development of this field by demonstrating the effectiveness of neural sequence models for premise selection and evaluating the LLMs’ generalization ability in theorem proving.
Ongoing research is delving into the integration of LMs and interactive proof assistants within the realm of mathematics. This involves generating proofs through tactic prediction (Polu and Sutskever, 2020b; Han et al., 2022; Lample et al., 2022; Jiang et al., 2022), automating formalization processes (Wu et al., 2022; Jiang et al., 2023b), and developing unified tools (Welleck and Saha, 2023). Due to the substantial computational demands of exploration, LMs applied in this domain have traditionally been limited in size. However, recent advancements have showcased potential in employing larger models (First et al., 2023; Jiang et al., 2023b). LLEMMA (Azerbayev et al., [n. d.]) demonstrates few-shot proof autoformalization and tactic prediction, offering a vast dataset of formal mathematics and an openly accessible model for further investigation in these directions. These endeavors establish crucial theoretical and practical frameworks for automating and enhancing tactic prediction.
The above series of progressive research work shows that utilizing the powerful computing power and extensive knowledge base of LLMs is the main development direction for future TP tasks. The more excellent models (such as GPT-4 (OpenAI, 2023)) obviously have better reasoning capabilities without any human intervention, but we also need to be clear about one thing, that is the inevitable hallucination problem in neural text generation models (Maynez et al., 2020). The hallucination problem refers to the generated content that is nonsensical or unfaithful to the provided source content (Ji et al., 2023). The occurrence of hallucinations at any step in the reasoning process will lead to errors in the final reasoning results. Therefore, solving the hallucination problem is also particularly important in future work related to TP.
3. PLMs-based Methods
Based on the transformer architecture (Vaswani et al., 2017) with self-attention mechanisms, the “pre-training and fine-tuning” paradigm has revolutionized the field of natural language processing. Within the context of Pre-trained Language Models (PLMs), numerous approaches have been proposed to tackle text generation problems (Noorbakhsh et al., 2021). Autoregressive LMs (ALMs, e.g., GPT-1 (Radford et al., 2018) and T5 (Raffel et al., [n. d.])) and Non-Autoregressive LMs (NALMs, e.g., BERT (Kenton and Toutanova, 2019) and Roberta (Liu et al., 2019)) are the two primary strategies for mathematics (Figure 3).
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.4em,font=,, where level=2text width=8.4em,font=,, where level=3text width=8.0em,font=,, where level=4text width=12em,font=,, [ Language Models, ver [ Pre-trained
Language
Models  (§3) [ Autoregression
LMs  (§3.1) [ GPT-$f$ (Polu and Sutskever, 2020a), EPT (Kim et al., 2020), Generate &\&& Rank (Shen et al., 2021), Thor (Jiang et al., 2022), HTPS (Lample et al., [n. d.]), Galactica (Taylor et al., 2022), MATH-PLM (Hendrycks et al., 2021b), LISA (Jiang et al., 2021),
PACT (Han et al., 2022), Minerva (Lewkowycz et al., 2022), LIME (Wu et al., [n. d.]) , leaf, text width=54.2em ] ] [ Non-Autoregression
LMs  (§3.2) [ Aristo (Clark et al., 2021), GenBERT (Geva et al., 2020), NF-NSM (Feng et al., 2022), MWP-BERT (Liang et al., 2022), MathBERT(Peng et al., 2021), TAGOP (Zhu et al., 2021), MT2Net (Zhao et al., 2022),
DeductReasoner (Jie et al., 2022), BERT-TD (Li et al., 2022) , leaf, text width=54.2em ] ] ] [ Large
Language
Models  (§4) [ Instruction
Learning  (§4.1) [ Instruction
Building [ Auto-explanation (NAKAMOTO et al., 2023), ProofNet (Azerbayev et al., 2023), WizardLM (Xu et al., 2023), Wizardmath (Luo et al., 2023), ControlMath (Chen et al., 2024c),
EURUS (Yuan et al., 2024), RefAug (Zhang et al., 2024a) , leaf, text width=44.6em ] ] [ Instruct
Tuning [ MathGLM (Yang et al., 2023), Goat (Liu and Low, 2023), Calculon (Muffo et al., [n. d.]), PaLM 2-L-Math (Liu et al., 2023b), LLEMMA (Azerbayev et al., [n. d.]), Qwen-Math (Yang et al., 2024),
REFT (Luong et al., 2024), InternLM-Math (Ying et al., 2024) , leaf, text width=44.6em ] ] [ In-context
Learning [ ScratchpadGPT (Nye et al., 2021), Codex-math (Drori et al., 2022), In-Context Sampling (Liu et al., 2022), ComplexPrompt (Fu et al., 2023),
ART (Zhang et al., 2023e), LAMBADA (Kazemi et al., 2022) , LogicSolver (Yang et al., 2022) , leaf, text width=44.6em ] ] ] [ Tool-based
Methods  (§4.2) [ Single-tool
Methods [ SymbLLM (He-Yueya et al., 2023), PoT (Chen et al., 2022a), Codex-math (Drori et al., 2022), MathPrompter (Imani et al., 2023), PAL (Gao et al., 2023), Code-scratchpad (Upadhyay et al., 2023),
MuMath-Code (Yin et al., 2024), LeanReasoner (Jiang et al., 2024), LINC (Olausson et al., 2023), , leaf, text width=44.6em ] ] [ Multi-tool
Methods [ Toolformer (Schick et al., 2023), Chameleon (Lu et al., 2023b), ART (Paranjape et al., 2023), ToolkenGPT (Hao et al., 2023b), CRITIC (Gou et al., 2023a), TALM (Parisi et al., 2022),
Tool-Documentation (Hsieh et al., 2023) , leaf, text width=44.6em ] ] ] [ Fundamental CoT
Methods  (§4.3) [ Foundation of
CoT [ Answer Rationales (Ling et al., 2017), MathQA (Amini et al., 2019), CoT (Wei et al., 2022), Complexity-based CoT (Fu et al., 2023), RFT (Yuan et al., 2023a), AFT (Wang et al., 2023a),
Thought Propagation (Yu et al., 2023a), PoT (Chen et al., 2022a), AOT (Sel et al., 2023), MAmmoTH (Yue et al., 2023), SymbCoT (Xu et al., 2024a) , leaf, text width=44.6em ] ] [ Construction of
CoT [ Zero-shot CoT (Kojima et al., [n. d.]), Auto-CoT (Zhang et al., 2023e), Complexity-based CoT (Fu et al., 2023), PromptPG-CoT (Lu et al., 2023c),
AutoMate CoT (Shum et al., 2023), BoostedPrompt (Pitis et al., 2023), CoT-Influx (Huang et al., 2024) , leaf, text width=44.6em ] ] ] [ Advanced CoT
Methods  (§4.4) [ Verify-based
Methods [ Code-based self-verification (Zhou et al., 2023d), VerifyCoT (Ling et al., 2023), DIVERSE (Li et al., 2023b), Verify-and-Edit (Zhao et al., 2023a),
Retrieval-CoT (He et al., 2022), SCREWS (Shridhar et al., 2023), SSC-CoT (Zhao et al., 2024a) , leaf, text width=44.6em ] ] [ Ensemble-based
Methods [ Self-Consistency (Wang et al., 2023d), Diversity-of-Thought (Naik et al., 2023), Complexity-based CoT (Fu et al., 2023), DIVERSE (Li et al., 2023b),
Self-check (Miao et al., 2023), MCR (Yoran et al., 2023), Rank-verifier (Cobbe et al., 2021), GRACE (Khalifa et al., 2023) , leaf, text width=44.6em ] ] [ Planing-based
Methods [ ToT (Yao et al., 2023b), TouT (Mo and Xin, 2023), ToT-system (Long, 2023), GoT (Yao et al., 2023a; Besta et al., 2023; Lei et al., 2023), RESPROMPT (Jiang et al., 2023a), RAP (Hao et al., 2023a), LATS (Zhou et al., 2023e),
LLM+P (Liu et al., 2023a), LLM+DP (Dagan et al., 2023), Self-refine (Madaan et al., 2023; Sun et al., 2023), ISR-LLM (Zhou et al., 2023b), Reflexion (Shinn et al., 2023), CR (Zhang et al., 2023c)
Step-Plan (Zhang et al., 2023b) , leaf, text width=44.6em ] ] [ Self-correct
Methods [ o1 (OpenAI, 2024), QWQ (Team, 2024), STaR (Zelikman et al., 2022), V-STaR (Hosseini et al., 2024), Quiet-STaR (Zelikman et al., 2024), REFT (Trung et al., 2024), SCoRe (Kumar et al., 2024), HGS-PRM (Ma et al., 2023)
MCTSr (Zhang et al., 2024b), CoRe (Zhu et al., 2022), Marco-o1 (Zhao et al., 2024c), PRM (Lightman et al., 2024), Step-DPO (Lai et al., 2024), Flow-DPO (Deng and Mineiro, 2024), OmegaPRM (Luo et al., 2024)
, leaf, text width=44.6em ] ] [ Socratic Teaching
Methods [ Socratic Models (Zeng et al., 2022), SOCRATIC QUESTIONING (Qi et al., 2023), Socratic prompt (Chang, 2023), SocratiQ (Al-Hossami et al., 2023b),
multi-turn Socratic advice (Ang et al., 2023), SocraticLLM (Ding et al., 2024), MATHDIAL (Macina et al., 2023) , leaf, text width=44.6em ] ] ] [ Multi-modal
Methods  (§4.5) [ Multi-modal
Reasoning [ GPT-4V (OpenAI, 2023), Qwen2-VL (Wang et al., 2024a), Gemini (Team et al., 2023), GPT-4o (Hurst et al., 2024), GLM-4V (GLM et al., 2024), LLaVA-o1 (Xu et al., 2024b), AtomThink (Xiang et al., 2024),
Math-LLaVA (Shi et al., 2024), M-STAR (Anonymous, 2024), UnAC (wang and Fu, 2024),
, leaf, text width=44.6em ] ] ] ] ]
Figure 3. Taxonomy of language models for mathematic.
3.1. Autoregression LMs
The distinctions between Autoregression LMs and Non-Autoregressive LMs in sequence generation tasks are depicted in Figure 4. Furthermore, Autoregression LMs can be categorized into two parts: causal decoder (e.g., GPT-1 (Radford et al., 2018) and GPT-2 (Radford et al., 2019)) and encoder-decoder (e.g., T5 (Raffel et al., [n. d.])). The generation mechanisms of these two structures for sequences are depicted on the left and in the middle of Figure 4, respectively. These two architectures have not only significantly propelled the advancement of PLMs but have also become the predominant foundational architectures for LLMs.
Many improved Autoregression LMs-based methods have been proposed to enhance the performance of PLMs in mathematical problems. GPT-$f$ (Polu and Sutskever, 2020a) presents an automated prover and proof assistant to explore the performance of LMs to automated theorem proving. To resolve challenges related to expression fragmentation and operand-context separation, Kim et al. (Kim et al., 2020) introduce the Expression-Pointer Transformer (EPT), a pure neural model that employs both an ‘Expression’ token and operand-context pointers during the generation of solution equations. Generate &\&& Rank (Shen et al., 2021) introduces a novel ranking task for MWPs, a multi-task framework built upon a generative PLM, effectively addressing the challenge of minor mistakes in mathematical expressions. Thor (Jiang et al., 2022) introduces a framework that integrates language models and automated TP. The latter is employed to selectively choose relevant premises from an extensive library to assist the language model in TP. HTPS (Lample et al., [n. d.]) presents HyperTree Proof Search, a novel algorithm leveraging online training from past proof searches, enhancing model performance and surpassing the previous SOTA GPT-$f$. Galactica (Taylor et al., 2022) proposes a working memory token approach that can achieve strong performance over existing methods on mathematical MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b) benchmarks.
Furthermore, MATH-PLM (Hendrycks et al., 2021b) introduces a new dataset named MATH, consisting of challenging competition mathematics problems. It is found that relying solely on increasing budgets and model parameter counts would be impractical for achieving robust mathematical reasoning. LISA (Jiang et al., 2021) gathers an extensive collection of lemmas and theorems extracted from the Isabelle standard library and the Archive of Formal Proofs (AFP). Using this vast corpus, it builds LMs that prove theorems effectively within the AFP. PACT (Han et al., 2022) extracts abundant self-supervised data from kernel-level proof terms, enabling joint training alongside the standard tactic prediction objective. This effectively enhances the model’s ability for TP. Minerva (Lewkowycz et al., 2022) trains a language model, initially pre-trained on general natural language data, through additional training on technical content to address Quantitative Reasoning Problems. It achieves a SOTA performance on technical benchmarks including MATH (Hendrycks et al., 2021b). LIME (Wu et al., [n. d.]) introduces a novel pre-training methodology that specifically learns inductive bias for mathematical reasoning. Observations indicate that models trained using LIME exhibit superior performance compared to standard transformers.
Figure 4. The distinctions between Autoregression LMs and Non-Autoregression LMs.
3.2. Non-Autoregression LMs
Unlike Autoregression LMs, Non-Autoregression LMs enable the model to generate all parts of a sequence simultaneously, without depending on previously generated content, as depicted on the right side of Figure 4. Notably, models like BERT (Kenton and Toutanova, 2019), Roberta (Liu et al., 2019), and related architectures utilize masked word representations, leveraging pre-trained context-aware embeddings as comprehensive semantic features. This significantly elevates the performance benchmarks of NLP tasks.
In the field of mathematics, researchers have proposed various methods for designing models to address mathematical reasoning and computational problems. For example, Aristo (Clark et al., 2021) fine-tunes BERT using scientific curriculum data, yielding promising results on science exams. GenBERT (Geva et al., 2020) and NF-NSM (Feng et al., 2022) enhance the numerical reasoning capabilities of models by incorporating numerical data into the training process of PLMs. MWP-BERT (Liang et al., 2022) further enhances the model’s capacity to represent and calculate numerical values by incorporating numeric attributes into symbol placeholders. MathBERT(Peng et al., 2021) employs additional joint training of text and formulas to effectively capture the semantic-level structural information of formulas. TAGOP (Zhu et al., 2021), MT2Net (Zhao et al., 2022), and DeductReasoner (Jie et al., 2022) utilize BERT or RoBERTa to extract the fundamental arithmetic relationships between quantities, enabling mathematical reasoning and operations. BERT-TD (Li et al., 2022) utilizes semantic encoding and contrastive learning to cluster problems with similar prototype equations, thereby enhancing the understanding of MWP patterns.
4. LLMs-based Methods
Large-scale language models (LLMs) are designed for processing and generating text akin to human communication (Touvron et al., 2023a; OpenAI, 2023). Mathematics is also a form of language, that communicates complex concepts and relationships through a structured system of symbols and notation, akin to the rules of a spoken language. Thus, a language model that grasps these mathematical rules can ”speak” the language of mathematics, proving to be a valuable asset for mathematicians. In numerous ways, a sophisticated language model like GPT-4 (OpenAI, 2023) becomes an invaluable tool in the field of mathematics (Azerbayev et al., [n. d.]; Luo et al., 2023). We classify the existing studies into four parts: instruction learning, tool-based methods, fundamental CoT methods, and advanced CoT methods (Figure 3).
4.1. Instruction Learning
Numerous approaches have been proposed to enhance the mathematical performance of models by instruction learning. These approaches are categorically delineated as instruction building, instruction tuning, and in-context learning, based on their distinctive characteristics.
Instruction Building.
Instruction Building refers to the process of designing and generating high-quality instruction data that guide models in improving specific capabilities, such as mathematical reasoning. After pre-training, these instructions are crucial for directing models towards more sophisticated understanding and application of mathematical concepts. A semi-supervised approach, as presented in Auto-explanation (NAKAMOTO et al., 2023), leverages LLMs to create datasets for automating the scoring of mathematical self-explanations in the context of mathematics education. Similarly, RefAug (Zhang et al., 2024a) leverages GPT-4o to construct reflective augmentation data for training model, thereby enhancing their performance on mathematical reasoning tasks. A challenging benchmark ProofNet (Azerbayev et al., 2023) is introduced to build a system of automated theorem proving. Meanwhile, prompt retrieval and distilled back translation methods are introduced for statement autoformalization. WizardLM (Xu et al., 2023) presents a groundbreaking method called Evol-Instruct, designed to autonomously generate high-quality instructions by LLMs themselves. Furthermore, Wizardmath (Luo et al., 2023) proposes a reinforced Evol-instruct Feedback method (RLEIF) to build more complexity of instruct dataset. This method effectively enhances the mathematical reasoning capabilities of Llama-2 by supervised fine-tuning and –Proximal Policy Optimization (PPO) training (Schulman et al., 2017). ControlMath (Chen et al., 2024c) introduces a novel approach that first generates diverse equations, which are then transformed into math word problems by a Problem Crafter LLM-agent. EURUS (Yuan et al., 2024) generates a large number of reasoning trajectories by leveraging CoT (Wei et al., 2022) and tool creation (Qian et al., 2023) methods. These trajectories within multi-turn interactions (Wang et al., 2023c) are organized into a preference tree structure. Then extensive instruction and preference data are collected from this tree to enhance the model’s mathematical reasoning capabilities by Direct Preference Optimization (DPO) (Rafailov et al., 2024), Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024), and Noise Contrastive Alignment (NCA) (Chen et al., 2024a).
Instruction Tuning.
Instruction tuning stands as a potent method to elevate the prowess of large models, aiding in steering the model towards generating outputs that align with human intent (Zhang et al., 2023a). Compared to the pre-training phase of LLMs, instruction tuning demands notably less data and computational resources, making it a prevalent strategy for enhancing a model’s domain-specific capabilities. For instance, MathGLM (Yang et al., 2023) showcases that even an LLM with 2B parameters can adeptly execute multi-digit arithmetic operations with limited training data. Goat (Liu and Low, 2023) discerns patterns in the tokenization of different numbers and proposes a consistent approach for their tokenization based on the LLaMA model. Similarly, Calculon (Muffo et al., [n. d.]) refines the model’s arithmetic skills by employing a digit decomposition technique to construct a fine-tuning dataset. In PaLM 2-L-Math (Liu et al., 2023b), three fine-tuning strategies were explored, revealing key insights: 1) The quality and style of step-by-step solutions significantly influence model performance; 2) Combining solution re-ranking and majority voting yields a more substantial performance enhancement compared to using them individually. Additionally, multi-task fine-tuning, separating solution generation and evaluation tasks, surpasses the baseline achieved by solution fine-tuning alone. Moving beyond the enhancement of mathematical capabilities through fine-tuning, LLEMMA (Azerbayev et al., [n. d.]) introduces the Proof-Pile-2 dataset, an amalgamation of mathematical texts and code. Continual pre-training with Code Llama empowers the model to leverage Python interpreters and formal theorem provers, demonstrating impressive performance on the MATH benchmark. To improve the effectiveness of instruction tuning, Luong et al. (Luong et al., 2024) introduce an approach that incorporates online reinforcement learning. This method enhances the model’s generalization by exploring multiple CoTs and using real-world answers as reward signals. Recently, more math-specific LLMs (Yang et al., 2024; Ying et al., 2024) have been proposed, with enhanced instruction tuning tailored for the mathematical domain. Qwen2.5-MATH (Yang et al., 2024) begins by pretraining on large-scale, high-quality mathematical data. In the post-training phase, it develops a reward model (RM), which is then utilized to the iterative evolution of data in supervised fine-tuning (SFT). Thereafter, both the SFT model and the RM undergo iterative training. During the inference stage, the RM guides the sampling process to optimize the model’s performance. In contrast, InternLM-Math (Ying et al., 2024) undergoes secondary pretraining on a base model using a mathematical corpus. In the post-training phase, it is fine-tuned with instructions through unifying CoT and code interpretation, reward modeling, data augmentation under an unified seq2seq format.
In-context Learning.
In-context Learning (ICL) (Brown et al., 2020; Kaplan et al., 2020) empowers LLMs to execute target tasks by presenting specific task examples as conditions during inference, without updating model parameters. Inspired by this, ScratchpadGPT (Nye et al., 2021) enhances its proficiency in multi-step computations by mandating the model to output intermediate computation steps into a designated “scratchpad.” Codex-math (Drori et al., 2022) refines Codex by generating programs through code and integrates few-shot learning to automatically create programs for solving MWPs. This method has substantially improved the previous state-of-the-art accuracy in automatic solutions by 73.1% across various benchmarks like MIT’s mathematics courses, Columbia University’s Computational Linear Algebra, and the MATH benchmark (Hendrycks et al., 2021b). Notably, few-shot learning contributed to a 10% accuracy boost. Recent studies also highlight the variability in in-context learning performance across different chosen examples (Liu et al., 2022). Fu et al. (Fu et al., 2023) and Zhang et al. (Zhang et al., 2023e) selected intricate and diverse examples to enhance reasoning performance. LAMBADA (Kazemi et al., 2022) introduces a backward reasoning method for mathematical inference, starting from conclusions and working back to axioms. It uses few-shot to guide LLMs in reverse problem-solving, achieving promising results. LogicSolver (Yang et al., 2022) first constructed a dataset of mathematical word problems, InterMWP, where each problem includes interpretable logical formulas based on algebraic knowledge. The relevant logical formulas in InterMWP is used as In-context to enhance both interpretability and accuracy.
4.2. Tool-based Methods
LLMs are designed to use tools, such as codes and calculators, to enhance their problem-solving abilities (Schick et al., 2023; Parisi et al., 2022).
Single-tool Methods.
To improve the performance of mathematical reasoning, math-specific tools such as symbolic solvers and programs are utilized for LLMs (He-Yueya et al., 2023; Chen et al., 2022a). For example, SymbLLM (He-Yueya et al., 2023) solved MWPs by combining LMs with symbolic solvers. They focus on automatically generating high-quality, step-by-step solutions to mathematical word problems, especially those encountered in mathematical applications. LeanReasoner (Jiang et al., 2024) and Lean Copilot (Song et al., 2024) integrates LLMs with the Lean theorem proving framework. Leveraging Lean’s symbolic solver reduces the risk of logical inconsistencies and enhances the LLMs’ capability to handle complex reasoning tasks through the extensive theorem libraries available in Lean. Similarly, LINC (Olausson et al., 2023) integrates LLMs with a first-order logic theorem prover. The LLM translates natural language premises and conclusions into First-Order Logic expressions. These expressions are then sent to an external theorem prover, such as Prover9, for symbolic deductive reasoning to determine the truth value of the conclusion. Furthermore, previous studies have explored the process of converting MWPs into code (Chen et al., 2022a; Drori et al., 2022). PoT (Chen et al., 2022a) proposes a fusion of CoT with programs, while Drori et al. (Drori et al., 2022) showcased the effectiveness of pre-training on text and fine-tuning on code. This approach successfully solves, explains and generates math problems at the university level. MathPrompter (Imani et al., 2023) employs a zero-shot CoT prompting technique to generate multiple algebraic expressions or Python functions in varied ways to solve the same math problem. PAL (Gao et al., 2023) introduces an innovative approach to bolster the performance of LLMs in mathematical problem-solving. This involves utilizing LLMs to comprehend natural language problems and generate programs as intermediate reasoning steps. To enhance the ability of LLMs to solve MWPs using code, MuMath-Code (Yin et al., 2024) constructed an augmented dataset that uses multi-perspective data augmentation methods and synthesizes code-nested solutions. In addition to solving math problems, programs can also play a role in tutoring math. For instance, Upadhyay explores the role of LLMs in tutoring systems (Upadhyay et al., 2023), introducing a “code scratchpad” alongside the traditional “language scratchpad” to enhance the model’s performance in tutoring steps, particularly using a grade school mathematics dataset.
Multi-tool Methods.
To facilitate the seamless integration of LLMs with diverse tools, several multi-tool approaches have been proposed to enable LLMs to learn how to use multiple tools simultaneously. Toolformer (Schick et al., 2023) adopts a self-supervised training approach to enable the utilization of different tools such as search engines, calculators and translation systems via simple API calls. This is accomplished by fine-tuning on a vast collection of sampled API calls and filtering based on their ability to reduce perplexity on subsequent tokens. Chameleon (Lu et al., 2023b), on the other hand, enhances LLMs with plug-and-play modules designed for compositional reasoning. It creates programs by combining various tools, including LLMs, off-the-shelf vision models, web search engines, Python functions, and heuristic-based modules, to tackle complex reasoning tasks. The Automatic Reasoning and Tool-use (ART) framework (Paranjape et al., 2023) leverages frozen LLMs to automatically generate intermediate reasoning steps as a program. This seamlessly incorporates external tools to support computations that surpass the core capabilities of LLMs. Meanwhile, ToolkenGPT (Hao et al., 2023b) adopts a strategy of learning “toolken” embeddings to represent each tool as a token. This empowers LLMs to effortlessly utilize tools similar to generating word tokens. It caters to a broader range of tools and utilizes extensive demonstration data to learn toolkit embeddings. Furthermore, CRITIC (Gou et al., 2023a) significantly enhances the outputs of LLMs by enabling them to verify and self-correct through interactions with external tools. Inspired by human cognition and critical thinking, CRITIC continuously refines text generated by LLMs through iterative interactions with tools like search engines and code interpreters. Tool Augmented Language Models (TALM) (Parisi et al., 2022) seamlessly integrate language models with non-differentiable tools. They employ an iterative “self-play” method to bootstrap performance based on a limited number of tool demonstrations. In contrast, Tool-Documentation (Hsieh et al., 2023) opted for tools based on documentation rather than relying on demonstrations. Current practices involve teaching LLMs through a few-shot demonstration of tool usage, a process prone to bias and overfitting. The alternative proposed by Tool-Documentation, utilizing tool documentation that provides descriptions of individual tool usage, is argued to be a more effective approach. ToRA (Gou et al., 2023b) employs a set of Tool-integrated Reasoning Agents, which integrate computational libraries and symbolic solvers, among other mathematical tools, to effectively solve complex MWPs.
4.3. Fundamental CoT Methods
A multitude of fundamental Chain-of-Thought (CoT) methods integrated with LLMs have been proposed to enhance the mathematical reasoning abilities of LLMs.
Foundation of CoT
In the initial stages of research, a limited body of work has leveraged the principles of CoT to enhance the mathematical capabilities of language models. Notably, Ling et al (Ling et al., 2017) propose a methodology involving the generation of a series of concise steps, termed ”Answer Rationales,” to guide the resolution of algebraic word problems. MathQA (Amini et al., 2019) suggests decomposing Math Word Problems into multiple steps corresponding to programmatic operations for resolution. Utilizing the in-context learning of LLMs, CoT (Wei et al., 2022) explicitly introduces the concept of the CoT for the first time, and substantiates its efficacy in enhancing reasoning abilities. This validation is also corroborated in Complexity-based CoT (Fu et al., 2023). Consequently, several methods (Chen et al., 2022a; Yuan et al., 2023a; Yu et al., 2023a; Wang et al., 2023a; Sel et al., 2023; Yue et al., 2023) leverage CoT methodologies to enhance the mathematical performance of models. RFT (Yuan et al., 2023a) proposes the application of rejection sampling finetuning (RFT) to gather more optimal reasoning paths, thereby enhancing mathematical reasoning performance. Thought Propagation (Yu et al., 2023a) considers gaining insights from analogous problems to assist in addressing the current issue. Wang et al. (Wang et al., 2023a) proposed a Fine-Tuning (AFT) paradigm that aligns the model to prioritize the generation of CoT with superior performance. SymbCoT (Xu et al., 2024a) integrates symbolic expressions and logical rules with CoT prompting, enhancing the logical rigor of CoT. Additionally, PoT (Chen et al., 2022a) introduces a Program of Thoughts, combining CoT and programming, and AOT (Sel et al., 2023) enhances reasoning abilities through algorithmic-style demonstrations of CoT. Furthermore, MAmmoTH (Yue et al., 2023) integrates CoT and PoT (Chen et al., 2022a) rationales to instruct large language models in utilizing code tools for solving mathematical problems.
Construction of CoT
To streamline the process of CoT creation, various approaches (Lu et al., 2023c; Zhang et al., 2023e; Fu et al., 2023; Wang et al., 2023a; Kojima et al., [n. d.]; Shum et al., 2023; Pitis et al., 2023; Sel et al., 2023; Yu et al., 2023a) have been introduced for its automatic generation. Zero-shot CoT (Kojima et al., [n. d.]) introduces a simple CoT example, “Let’s think step by step,” which effectively enhances the model’s reasoning capabilities. Auto-CoT (Zhang et al., 2023e) selects representative questions through clustering and answers them using “Let’s think step by step,” concatenating the question-answer pairs one by one to automatically generate CoT. In the case of Complexity-based CoT (Fu et al., 2023), the demonstration is chosen by simply selecting the reasoning chain with the highest number of inference steps sampled from the model. Several methods aim to enhance the quality of CoT generated by models. CoT-Influx (Huang et al., 2024) trains a pruner specifically optimized for enhancing the mathematical capabilities of CoT generation, and validates the effectiveness of the pruner in improving the mathematical performance of LLMs on the GSM8K (Cobbe et al., 2021). Using reinforcement learning, PromptPG-CoT (Lu et al., 2023c) trains a model with policy gradient to assist GPT-3 in selecting suitable CoT demonstrations. Similarly, AutoMate CoT (Shum et al., 2023) employs a variance-reduced policy gradient strategy to estimate the significance of each example in a black box language model, thereby selecting more effective examples. BoostedPrompt (Pitis et al., 2023) presents an iterative prompt ensemble method that enhances prompts when the current demonstration faces challenges in handling specific problems.
4.4. Advanced CoT Methods
To further enhance the capability of CoT in LLMs, advanced CoT methods have been proposed, including Verify-based Methods, Ensemble-based Methods, Planning-based Methods, and Socratic Teaching Methods.
Verify-based Methods.
LLMs often produce incorrect reasoning steps, which can lead to a series of cascading errors in their solutions. Implementing verification and refining the reasoning process based on feedback can significantly reduce these errors. This approach is akin to human reflection and involves critically evaluating each step of the reasoning process. Various verify-based methods (Zhou et al., 2023d; Ling et al., 2023; Li et al., 2023b; Zhao et al., 2023a; He et al., 2022; Shridhar et al., 2023) have been proposed to address these issues. Zhou et al. (Zhou et al., 2023d) propose a code-based self-verification approach, utilizing a zero-shot prompt to encourage the GPT-4 Code Interpreter to employ code for self-verifying its responses, enhancing its mathematical reasoning capabilities. To mitigate the challenge of validating the entire deductive reasoning process, VerifyCoT (Ling et al., 2023) introduces a deductive reasoning form, ensuring that each reasoning step strictly relies on the preceding steps. Furthermore, DIVERSE (Li et al., 2023b) independently verifies each reasoning step and a voting mechanism to eliminate incorrect answers. Both Verify-and-Edit (Zhao et al., 2023a) and Retrieval-CoT (He et al., 2022) utilizes external retrieval tools to support the model in validating reasoning rationales. The key difference is the former edits rationales using retrieved information, while the latter helps the model self-rethink to improve performance on complex reasoning tasks. Both methods effectively reduce factual mistakes during the reasoning process. Shridhar et al. (Shridhar et al., 2023) summarize the use of the revisions approach in the SCREWS framework, which includes a selection module to choose between the original and modified reasoning steps. SSC-CoT (Zhao et al., 2024a) found that prompting the model from critical intermediate steps can enhance the success rate of solving mathematical problems. Therefore, it emphasizes verifying these key steps during the inference process.
Ensemble-based Methods.
Due to the inherent stochastic nature of LLMs, which output probability distributions over predicted words, they may randomly generate incorrect reasoning steps and outcomes. To tackle this challenge, some methods (Wang et al., 2023d; Naik et al., 2023; Fu et al., 2023; Li et al., 2023b; Miao et al., 2023; Yoran et al., 2023; Cobbe et al., 2021; Khalifa et al., 2023) leverage the concept of ensemble learning, employing techniques such as voting and ranking to eliminate uncertainties in the reasoning process. Self-Consistency (Wang et al., 2023d) employs multiple reasoning paths and selects the final response through a simple majority vote. Similarly, Diversity-of-Thought (Naik et al., 2023) generates diverse reasoning paths by altering prompts and aggregates the responses via the majority vote. Complexity-based CoT(Fu et al., 2023) favors answers derived from more intricate reasoning paths. DIVERSE (Li et al., 2023b) uses a weighted voting mechanism to filter out incorrect answers. Nevertheless, these voting-based methods often overlook the potentially useful information within unsuccessful CoT reasoning and lack an efficient integration of multiple reasoning chains to improve performance. Self-check (Miao et al., 2023) addresses this by incorporating reasoning steps into the voting mechanism, ensuring both consistent answers and reliable reasoning. MCR (Yoran et al., 2023) takes a step further by consolidating information across various reasoning chains, acquiring the most relevant facts, facilitating a more comprehensive analysis to make successful reasoning. Additionally, there are a few methods based on ranking. Rank-verifier (Cobbe et al., 2021) proposes using a ranking system to judge the correctness of model completions. Furthermore, GRACE (Khalifa et al., 2023) leverages a discriminator trained via contrastive learning to rank each reasoning step.
Planing-based Methods.
The original structure of the Chain of Thought (CoT) is sequential, facing limitations when handling highly complex problems and lacking the ability for retrospective correction. To make the CoT chain structure more systematic and intricate, certain planning-based methods (Yao et al., 2023b; Mo and Xin, 2023; Long, 2023; Yao et al., 2023a; Besta et al., 2023; Lei et al., 2023; Jiang et al., 2023a; Hao et al., 2023a; Zhou et al., 2023e; Liu et al., 2023a; Dagan et al., 2023; Madaan et al., 2023; Sun et al., 2023; Zhou et al., 2023b; Shinn et al., 2023) have been proposed. These methods achieve this by altering the organizational structure of reasoning steps or incorporating mechanisms for refinement and reflection. The Tree-of-Thought (ToT) (Yao et al., 2023b) proposes organizing reasoning paths into a tree structure, where each node represents a reasoning step, and edges denote dependencies between nodes. During the inference process, ToT can use self-assessment to determine future actions. This structure facilitates both forward and backward exploration, employing either deep-first or breadth-first search techniques. TouT (Mo and Xin, 2023) effectively employs Monte Carlo Dropout to assess uncertainty scores associated with diverse local responses of language models at intermediate steps, enhancing response precision through integration with global search algorithms. Long et al. (Long, 2023) introduced a multi-module ToT system, where the state of the ToT thought chain process can be stored and retraced through its storage module. Furthermore, Yao et al. (Yao et al., 2023a), Besta et al. (Besta et al., 2023), and Lei et al. (Lei et al., 2023) introduced a novel thought structure called the Graph of Thought (GoT). Similar to ToT, each graph node represents a distinct thought step or state, with edges indicating dependencies between nodes. These methods can effectively integrate diverse reasoning thoughts, fostering collaborative outcomes and improving reasoning by incorporating feedback loops. This contributes to an overall enhancement in the inference capacity of the thought network. CR (Zhang et al., 2023c) proposes a method where LLMs assume three roles—Proposer, Verifier, and Reporter—to collaboratively solve problems. This approach differs from CoT and ToT in its ability to dynamically store and utilize all historically validated inference results for execution and combination, a concept akin to GoT. Additionally, RESPROMPT (Jiang et al., 2023a) introduces Residual Connection Prompting, which converts the input prompt into complex reasoning graphs. Inspired by Monte Carlo tree search (MCTS) (Chaslot et al., 2008), RAP (Hao et al., 2023a) utilizes the MCTS to traverse tree structures during the reasoning process. This approach effectively balances global and local searches to obtain a high-reward reasoning path. On the other hand, LATS (Zhou et al., 2023e) employs LLMs as agents, value functions, and optimizers, repurposing their inherent strengths to enhance decision-making. Step-Plan (Zhang et al., 2023b) propose a step-by-step planning approach for intermediate solution generation, which strategically plans the generation of the next solution step. Liu et al. (Liu et al., 2023a) and Dagan et al. (Dagan et al., 2023) introduced LLM+P and LLM+DP, incorporating the generation of Planning Domain Definition Language (PDDL) to break down complex problems and execute plans using specialized models. Furthermore, Self-Refine, as proposed by Madaan et al. (Madaan et al., 2023) and Sun et al. (Sun et al., 2023), focuses on error correction and summarizing past experiences. Specifically, Self-Refine performs pre-execution validation, providing feedback to the model for refinement in case of errors. Zhou et al. (Zhou et al., 2023b) then took a step further by merging the iterative self-refine with PDDL in ISR-LLM. Additionally, Shinn et al. (Shinn et al., 2023) introduced re-flexion to rectify errors stemming from previous actions.
Self-correct Methods.
Recently, a lot of Self-correct Methods have been proposed, which introduce a self-correction mechanism on top of CoT. Combined with techniques like MCTS and RL, these methods enable the model to iteratively enhance itself. The success of OpenAI’s o1 (OpenAI, 2024) marks a significant advancement in high-level reasoning by employing long CoT and self-correction mechanisms. Building on this, QwQ (Team, 2024), Marco-o1 (Zhao et al., 2024c) and o3 have also been introduced, further validating the effectiveness of long CoT and self-correction. STaR (Zelikman et al., 2022) leverages existing LLM reasoning to iteratively create a high-quality dataset, refining the model through self-generated inferences. Similar to STaR, REFT (Trung et al., 2024) does not employ SFT, and instead refines the model using the PPO in an online reinforcement learning setting. V-STaR (Hosseini et al., 2024) builds upon STaR by retaining error trajectories during the iterative generation process. Quiet-STaR (Zelikman et al., 2024) enables LLMs to generate a “rationale” before predicting the next token, explaining the anticipated content of the future text. CoRe (Zhu et al., 2022) mimics the human brain’s dual-system model, separating reasoning into generation and verification, and it uses MCTS for feedback during inference. Additionally, SCoRe (Kumar et al., 2024) introduces a two-stage reinforcement strategy to mitigate distribution shifts and behavior collapse, facilitating self-correction. MCTSr (Zhang et al., 2024b) proposes a method based on Monte Carlo Tree Self-Refine that enhances LLM’s performance on the Mathematical Olympiad, matching GPT-4. Flow-DPO (Deng and Mineiro, 2024) introduces a method where two LLM agents collaborate to generate step-level preference data, which is used to train the LLM using the DPO algorithm. Lightman et al. (Lightman et al., 2024) presents a Process Reward Model (PRM) and compares it with the Outcome Reward Model. Step-DPO (Lai et al., 2024) also introduces a method in which each step of the reasoning process is optimized based on preferences, rather than evaluating the entire answer as a whole. OmegaPRM (Luo et al., 2024) proposes an automated process supervision method with a divide-and-conquer style MCTS algorithm, which significantly improves Gemini’s proficiency in mathematical reasoning. Moreover, HGS-PRM (Ma et al., 2023) proposes combining Heuristic Greedy Search with PRM to optimize the reasoning paths.
Socratic Teaching Methods.
The Socratic teaching method (Nelson, 1980), derived from the philosophies of ancient Greek thinker Socrates, provides a unique perspective when considering the integration of LMs in mathematics education (Davies et al., 2021). This method, emphasizing questioning and dialogue, aims to guide learners toward profound thinking (Davis and Steinglass, 1997; Brickhouse and Smith, 2009; Ding et al., 2024). In realm of mathematics education, employing this method can aid students in grasping concepts with clarity, fostering more rigorous logical application, and promoting reflection and self-discovery. For LLMs, research shows that the Socratic teaching method can also improve their abilities to a certain extent. For example, the framework of Socratic Models (SMs) (Zeng et al., 2022) significantly broadens the scope of interactive capabilities. SMs facilitate the amalgamation of multiple LLMs to tackle novel multi-modal tasks using structured Socratic dialogues. In Chang et al. (Chang, 2023), techniques including definition, dialectics, cross-examination, induction, and counterfactual reasoning are utilized. These techniques contribute to enhancing the LLMs’ comprehension of user intentions, thereby improving model’s effectiveness in addressing user queries. The inductive, deductive, and analogical reasoning techniques in the Socratic teaching method establish a comprehensive analytical framework for solving mathematical problems (Chang, 2023; Al-Hossami et al., 2023a), ensuring a strong and concise ability to address the complexities of mathematical problem-solving. MATHDIAL (Macina et al., 2023) enables LLMs to engage in multi-round dialogues as both teacher and student, creating a conversational dataset that incorporates rich educational socratic attributes and background in mathematical reasoning problems. Moreover, contributions from Ang et al. (Ang et al., 2023) and Al et al. (Al-Hossami et al., 2023b) provided invaluable datasets and evaluation methodologies for Socratic questioning. These contributions underscore the practical applications and potential impact of Socratic methods in counseling and coaching, showcasing their diverse applicability. Overall, the integration of Socratic teaching techniques in LLMs offers novel perspectives and tools for mathematics education. Unlike CoT (Wei et al., 2022), Socratic questioning explicitly guides the thinking process, stimulating effective recursive thinking and displaying greater resilience to errors in the thinking process (Qi et al., 2023). Ranging from basic conversational guidance to intricate problem-solving, these methods demonstrate the effective fusion of modern technology with classical ideas to bolster the learning and comprehension abilities of LLMs.
4.5. Multi-modal Methods
Building on LMs, some multimodal approaches incorporate visual encoders, enabling the models to handle mathematics problems that rely on visual elements, such as geometry, topology, tabular statistics, etc. GPT-4V (OpenAI, 2023) is the first successfully proposed Multimodal LLMs (MLLMs) that demonstrate strong visual reasoning capabilities to handle certain image-based mathematical problems. Subsequently, a lot of other high-performing MLLMs have been introduced, including Gemini (Team et al., 2023), Qwen2-VL (Wang et al., 2024a), CogVLM (Wang et al., 2023b), GPT-4o (Hurst et al., 2024), GLM-4V (GLM et al., 2024) and InternVL (Chen et al., 2024d), etc. UnAC (wang and Fu, 2024) introduces a multimodal prompting approach that decomposes the reasoning process into three stages: understanding the image, abstracting image information, and systematically verifying the reasoning steps. Math-LLaVA (Shi et al., 2024) introduces a high-quality, diverse multimodal mathematical dataset called MathV360K, which significantly improves the model’s multimodal mathematical reasoning capabilities. LLaVA-o1 (Xu et al., 2024b) and AtomThink (Xiang et al., 2024) introduce long CoT into the domain of MLLMs. M-STAR (Anonymous, 2024) investigates the application of self-evolving training in multimodal reasoning by combining continuous self-evolving training, reward model re-ranking, and adaptive exploration strategies.
Table 1. The statistics information of mathematical datasets. Level: EE = Elementary, MM = Middle School, HH = High School, UU = University, CC = Competition, HH = Hybrid, OO = Others. Modality: TT = Text, MM = Multimodal
[…]
VERBPHYSICS (Forbes and Choi, 2017) 733 1,096 1,828 3,657 EN Calculation Training Formula - TT
Clinical (Spithourakis et al., 2016; Spithourakis and Riedel, 2018) 11,170 1,625 3,220 16,015 EN Calculation Training Formula OO TT
Scientific (Spithourakis and Riedel, 2018) 14,694 2,037 4,231 20,962 EN Calculation Training Formula OO TT
DoQ (Elazar et al., 2019) 587 5,418 6,007 12,012 EN Calculation Training Formula - TT
DROP (Dua et al., 2019) 77,409 9,536 9,622 96,567 EN Calculation Training Text EE TT
AddSub (Hosseini et al., 2014) - - - 395 EN MWP Training Formula EE TT
SingleOp (Roy et al., 2015) 265 107 159 531 EN MWP Training Formula EE TT
SingleEq (Koncel-Kedziorski et al., 2015) - - - 508 EN MWP Training Formula EE TT
MultiArith (Roy and Roth, 2015) 420 - 180 600 EN MWP Training Formula - TT
Alg514 (Kushman et al., 2014) - - - 514 EN MWP Training Formula - TT
Math23k (Wang et al., 2017) 22,162 - 1000 23,162 CH MWP Training Formula EE TT
AQuA (Ling et al., 2017) 97,467 254 254 97,975 EN MWP Training Text UU TT
GSM8K (Cobbe et al., 2021) 7,473 1,319 8,792 EN MWP Training Text EE TT
SVAMP (Patel et al., 2021) 700 300 1,000 EN MWP Training Formula EE TT
DRAW (Upadhyay and Chang, 2015) - - - 1,000 EN MWP Training Formula OO TT
Dolphin1878 (Shi et al., 2015) - 374 1,504 1,878 EN MWP Training Formula - TT
HMWP (Qin et al., 2020) - - - 5,470 CH MWP Training Formula - TT
ArMATH (Alghamdi et al., 2022) - - - 6,000 Arabic MWP Training Formula EE TT
TabMWP (Lu et al., 2023c) - - - 38,431 EN MWP Training Text EEMM TT
TAL-SCQ5K  [^1] https://github.com/math-eval/TAL-SCQ5K 3,000 - 2,000 5,000 CH/EN MWP Training Text CC TT
REALFP (Kalyan et al., 2021) 185 185 558 928 EN MWP Training Formula CC TT
SYNTHFP (Kalyan et al., 2021) 10,000 1,000 1,000 12,000 EN MWP Training Formula OO TT
MultiHiertt (Zhao et al., 2022) 7,830 1,044 1,566 10,440 EN MWP Training Formula OO MM
MATHPILE (Wang et al., 2023e) - - - 903,180 EN MWP Training Text HH TT
OpenWebMath (Paster et al., 2023) - - - - EN MWP Training Formula HH TT
MathQA (Amini et al., 2019) 29,837 4,475 28,985 37,297 EN MQA Training Formula - TT
FinQA (Chen et al., 2021a) 6,251 883 1,147 8,281 EN MQA Training Formula OO TT
TAT-QA (Zhu et al., 2021) - - - 16,552 EN MQA Training Text OO MM
MML (Grabowski et al., 2015) - - - 57,882 EN TP Training Formula - TT
HolStep (Kaliszyk et al., 2017) 2,013,046 - 196,030 2,209,076 EN TP Training Formula - TT
CoqGym (Yang and Deng, 2019) - - - 71,000 EN TP Training Formula - TT
HOList (Bansal et al., 2019) - - - 29,462 EN TP Training Formula - TT
IsarStep (Li et al., 2021) 820,000 5,000 5,000 830,000 EN TP Training Formula UU TT
LISA (Jiang et al., 2021) - - - 183,000 EN TP Training Formula - TT
NaturalProofs (Welleck et al., 2021) 32,000 EN TP Training Text OO TT
LeanStep (Han et al., 2022) - - - 21,606,000 EN TP Training Formula - TT
NumGLUE (Mishra et al., 2022b) - - - 101,835 EN Calculation Benchmark Text HH TT
Dophin18k (Huang et al., 2016) - - - 18,460 EN MWP Benchmark Text EE TT
MAWPS (Koncel-Kedziorski et al., 2016) - - - 3,320 EN MWP Benchmark Formula HH TT
ASDiv (Miao et al., 2020) - - - 2,305 EN MWP Benchmark Formula EE TT
MATH (Hendrycks et al., 2021b) 7,500 5,000 12,500 EN MWP Benchmark Text HHOO TT
MGSM (Shi et al., 2022) - - - - Multilingual MWP Benchmark Text EE TT
Mathematics (Saxton et al., 2019) 2,000,000 100,000 2,100,000 EN MWP Benchmark Formula - TT
MMLU-Math (Hendrycks et al., 2021a) - - - 906 EN WMP Benchmark Formula EEHHCC TT
AGIEval (Zhong et al., 2023) - - - 469/220 CH/EN MWP Benchmark Formula HHCCOO TT
AIME [^2] https://www.kaggle.com/datasets/hemishveeraboina/aime-problem-set-1983-2024 - - - 933 EN MWP Benchmark Formula CC TT
MATHTRAP (Zhao et al., 2024b) - - - 105/155 EN MWP Benchmark Formula HH TT
MathVista (Lu et al., 2023a) - - - 6,141 EN MWP Benchmark Formula EEMMHHUU MM
Math-V (Wang et al., 2024b) - - - 3,040 EN MWP Benchmark Formula EEHHOO MM
CMM-Math (Liu et al., 2024) 22,248 - 5,821 28,069 CH MWP Benchmark Text EEMMHH MM
MathVerse (Zhang et al., 2025) - - - 2,612 EN MWP Benchmark Formula HH MM
GeoQA (Chen et al., 2021b) 3,499 745 754 4,998 EN MWP Benchmark Formula MM MM
M3CoT (Chen et al., 2024b) - - - 11,459 EN MWP Benchmark Formula OO MM
MR-MATH (Xia et al., 2024) - - - - EN MWP Benchmark Formula EE TT
MMMU-Math (Yue et al., 2024) - - - 505 EN MWP Benchmark Text UU MM
We-Math (Qiao et al., 2024) - - - 1700 EN MWP Benchmark Formula EEHHUU MM
U-MATH (Chernyshev et al., 2024) - - - 1,080 EN MWP Benchmark Formula UU MM
INT (Wu et al., 2021) - - - - EN TP Benchmark Formula - TT
miniF2F (Zheng et al., 2023) - 244 244 488 EN TP Benchmark Formula EEHHOO TT
FOLIO (Han et al., 2024) - - - 1,430 EN TP Benchmark Formula HH TT
Aggregate (Roy and Roth, 2018) - - - 1,492 EN MWP Augmented Formula HH TT
MathQA-Python (Austin et al., 2021) 19,209 2,822 2,822 23,914 EN MWP Augmented Code - TT
Math50k (Li et al., 2023a) - - - 50,000 EN WMP Augmented Text - TT
PRM800K (Lightman et al., 2024) - - - 2,868 EN WMP Augmented Text HHOO TT
MetaMathQA (Yu et al., 2023b) - - - 395,000 EN MWP Augmented Text HH TT
Lila (Mishra et al., 2022a) - - - 134,000 EN MWP Augmented Code HH TT
PEN (Kim et al., 2022) - - - 3,581 EN MWP Augmented Formula HH TT
miniF2F+informal (Jiang et al., 2023b) - 244 244 488 EN TP Augmented Formula EEHHOO TT
NaturalProofs-Gen (Welleck et al., 2022) 12,500 1,000 1,000 14,500 EN TP Augmented Text OO TT
[…]
5. Datasets
To train and evaluate the arithmetic and math ability of language models, various math word problems (MWPs) datasets (Saxton et al., 2019; Hendrycks et al., 2021b; Cobbe et al., 2021; Kushman et al., 2014) are designed. In this paper, we organize the frequently used datasets for mathematical language models by dividing them into training, benchmark and data augmentation datasets (Table 1).
5.1. Training Datasets
5.1.1. Mathematical Calculation
Several studies have introduced datasets aiming to identify numerical information within text (Spithourakis et al., 2016; Spithourakis and Riedel, 2018; Elazar et al., 2019), specifically targeting the prediction of attribute values. The Clinical Data (Spithourakis et al., 2016; Spithourakis and Riedel, 2018) consists of clinical records sourced from the London Chest Hospital. Each patient record includes a text report and structured KB tuples detailing 20 potential numeric attributes like age and gender. Scientific Data (Spithourakis and Riedel, 2018) encompasses paragraphs extracted from Cornell’s ARXIV repository, featuring over half a million converted papers across 37 scientific sub-fields. Additionally, Elazar et al. (Elazar et al., 2019) proposed the Distributions over Quantities (DoQ) dataset, comprising empirical counts of scalar attribute values linked to more than 350K nouns, adjectives, and verbs across 10 different attributes. Furthermore, certain studies consider the relational knowledge between various numbers. For instance, VERBPHYSICS (Forbes and Choi, 2017) aggregates crowdsourced data on actions and objects, including knowledge of grounded object pairs and actions related to them. More intricately, DROP (Dua et al., 2019) dataset necessitates Discrete Reasoning Over the content of Paragraphs. In this benchmark of 55,000 adversarial questions, a system resolves references within a question across multiple positions and performs operations like addition, counting, or sorting.
5.1.2. Math Word Problems
A large number of datasets are proposed for MWPs. Hosseini et al. (Hosseini et al., 2014) curated AddSub, primarily focusing on addition and subtraction problems. This dataset identifies relevant variables and their respective values to translate sentences into problem statements, represented as equations. SingleOp (Roy et al., 2015), encompasses a wider range of mathematical operations including multiplication and division. Its purpose is to facilitate reasoning about quantities articulated in natural language. DRAW (Upadhyay and Chang, 2015) comprises 1000 algebra word problems semi-automatically annotated for evaluating automatic solvers. It features gold coefficient alignments crucial for uniquely identifying equation system derivations. The Alg514 (Kushman et al., 2014) dataset comprises 514 linear algebra problems structured around 28 equation templates. Meanwhile, MultiArith (Roy and Roth, 2015) is designed to handle arithmetic problems involving multiple steps and operations, without predefined templates. MATHPILE (Wang et al., 2023e) is a diverse and high-quality math-centric corpus comprising about 9.5 billion tokens. OpenWebMath (Paster et al., 2023) is an open dataset inspired by these works containing 14.7B tokens of mathematical webpages from Common Crawl. Several studies have introduced methods focused on predicting equations through semantic parsing. SingleEq (Koncel-Kedziorski et al., 2015) revolves around grade-school algebra word problems. It comprises samples that can be correlated to a single equation involving multiple mathematical operations on one variable, structured as a parsing tree. In contrast, Dolphin1878 (Shi et al., 2015) contains 1,878 word problems in mathematics, wherein each sample might encompass multiple equations and variables depicted through multiple trees. AllArith (Roy and Roth, 2017) offers a concise representation of the dependencies among number units mentioned in a given problem, referred to as Unit Dependency Graphs (UDGs). To augment dataset size, Math23k (Wang et al., 2017) gathers 23,161 problems labeled with structured equations and corresponding answers. GSM8K (Cobbe et al., 2021) comprises 8.5K high-quality grade school math problems crafted by human problem writers, where the problems typically involve 2 to 8 steps for resolution. HMWP (Qin et al., 2020) aims to enhance dataset diversity by extracting MWPs from a Chinese K12 problem bank. This effort sought to validate the universality of math word problem solvers and expand research in MWPs to better align with real-world scenarios. In total, it comprises 5,491 MWPs, including 2,955 single-variable linear MWPs, 1,636 two-variable linear MWPs, and 900 single-variable nonlinear MWPs. Additionally, MathQA (Amini et al., 2019) presents a diverse collection of 37,000 English multiple-choice MWPs spanning various mathematical domains. Existing research endeavors to enhance the comprehensibility and precision of intricate reasoning by providing fine-grained annotations of reasoning processes (Upadhyay and Chang, 2017; Ling et al., 2017; Chen et al., 2021a; Lightman et al., 2024; Zhao et al., 2022). DRAW-1K (Upadhyay and Chang, 2017) introduces a novel dataset featuring 1,000 general algebra word problems, each annotated with derivations. AQUA (Ling et al., 2017) structures each question into four components: problem description, multiple-choice answer options, rationale, and correct option label. FinQA (Chen et al., 2021a) presents an expert-annotated dataset comprising 8,281 financial QA pairs, meticulously annotated with numerical reasoning processes to ensure comprehensive explication. PRM800K (Lightman et al., 2024) assigns labels (positive, negative, or neutral) to each step in solving MATH problems sampled by a large-scale generator. The training set encompasses 800K step-level labels across 75K solutions to 12K problems. Additionally, Patel et al. (Patel et al., 2021) introduced the SVAMP challenge set to establish a more robust evaluation framework for methods designed to tackle elementary-level MWPs. Alghamdi et al. (Alghamdi et al., 2022) contributed the inaugural large-scale dataset, ArMATH, for Arabic MWPs. This dataset comprises 6,000 samples of primary-school math problems written in Modern Standard Arabic. Kalyan et al. (Kalyan et al., 2021) proposed two datasets, REALFP and SYNTHFP, for a novel task called Fermi Problems (FPs). This task necessitates the amalgamation of various reasoning abilities, including suitable abstractions, commonsense knowledge, and creative synthesis. Recently, mathematical reasoning has incorporated multiple input modalities such as textual, visual, and tabular data (Lu et al., 2023c; Zhu et al., 2021; Zhao et al., 2022). For instance, TabMWP (Lu et al., 2023c) presents a new dataset comprising 38,431 open-domain grade-level problems that necessitate mathematical reasoning across both textual and tabular data. Zhu et al. (Zhu et al., 2021) introduced a large-scale QA dataset, TAT-QA, encompassing both tabular and textual data. This dataset often requires numerical reasoning—such as addition, subtraction, multiplication, division, counting, comparison/sorting, and their combinations—based on real financial reports to infer answers. Additionally, MultiHiertt (Zhao et al., 2022) is a proposed dataset consisting of QA pairs derived from Multi Hierarchical Tabular and Textual data, where each document in this dataset contains multiple tables and longer unstructured texts.
5.1.3. Theorem Proving
Current research delves into exploring mathematical theorem datasets (e.g., INT (Wu et al., 2021), Feit-Thompson (Huang et al., 2019), and IsarStep (Li et al., 2021)) to develop novel machine learning-based theorem-proving strategies. Huang et al. (Huang et al., 2019) introduced the Feit-Thompson dataset, encompassing 1,602 lemmas, expanding into 83,478 proof states for the Feit-Thompson theorem (Gonthier et al., 2013). IsarStep (Li et al., 2021) constitutes a non-synthetic dataset extracted from the largest repository of proofs handwritten by human experts in a theorem prover. Furthermore, NaturalProofs (Welleck et al., 2021) comprises 32,000 theorem statements and proofs, 14,000 definitions, and 2,000 other types of pages (e.g., axioms, corollaries). This dataset draws from three domains: broad-coverage data sourced from ProofWiki, deep-coverage data from the Stacks, and low-resource data extracted from mathematics textbooks. Several formal mathematical libraries (Megill and Wheeler, 2019; Rudnicki, 1992; Grabowski et al., 2015) encompass a range of formal languages tailored for theorem proving (TP). Examples include MML (Grabowski et al., 2015), Coq (Bertot and Castéran, 2013), Lean (Han et al., 2022), Isabelle (Wenzel et al., 2008), and Hol light (Slind and Norrish, 2008). For instance, MML (Grabowski et al., 2015), also known as the Mizar Mathematical Library, was established to systematically construct a centralized, reusable knowledge base reflecting the standard foundations of mathematics, namely classical first-order logic and set theory. Yang et al. (Yang and Deng, 2019) compiled the CoqGym dataset, comprising 71,000 human-written proofs from 123 projects developed using the Coq proof assistant. LeanStep (Han et al., 2022) serves as the tactic proof dataset for the Lean theorem prover, featuring high-level human-written tactics alongside kernel-level proof terms. LISA (Jiang et al., 2021), one of the largest proof corpora for interactive theorem provers, contains 183,000 theorems and 2.16 million proof steps. This dataset is extracted from the Archive of Formal Proofs and the standard library of Isabelle. Kaliszyk et al. (Kaliszyk et al., 2017) introduced the HolStep dataset rooted in Higher-Order Logic (HOL) proofs. HOList (Bansal et al., 2019) includes almost 30,000 theorems and proofs across three corpora: core, complex, and flyspeck. The core corpus contains fundamental theorems necessary for defining tactics, while the complex corpus consists of theorems related to complex calculus. Flyspeck contains the majority of the lemmas and theorems related to the Kepler conjecture (Hales et al., 2017).
5.2. Benchmark Datasets
Various benchmark datasets have been introduced to assess the performance of different mathematical algorithms, such as MAWPS (Koncel-Kedziorski et al., 2016). Huang et al. (Huang et al., 2016) developed Dolphin18K, a diverse collection containing over 18,000 annotated MWPs. Mathematics (Saxton et al., 2019) encompasses problems spanning arithmetic, algebra, probability, and calculus. These problems involve sequential questions and answers presented in a free-form textual input/output format. For assessing arithmetic understanding, Mishra et al. (Mishra et al., 2022b) introduced NumGLUE, a multi-task benchmark that evaluates AI systems across eight distinct tasks. miniF2F (Zheng et al., 2023) is a dataset designed to create a unified cross-system benchmark for neural theorem proving. It comprises 488 formal Olympiad-level mathematics problem statements, encompassing Metamath, Lean, Isabelle, and HOL Light. FOLIO (Han et al., 2024) is a human-annotated, logically complex and diverse dataset for reasoning in natural language (NL), equipped with 1,430 first-order logic (FOL) annotations. Several datasets consider varying difficulty levels across educational stages, ranging from elementary to high school (Miao et al., 2020; Shi et al., 2022; Hendrycks et al., 2021b). Miao et al. (Miao et al., 2020) introduced ASDiv, comprising 2,305 MWPs categorized by problem type and elementary school grade level. The Multilingual Grade School Math Benchmark (MGSM) (Shi et al., 2022) leverages problems from GSM8K, translating them into 10 languages with the assistance of human annotators. The MATH dataset (Hendrycks et al., 2021b) contains 12,500 problems sourced from high school math competitions, where each problem within MATH is equipped with a step-by-step solution. Furthermore, several datasets serve to evaluate foundation models (Zhong et al., 2023; Yuan et al., 2023b). AGIEval (Zhong et al., 2023) functions as a human-centric benchmark tailored to assess foundation models’ general abilities in tasks related to human cognition and problem-solving. Yuan et al. (Yuan et al., 2023b) introduced the arithmetic dataset MATH 401, designed to test the latest large language models. With the development of multimodal LLMs (MLLMs), benchmarks for multimodal mathematical reasoning are proposed (Yue et al., 2024; Qiao et al., 2024; Chernyshev et al., 2024; Chen et al., 2024b; Zhang et al., 2025; Wang et al., 2024b). For example, MathVista (Lu et al., 2023a) is a popular English mathematical benchmark for MLLMs. Liu et al. (Liu et al., 2024) released a Chinese multimodal mathematical reasoning dataset to enhance and measure the performance of Multimodal LLMs. Moreover, MMMU-Math (Yue et al., 2024) and U-Math (Chernyshev et al., 2024) collect university-level problems to further evaluate the performance.
5.3. Augmented Datasets
Augmented datasets serve to enhance existing datasets by incorporating additional samples or information. Roy et al. (Roy and Roth, 2018) introduced Aggregate by augmenting the AllArith dataset with 661 word problems from Perturb. Perturb contains problems that were manually pruned either for not yielding the desired solution $a$ or being too dissimilar from the input problem $p$. Large-scale language models like ChatGPT and GPT-4 have been utilized to expand datasets. For instance, MetaMathQA (Yu et al., 2023b) rephrases questions from multiple perspectives without introducing extra knowledge to bootstrap mathematical questions. The Math5K dataset (Li et al., 2023a) comprises 50,000 problem-solution pairs generated using GPT-4. These datasets leverage advanced language models to enrich the existing collection of problems and solutions. Several studies have expanded existing datasets by including supplementary information, such as code programs (Austin et al., 2021; Mishra et al., 2022a) and explanations (Kim et al., 2022). MathQA-Python (Austin et al., 2021) is a Python adaptation of the MathQA benchmark, comprising 23,914 problems. This dataset aims to evaluate models’ capabilities in synthesizing code from intricate textual descriptions. Lila (Mishra et al., 2022a) is a comprehensive mathematical reasoning benchmark composed of 23 diverse tasks. It extends 20 existing datasets by gathering task instructions and solutions as Python programs, offering not only correct answers but also explainable solutions. PEN (Kim et al., 2022) focuses on providing plausible and accurate explanations for solving algebraic word problems present in three benchmark datasets: ALG514, DRAW-1K, and MAWPS. Moreover, some studies have contributed statements and proofs based on existing datasets for theorem proving (Jiang et al., 2023b; Welleck et al., 2022). For example, miniF2F+informal (Jiang et al., 2023b) is a dataset consisting of manually curated informal statements and proofs aligned with formal statements from the miniF2F dataset. NaturalProofs-Gen (Welleck et al., 2022) adapts data from NATURALPROOFS, including theorem statements, proofs, definitions, and additional pages sourced from ProofWiki. Each proof follows a multi-step structure and references a variable number of sources.
Table 2. Comparison of typical mathematical LMs. Pre-T: Pre-Training, FT: Finetuning, MM: Multimodal, Symbolic-S: Symbolic solvers, F-CoT: Foundational CoT, A-CoT: Advanced CoT, RL: Reinforce Learning
[…]
GPT-$f$ (Polu and Sutskever, 2020a) - Pre-T Symbolic-S
LISA (Jiang et al., 2021) - Pre-T Symbolic-S
MATH-PLM (Hendrycks et al., 2021b) GPT-2, GPT-3 Pre-T F-CoT
Minerva (Lewkowycz et al., 2022) PaLM Pre-T
MWP-BERT (Liang et al., 2022) BERT Pre-T
ControlMath (Chen et al., 2024c) LLaMA-2, Mistral FT F-CoT
EURUS (Yuan et al., 2024) Mistral, CodeLLaMA FT F-CoT DPO, KTO, NCA
MathGLM (Yang et al., 2023) GLM FT
WizardLM (Xu et al., 2023) LLaMA FT
Wizardmath (Luo et al., 2023) LLaMA-2 FT PPO
PaLM 2-L-Math (Liu et al., 2023b) PaLM-2-L FT F-CoT
Codex-math (Drori et al., 2022) Codex Prompt Program
Toolformer (Schick et al., 2023) GPT-J FT Calculator
LeanReasoner (Jiang et al., 2024) CodeLLaMA FT Symbolic-S
LINC (Olausson et al., 2023) StarCoder+ Prompt Symbolic-S F-CoT
MuMath-Code (Yin et al., 2024) Llama-2 FT Program
MAmmoTH (Yue et al., 2023) - FT Program F-CoT
InternLM-Math (Ying et al., 2024) InternLM2 Pre-T, FT Program F-CoT PPO
Qwen-Math (Yang et al., 2024) - Pre-T, FT Program F-CoT PPO
PAL (Gao et al., 2023) - Prompt Program F-CoT
CoT (Wei et al., 2022) - Prompt F-CoT
PromptPG-CoT (Lu et al., 2023c) - Prompt
BoostedPrompt (Pitis et al., 2023) - Prompt F-CoT
ToT (Yao et al., 2023b) - Prompt A-CoT
PoT (Chen et al., 2022a) - Prompt Program F-CoT
Self-check (Miao et al., 2023) - Prompt A-CoT
Self-Consistency (Wang et al., 2023d) - Prompt A-CoT
Diversity-of-Thought (Naik et al., 2023) - Prompt A-CoT
RAP (Hao et al., 2023a) - Prompt A-CoT
LATS (Zhou et al., 2023e) - Prompt A-CoT
LLM+P (Liu et al., 2023a) - Prompt Symbolic-S A-CoT
LLM+DP (Dagan et al., 2023) - Prompt Symbolic-S A-CoT
ISR-LLM (Zhou et al., 2023b) - Prompt A-CoT
MCR (Yoran et al., 2023) - Prompt A-CoT
Rank-verifier (Cobbe et al., 2021) - Prompt A-CoT
GRACE (Khalifa et al., 2023) - Prompt A-CoT
Reflexion (Shinn et al., 2023) - Prompt A-CoT
CR (Zhang et al., 2023c) - Prompt A-CoT
Step-Plan (Zhang et al., 2023b) - Prompt A-CoT
STaR (Zelikman et al., 2022) GPT-J FT A-CoT
V-STaR (Hosseini et al., 2024) LLaMA2, CodeLLaMA FT A-CoT DPO
Quiet-STaR (Zelikman et al., 2024) Mistral Pre-T A-CoT
REFT (Luong et al., 2024) Galactica, CodeLLaMA FT A-CoT PPO
SCoRe (Kumar et al., 2024) Gemini FT A-CoT SCoRe
HGS-PRM (Ma et al., 2023) LLaMA2, WizardMath FT A-CoT PPO
MCTSr (Zhang et al., 2024b) LLaMa-3 FT A-CoT
CoRe (Zhu et al., 2022) GPT-J, DeBERTa-large FT A-CoT
Marco-o1 (Zhao et al., 2024c) Qwen2 FT A-CoT
Step-DPO (Lai et al., 2024) Qwen2 FT A-CoT DPO
Flow-DPO (Deng and Mineiro, 2024) LLaMA3, Phi-3 FT A-CoT DPO
OmegaPRM (Luo et al., 2024) Gemini Pro, Gemma2 FT A-CoT PRM
SocraticLLM (Ding et al., 2024) Qwen2 FT A-CoT
MATHDIAL (Macina et al., 2023) T5 FT A-CoT
QwQ (Team, 2024) Qwen2.5 Pre-T Long CoT
o1 (OpenAI, 2024) - Pre-T Long CoT
GPT-4o (Hurst et al., 2024) - Pre-T A-CoT
GPT-4V (OpenAI, 2023) - Pre-T A-CoT
Qwen2-VL (Wang et al., 2024a) Qwen2-VL Pre-T F-CoT
LLaVA-o1 (Xu et al., 2024b) LLaVA FT Long CoT
Gemini (Team et al., 2023) Gemini Pre-T A-CoT
GLM-4V (GLM et al., 2024) ChatGLM Pre-T A-CoT
AtomThink (Xiang et al., 2024) LLaVA, EMOVA FT Long CoT PRM
Math-LLaVA (Shi et al., 2024) LLaVA FT F-CoT
M-STAR (Anonymous, 2024) MiniCPM-V FT A-CoT PRM
UnAC (wang and Fu, 2024) GPT-4V, Gemini, LLaVA Prompt F-CoT
[…]
Table 3. Performance of various mathematical LMs in terms of accuracy.
[…]
MATH-PLM (Hendrycks et al., 2021b) 6.9 - - - - - - - -
Minerva (Lewkowycz et al., 2022) 50.3 78.5 75.0 - - - - - -
MWP-BERT (Liang et al., 2022) - - - - - 76.6 - - -
ControlMath (Chen et al., 2024c) 9.5 49.3 - - - - 51.6 - -
EURUS (Yuan et al., 2024) 41.7 62.8 - 93.0 - - 90.4 - -
Wizardmath (Luo et al., 2023) 22.7 81.6 - - - - - -
Codex-math (Drori et al., 2022) 81.8 - - - - - - - -
Toolformer (Schick et al., 2023) - - - 40.4 - - 29.4 - -
LeanReasoner (Jiang et al., 2024) - - - - - - - 82.6 -
LINC (Olausson et al., 2023) - - - - - - - 72.5 -
MuMath-Code (Yin et al., 2024) 55.1 90.7 - - - - - - -
MAmmoTH (Yue et al., 2023) 44.2 76.7 - - - - - - 61.4
InternLM-Math (Ying et al., 2024) 37.7 82.6 - - - - - - -
Qwen-Math (Yang et al., 2024) 66.8 91.6 82.8 - 63.3 86.3 - - -
PAL (Gao et al., 2023) - 72.0 - 79.6 - - 79.4 - -
BoostedPrompt (Pitis et al., 2023) - 87.1 - - - - - - 63.8
PoT (Chen et al., 2022a) - 80.0 - - - 89.1 - - 58.6
Self-check (Miao et al., 2023) 51.3 88.1 - - 81.2 - - - -
Self-Consistency (Wang et al., 2023d) - 78.0 - 87.8 - - 86.8 - 52.0
Diversity-of-Thought (Naik et al., 2023) - 96.3 - - - - - - 81.7
GRACE (Khalifa et al., 2023) - 36.3 - - - 84.4 68.6 - -
CR (Zhang et al., 2023c) 72.2 - - - - - - 98.0 -
REFT (Luong et al., 2024) - 75.3 - - - 71.8 79.2 - -
SCoRe (Kumar et al., 2024) 64.4 - - - - - - - -
HGS-PRM (Ma et al., 2023) 13.7 65.4 - - - - - - -
MCTSr (Zhang et al., 2024b) 58.2 96.6 - - 11.8 - - - -
CoRe (Zhu et al., 2022) - 63.2 - 90.5 - - - - -
Step-DPO (Lai et al., 2024) 70.8 94.0 - - - - - - -
Flow-DPO (Deng and Mineiro, 2024) 38.6 71.3 - - - - - - -
OmegapRM (Luo et al., 2024) 69.4 93.6 - - - - - - -
QwQ (Team, 2024) 90.6 - - - 50.0 - - - -
GLM-4-9B-Chat (GLM et al., 2024) 50.6 79.6 72.4 - - - - - -
[…]
o1 (OpenAI, 2024) 94.8 - 92.3 - 83.3 78.1 73.9 - -
GPT-4o (Hurst et al., 2024) 60.3 - 88.0 - 13.4 69.2 63.8 29.02 30.4
GPT-4V (OpenAI, 2023) - - - - - 56.8 49.9 - 22.8
Qwen2-VL (Wang et al., 2024a) - - 46.2 - - 64.5 70.5 43.0 25.9
LLaVA-o1 (Xu et al., 2024b) 64.0 - - - - - 54.8 - -
Gemini (Team et al., 2023) 67.7 - 85.9 - - 62.2 63.9 41.9 17.7
GLM-4V (GLM et al., 2024) - - - - - 47.2 - - -
AtomThink (Xiang et al., 2024) - - - - - - 53.3 - 40.5
Math-LLaVA (Shi et al., 2024) - - - - 38.3 38.3 46.6 - 15.7
M-STAR (Anonymous, 2024) - - - - - - 59.5 - -
UnAC (wang and Fu, 2024) - - - - - 59.7 56.6 - -
[…]
6. Analysis and Discussion
To gain a deeper understanding of the advantages of typical mathematical LMs, we conduct a comprehensive comparison of their characteristics and performance (Table 2 and Table 3). Table 2 provides a detailed overview of mathematical LLMs, their architectures, training strategies, reasoning capabilities, and other advanced features. From the table, we observe the following findings. First, CoT has become a critical enabler for mathematical reasoning. Additionally, with the introduction of OpenAI’s o1 model, an increasing number of studies are focusing on Long CoT. Second, only a few models (e.g., LLaVA, M-STAR) support multimodal input (text + image). This indicates that most mathematical LLMs focus solely on text-based reasoning. As tasks increasingly require interpreting diagrams and visual data, multimodal models could play a more prominent role in future developments. Third, models like Toolformer and LeanReasoner integrate external tools (e.g., calculators or symbolic solvers). These tools complement the LLMs’ reasoning by handling computations or symbolic operations that the models alone might struggle with. Fourth, reinforcement learning (RL) is primarily used to refine reasoning strategies and align model outputs with user expectations or task requirements. However, only a small subset of models uses RL techniques, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024) or Proximal Policy Optimization (PPO) (Schulman et al., 2017), due to the instability of RL. Fifth, many recent models (e.g., CoT, BoostedPrompt, PoT) rely on prompt engineering to tackle complex problems without requiring heavy computational resources for fine-tuning. Fine-tuning is essential for most advanced mathematical models (e.g., WizardMath, EURUS) to enhance mathematical reasoning capabilities. Table 3 compares the performance of various mathematical LMs across different datasets. The methods can be classified into two categories: Text-based Methods and Multi-modal Methods. Such benchmarks are widely recognized and serve to assess mathematical capabilities in diverse contexts. Here are some conclusions based on the data: First, o1 stands out as the most accurate model across most datasets, achieving the highest scores in MATH (94.8%), MMLU (92.3%), AIME (83.3%), MMMU (78.1%) and MathVista (73.9%). It indicates that slow thinking with long CoT improves the performance of mathematical reasoning effectively. Second, methods like EURIUS and PoT show dataset-specific strengths, suggesting that tailored approaches can yield better results for certain problem types (e.g., formal proofs, complex reasoning). For example, CR achieves the best performance on the FOLIO Dataset with 98.0%, highlighting its capability for handling formal logic problems. Third, some Multi-modal benchmarks, such as CMM-Math (Liu et al., 2024) and Math-V (Wang et al., 2024b), show relatively low accuracy across all methods, indicating a need for further research and improvement in these areas.
7. Challenges and Further Directions
Faithfulness.
Mathematical LLMs reside in the phenomena of hallucination and faithfulness, where the generated output may lack factual accuracy or logical grounding (Rawte et al., 2023; Ji et al., 2023). This phenomenon leads to the production of erroneous or misleading mathematical results, undermining the reliability of the model’s outputs. Some studies explored to address this problem by integrating extra knowledge (He et al., 2022), reinforcement learning from human feedback (Lightman et al., 2024), tools (Schick et al., 2023; Parisi et al., 2022), and verify-based methods (Zhou et al., 2023d; Ling et al., 2023; Li et al., 2023b; Zhao et al., 2023a; He et al., 2022; Shridhar et al., 2023). However, the improvement of the hallucination phenomena is limited, which influences the trustworthiness and utility of mathematical language models in practical applications and scholarly pursuits.
Multi-modal.
In particular, math problems (e.g., geometry problems (Gelernter et al., 1960)) involve not only textual information but also various modalities such as diagrams, graphs, or mathematical symbols (Lu et al., 2021b). While existing models excel in processing textual information, interpreting and reasoning across multiple modalities concurrently remains a formidable task. Few multi-modal mathematical datasets and methods are proposed for this task (Zhao et al., 2022; Lu et al., 2021b). Bridging the gap between textual representations and visual/mathematical elements necessitates robust mechanisms that can effectively capture and synthesize information from disparate sources, ultimately leading to comprehensive and accurate problem-solving capabilities. In fact, the multi-modal information in mathematics is much more complex than general multi-modal tasks like vision question answering (Antol et al., 2015) and image captioning (Hossain et al., 2019). Achieving proficiency in handling multi-modal mathematical problems stands as a pivotal yet intricate objective in advancing the competency of mathematical LLMs.
Uncertainty.
The uncertainty of LLMs (Gawlikowski et al., 2023; Duan et al., 2023) leads to the ambiguity and variability problem of mathematical problems. While these models excel in deterministic tasks, their handling of uncertainty, such as probabilistic reasoning or dealing with incomplete or vague information, poses a significant challenge. Mathematical problems often entail nuanced interpretations, fuzzy constraints, or scenarios where a single precise solution might not exist. Several studies investigated this problem via controlled generation technologies (Zhou et al., 2023a). However, ensuring that LLMs can navigate and appropriately account for these uncertainties while providing accurate and contextually relevant solutions remains a complex task.
Evaluation.
It is still a challenge to evaluate mathematical LMs with robust and comprehensive evaluation metrics that adequately capture the models’ performance across various mathematical tasks. Traditional evaluation metrics in natural language processing might not effectively capture the intricacies of mathematical reasoning and problem-solving. Designing evaluation benchmarks (Saxton et al., 2019; Hendrycks et al., 2021b; Cobbe et al., 2021; Kushman et al., 2014) and metrics (Wang and Komatsuzaki, 2021; Chung et al., 2022; Thoppilan et al., 2022; Yuan et al., 2023b) that encompass a wide spectrum of mathematical tasks, spanning from basic arithmetic to complex theorem proving, while accounting for linguistic fluency and mathematical accuracy, remains a significant challenge. Addressing these challenges is crucial to ascertain the reliability, efficacy, and generalizability of mathematical LMs, fostering advancements in this burgeoning field.
Creation.
While previous models exhibit remarkable capabilities in understanding and manipulating existing mathematical concepts, their ability to autonomously devise and rigorously prove entirely new theorems presents a formidable hurdle. The development of novel mathematical theorems demands not only profound mathematical reasoning but also creative and insightful problem-solving abilities, aspects that necessitate a deeper understanding of abstract mathematical concepts beyond what the models have been trained on. Davies et al. (Davies et al., 2021) applied machine learning techniques to discover potential patterns and relationships among mathematical entities. Recently, Bernardino et al. (Bernardino Romera-Paredes et al., 2023) proposed FunSearch to find the first discoveries made for established open problems using LLMs. Bridging the gap between the models’ proficiency in assimilating existing mathematical knowledge and their capability to generate novel and impactful theorems represents a significant frontier in leveraging LLMs for the advancement of mathematical discovery.
Application.
While mathematical LMs exhibit promising potential in autonomously solving math problems, their deployment in educational settings as teaching aids or tutors necessitates addressing several pivotal challenges. Tailoring these models to serve as effective educational tools demands not only mathematical proficiency but also adeptness in pedagogy and instructional methodologies. Few studies applied Socratic questioning for mathematical teaching (Zeng et al., 2022; Shridhar et al., 2022). Customizing LLMs to cater to diverse learning styles, adapting explanations to suit different proficiency levels, and fostering an interactive and engaging learning environment is interesting.
Data scarcity.
The training data significantly influences the performance of language models, particularly in LLMs (Ding et al., 2023). High-quality and diverse training data can assist the model in enhancing its mathematical reasoning capabilities (Luo et al., 2023). As discussed in Section 4.1, while there have been limited studies on constructing instruction data through LLMs, these efforts have only considered building from a small set of mathematical reasoning datasets, such as GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). High-quality and diverse mathematical instruction data remains scarce. It is necessary to explore additional forms and construction methods for mathematical train data. Additionally, the generation of mathematical training datasets in a multimodal context is also a promising direction.
8. Conclusions
The survey elucidates the pivotal role of mathematical LMs in reshaping the landscape of mathematical problem-solving, leveraging a spectrum of models, from PLMs to LLMs, to address diverse mathematical tasks. Our taxonomical delineation of mathematical tasks and methods provides a systematic framework for comprehending the intricacies of LMs-based methodologies, distinguishing between arithmetic calculation, mathematical reasoning, and various algorithmic approaches employed in these models. We also outline the features and effectiveness of current mathematical LMs to gain a deeper understanding of these techniques. We gather over 60 diverse mathematical datasets, carefully organized into training, benchmark, and augmented categories, highlighting how crucial data is in pushing forward mathematical research and supporting insightful inquiry in various mathematical fields. Additionally, by thoughtfully tackling challenges like faithfulness, multi-modality, uncertainty, evaluation, theorem creation, application, and data scarcity, this survey opens doors for future studies aimed at enhancing the capabilities of mathematical LMs. By exploring the current advancements, obstacles, and paths for future research, we see this thorough overview as a key resource for fostering innovation and guiding the future of mathematical LMs research, all while contributing to the exciting developments in mathematics and artificial intelligence.
Acknowledgements.
The authors wish to thank the reviewers for their helpful comments and suggestions. This research is funded by the National Science and Technology Major Project (No. 2021ZD0114002), the National Nature Science Foundation of China (No. 62477010, No.62307028), the Science and Technology Commission of Shanghai Municipality Grant (No. 22511105901, No. 21511100402), Shanghai Science and Technology Innovation Action Plan (No. 24YF2710100) and Shanghai Special Project to Promote High-quality Industrial Development (No. RZ-CYAI-01-24-0288).
  • (1)
  • Al-Hossami et al. (2023a) Erfan Al-Hossami, Razvan Bunescu, Justin Smith, and Ryan Teehan. 2023a. Can Language Models Employ the Socratic Method? Experiments with Code Debugging. arXiv:2310.03210 [cs.CL]
  • Al-Hossami et al. (2023b) Erfan Al-Hossami, Razvan Bunescu, Ryan Teehan, Laurel Powell, Khyati Mahajan, and Mohsen Dorodchi. 2023b. Socratic questioning of novice debuggers: A benchmark dataset and preliminary evaluations. In BEA. 709–726.
  • Alghamdi et al. (2022) Reem Alghamdi, Zhenwen Liang, and Xiangliang Zhang. 2022. ArMATH: a Dataset for Solving Arabic Math Word Problems. In LREC. 351–362.
  • Amini et al. (2019) Aida Amini, Saadia Gabriel, Shanchuan Lin, Rik Koncel-Kedziorski, Yejin Choi, and Hannaneh Hajishirzi. 2019. MathQA: Towards Interpretable Math Word Problem Solving with Operation-Based Formalisms. In NAACL. 2357–2367.
  • Ang et al. (2023) Beng Heng Ang, Sujatha Das Gollapalli, and See Kiong Ng. 2023. Socratic Question Generation: A Novel Dataset, Models, and Evaluation. In EACL. 147–165.
  • Anonymous (2024) Anonymous. 2024. Diving into Self-Evolve Training for Multimodal Reasoning. In Submitted to The Thirteenth International Conference on Learning Representations.
  • Antol et al. (2015) Stanislaw Antol, Aishwarya Agrawal, Jiasen Lu, Margaret Mitchell, Dhruv Batra, C Lawrence Zitnick, and Devi Parikh. 2015. Vqa: Visual question answering. In ICCV. 2425–2433.
  • Austin et al. (2021) Jacob Austin, Augustus Odena, Maxwell Nye, Maarten Bosma, Henryk Michalewski, David Dohan, Ellen Jiang, Carrie Cai, Michael Terry, Quoc Le, et al. 2021. Program synthesis with large language models. arXiv (2021).
  • Azerbayev et al. (2023) Zhangir Azerbayev, Bartosz Piotrowski, Hailey Schoelkopf, Edward W Ayers, Dragomir Radev, and Jeremy Avigad. 2023. Proofnet: Autoformalizing and formally proving undergraduate-level mathematics. arXiv (2023).
  • Azerbayev et al. ([n. d.]) Zhangir Azerbayev, Hailey Schoelkopf, Keiran Paster, Marco Dos Santos, Stephen McAleer, Albert Q Jiang, Jia Deng, Stella Biderman, and Sean Welleck. [n. d.]. LLEMMA: AN OPEN LANGUAGE MODEL FOR MATHEMATICS. Minerva 8 ([n. d.]), 164B.
  • Bansal et al. (2019) Kshitij Bansal, Sarah Loos, Markus Rabe, Christian Szegedy, and Stewart Wilcox. 2019. Holist: An environment for machine learning of higher order logic theorem proving. In ICML. 454–463.
  • Berg-Kirkpatrick and Spokoyny (2020) Taylor Berg-Kirkpatrick and Daniel Spokoyny. 2020. An empirical investigation of contextualized number prediction. In EMNLP. 4754–4764.
  • Bernardino Romera-Paredes et al. (2023) Alexander Novikov Bernardino Romera-Paredes, Mohammadamin Barekatain et al. 2023. Mathematical discoveries from program search with large language models. Nature (2023).
  • Bertot and Castéran (2013) Yves Bertot and Pierre Castéran. 2013. Interactive theorem proving and program development: Coq’Art: the calculus of inductive constructions.
  • Besta et al. (2023) Maciej Besta, Nils Blach, Ales Kubicek, Robert Gerstenberger, Lukas Gianinazzi, Joanna Gajda, Tomasz Lehmann, Michal Podstawski, Hubert Niewiadomski, Piotr Nyczyk, and Torsten Hoefler. 2023. Graph of Thoughts: Solving Elaborate Problems with Large Language Models.
  • Bobrow et al. (1964) Daniel Bobrow et al. 1964. Natural language input for a computer problem solving system. (1964).
  • Briars and Larkin (1984) Diane J Briars and Jill H Larkin. 1984. An integrated model of skill in solving elementary word problems. Cognition and instruction 1, 3 (1984), 245–296.
  • Brickhouse and Smith (2009) Thomas C Brickhouse and Nicholas D Smith. 2009. Socratic teaching and Socratic method. (2009).
  • Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. NeurIPS 33 (2020), 1877–1901.
  • Chang (2023) Edward Y Chang. 2023. Prompting large language models with the socratic method. In CCWC. 0351–0360.
  • Chaslot et al. (2008) Guillaume Chaslot, Sander Bakkes, Istvan Szita, and Pieter Spronck. 2008. Monte-carlo tree search: A new framework for game ai. In AAAI, Vol. 4. 216–217.
  • Chen et al. (2024a) Huayu Chen, Guande He, Lifan Yuan, Ganqu Cui, Hang Su, and Jun Zhu. 2024a. Noise contrastive alignment of language models with explicit rewards. arXiv (2024).
  • Chen et al. (2021b) Jiaqi Chen, Jianheng Tang, Jinghui Qin, Xiaodan Liang, Lingbo Liu, Eric P Xing, and Liang Lin. 2021b. GeoQA: A geometric question answering benchmark towards multimodal numerical reasoning. arXiv (2021).
  • Chen et al. (2024c) Nuo Chen, Ning Wu, Jianhui Chang, and Jia Li. 2024c. ControlMath: Controllable Data Generation Promotes Math Generalist Models. arXiv (2024).
  • Chen et al. (2024b) Qiguang Chen, Libo Qin, Jin Zhang, Zhi Chen, Xiao Xu, and Wanxiang Che. 2024b. M3CoT: A Novel Benchmark for Multi-Domain Multi-step Multi-modal Chain-of-Thought. In ACL. 8199–8221.
  • Chen et al. (2022a) Wenhu Chen, Xueguang Ma, Xinyi Wang, and William W Cohen. 2022a. Program of thoughts prompting: Disentangling computation from reasoning for numerical reasoning tasks. arXiv (2022).
  • Chen et al. (2022b) Yanda Chen, Ruiqi Zhong, Sheng Zha, George Karypis, and He He. 2022b. Meta-learning via Language Model In-context Tuning. In ACL. 719–730.
  • Chen et al. (2021a) Zhiyu Chen, Wenhu Chen, Charese Smiley, Sameena Shah, Iana Borova, Dylan Langdon, Reema Moussa, Matt Beane, Ting-Hao Huang, Bryan R Routledge, et al. 2021a. FinQA: A Dataset of Numerical Reasoning over Financial Data. In EMNLP. 3697–3711.
  • Chen et al. (2024d) Zhe Chen, Jiannan Wu, Wenhai Wang, Weijie Su, Guo Chen, Sen Xing, Muyan Zhong, Qinglong Zhang, Xizhou Zhu, Lewei Lu, et al. 2024d. Internvl: Scaling up vision foundation models and aligning for generic visual-linguistic tasks. In CVPR. 24185–24198.
  • Chernyshev et al. (2024) Konstantin Chernyshev, Vitaliy Polshkov, Ekaterina Artemova, Alex Myasnikov, Vlad Stepanov, Alexei Miasnikov, and Sergei Tilga. 2024. U-MATH: A University-Level Benchmark for Evaluating Mathematical Skills in LLMs. arXiv:2412.03205
  • Chowdhery et al. (2022) Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, et al. 2022. Palm: Scaling language modeling with pathways. arXiv (2022).
  • Chu et al. (2023) Zheng Chu, Jingchang Chen, Qianglong Chen, Weijiang Yu, Tao He, Haotian Wang, Weihua Peng, Ming Liu, Bing Qin, and Ting Liu. 2023. A Survey of Chain of Thought Reasoning: Advances, Frontiers and Future. arXiv:2309.15402 [cs.CL]
  • Chung et al. (2022) Hyung Won Chung, Le Hou, Shayne Longpre, Barret Zoph, Yi Tay, William Fedus, Yunxuan Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, et al. 2022. Scaling instruction-finetuned language models. arXiv (2022).
  • Clark et al. (2021) Peter Clark, Oren Etzioni, Daniel Khashabi, Tushar Khot, Bhavana Dalvi Mishra, Kyle Richardson, Ashish Sabharwal, Carissa Schoenick, Oyvind Tafjord, Niket Tandon, Sumithra Bhakthavatsalam, Dirk Groeneveld, Michal Guerquin, and Michael Schmitz. 2021. From ’F’ to ’a’ on the N.Y. Regents Science Exams: An Overview of the Aristo Project. arXiv:1909.01958
  • Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. 2021. Training Verifiers to Solve Math Word Problems. arXiv:2110.14168
  • Couperus (2023) Jelle Couperus. 2023. Large Language Models and Mathematical Understanding. Master’s thesis.
  • Dagan et al. (2023) Gautier Dagan, Frank Keller, and Alex Lascarides. 2023. Dynamic Planning with a LLM. arXiv (2023).
  • Davies et al. (2021) Alex Davies, Petar Veličković, Lars Buesing, Sam Blackwell, Daniel Zheng, Nenad Tomašev, Richard Tanburn, Peter Battaglia, Charles Blundell, András Juhász, et al. 2021. Advancing mathematics by guiding human intuition with AI. Nature 600, 7887 (2021), 70–74.
  • Davis and Steinglass (1997) Peggy Cooper Davis and Elizabeth Ehrenfest Steinglass. 1997. A dialogue about Socratic teaching. NYU Rev. L. & Soc. Change 23 (1997), 249.
  • Deng and Mineiro (2024) Yihe Deng and Paul Mineiro. 2024. Flow-DPO: Improving LLM Mathematical Reasoning through Online Multi-Agent Learning. arXiv (2024).
  • Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. arXiv:1810.04805
  • Ding et al. (2023) Ning Ding, Yulin Chen, Bokai Xu, Yujia Qin, Shengding Hu, Zhiyuan Liu, Maosong Sun, and Bowen Zhou. 2023. Enhancing Chat Language Models by Scaling High-quality Instructional Conversations. In EMNLP. 3029–3051.
  • Ding et al. (2024) Yuyang Ding, Hanglei Hu, Jie Zhou, Qin Chen, Bo Jiang, and Liang He. 2024. Boosting Large Language Models with Socratic Method for Conversational Mathematics Teaching. In CIKM. 3730–3735.
  • Drori et al. (2022) Iddo Drori, Sarah Zhang, Reece Shuttleworth, Leonard Tang, Albert Lu, Elizabeth Ke, Kevin Liu, Linda Chen, Sunny Tran, Newman Cheng, et al. 2022. A neural network solves, explains, and generates university math problems by program synthesis and few-shot learning at human level. PNAS 119, 32 (2022), e2123433119.
  • Dua et al. (2019) Dheeru Dua, Yizhong Wang, Pradeep Dasigi, Gabriel Stanovsky, Sameer Singh, and Matt Gardner. 2019. DROP: A Reading Comprehension Benchmark Requiring Discrete Reasoning Over Paragraphs. In NAACL. 2368–2378.
  • Duan et al. (2023) Jinhao Duan, Hao Cheng, Shiqi Wang, Chenan Wang, Alex Zavalny, Renjing Xu, Bhavya Kailkhura, and Kaidi Xu. 2023. Shifting attention to relevance: Towards the uncertainty estimation of large language models. arXiv (2023).
  • Elazar et al. (2019) Yanai Elazar, Abhijit Mahabal, Deepak Ramachandran, Tania Bedrax-Weiss, and Dan Roth. 2019. How Large Are Lions? Inducing Distributions over Quantitative Attributes. In ACL. 3973–3983.
  • Ethayarajh et al. (2024) Kawin Ethayarajh, Winnie Xu, Niklas Muennighoff, Dan Jurafsky, and Douwe Kiela. 2024. Kto: Model alignment as prospect theoretic optimization. arXiv (2024).
  • Feigenbaum et al. (1963) Edward A Feigenbaum, Julian Feldman, et al. 1963. Computers and thought. Vol. 7.
  • Feng et al. (2022) Yu Feng, Jing Zhang, Xiaokang Zhang, Lemao Liu, Cuiping Li, and Hong Chen. 2022. Injecting Numerical Reasoning Skills into Knowledge Base Question Answering Models. arXiv:2112.06109
  • First et al. (2023) Emily First, Markus N Rabe, Talia Ringer, and Yuriy Brun. 2023. Baldur: whole-proof generation and repair with large language models. arXiv (2023).
  • Fletcher (1985) Charles R Fletcher. 1985. Understanding and solving arithmetic word problems: A computer simulation. Behavior Research Methods, Instruments, & Computers 17, 5 (1985), 565–571.
  • Forbes and Choi (2017) Maxwell Forbes and Yejin Choi. 2017. Verb Physics: Relative Physical Knowledge of Actions and Objects. In ACL. 266–276.
  • Fu et al. (2023) Yao Fu, Hao Peng, Ashish Sabharwal, Peter Clark, and Tushar Khot. 2023. Complexity-Based Prompting for Multi-step Reasoning. In ICLR.
  • Gao et al. (2023) Luyu Gao, Aman Madaan, Shuyan Zhou, Uri Alon, Pengfei Liu, Yiming Yang, Jamie Callan, and Graham Neubig. 2023. Pal: Program-aided language models. In ICML. 10764–10799.
  • Gawlikowski et al. (2023) Jakob Gawlikowski, Cedrique Rovile Njieutcheu Tassi, Mohsin Ali, Jongseok Lee, Matthias Humt, Jianxiang Feng, Anna Kruspe, Rudolph Triebel, Peter Jung, Ribana Roscher, et al. 2023. A survey of uncertainty in deep neural networks. Artificial Intelligence Review 56, Suppl 1 (2023), 1513–1589.
  • Gelernter et al. (1960) Herbert Gelernter, James R Hansen, and Donald W Loveland. 1960. Empirical explorations of the geometry theorem machine. In western joint IRE-AIEE-ACM computer conference. 143–149.
  • Geva et al. (2020) Mor Geva, Ankit Gupta, and Jonathan Berant. 2020. Injecting Numerical Reasoning Skills into Language Models. In ACL, Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel Tetreault (Eds.). 946–958.
  • GLM et al. (2024) Team GLM, Aohan Zeng, Bin Xu, Bowen Wang, Chenhui Zhang, Da Yin, Dan Zhang, Diego Rojas, Guanyu Feng, Hanlin Zhao, et al. 2024. Chatglm: A family of large language models from glm-130b to glm-4 all tools. arXiv (2024).
  • Gonthier et al. (2013) Georges Gonthier, Andrea Asperti, Jeremy Avigad, Yves Bertot, Cyril Cohen, François Garillot, Stéphane Le Roux, Assia Mahboubi, Russell O’Connor, Sidi Ould Biha, et al. 2013. A machine-checked proof of the odd order theorem. In ITP. 163–179.
  • Gou et al. (2023a) Zhibin Gou, Zhihong Shao, Yeyun Gong, Yelong Shen, Yujiu Yang, Nan Duan, and Weizhu Chen. 2023a. CRITIC: Large Language Models Can Self-Correct with Tool-Interactive Critiquing. arXiv:2305.11738
  • Gou et al. (2023b) Zhibin Gou, Zhihong Shao, Yeyun Gong, Yujiu Yang, Minlie Huang, Nan Duan, Weizhu Chen, et al. 2023b. Tora: A tool-integrated reasoning agent for mathematical problem solving. arXiv (2023).
  • Grabowski et al. (2015) Adam Grabowski, Artur Korniłowicz, and Adam Naumowicz. 2015. Four Decades of Mizar: Foreword. Journal of Automated Reasoning 55 (2015), 191–198.
  • Hales et al. (2017) Thomas Hales, Mark Adams, Gertrud Bauer, Tat Dat Dang, John Harrison, Hoang Le Truong, Cezary Kaliszyk, Victor Magron, Sean McLaughlin, Tat Thang Nguyen, et al. 2017. A formal proof of the Kepler conjecture. In Forum of mathematics, Pi, Vol. 5. e2.
  • Han et al. (2022) Jesse Michael Han, Jason Rute, Yuhuai Wu, Edward Ayers, and Stanislas Polu. 2022. Proof Artifact Co-Training for Theorem Proving with Language Models. In ICLR.
  • Han et al. (2024) Simeng Han, Hailey Schoelkopf, Yilun Zhao, Zhenting Qi, Martin Riddell, Wenfei Zhou, James Coady, David Peng, Yujie Qiao, Luke Benson, Lucy Sun, Alexander Wardle-Solano, Hannah Szabó, Ekaterina Zubova, Matthew Burtell, Jonathan Fan, Yixin Liu, Brian Wong, Malcolm Sailor, Ansong Ni, Linyong Nan, Jungo Kasai, Tao Yu, Rui Zhang, Alexander R. Fabbri, Wojciech Kryscinski, Semih Yavuz, Ye Liu, Xi Victoria Lin, Shafiq Joty, Yingbo Zhou, Caiming Xiong, Rex Ying, Arman Cohan, and Dragomir Radev. 2024. FOLIO: Natural Language Reasoning with First-Order Logic. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, EMNLP 2024, Miami, FL, USA, November 12-16, 2024, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.). Association for Computational Linguistics, 22017–22031. https://aclanthology.org/2024.emnlp-main.1229
  • Hao et al. (2023a) Shibo Hao, Yi Gu, Haodi Ma, Joshua Jiahua Hong, Zhen Wang, Daisy Zhe Wang, and Zhiting Hu. 2023a. Reasoning with language model is planning with world model. arXiv (2023).
  • Hao et al. (2023b) Shibo Hao, Tianyang Liu, Zhen Wang, and Zhiting Hu. 2023b. ToolkenGPT: Augmenting Frozen Language Models with Massive Tools via Tool Embeddings. arXiv:2305.11554
  • Hao et al. (2022) Yihan Hao, Mingliang Zhang, Fei Yin, and Linlin Huang. 2022. PGDP5K: A Diagram Parsing Dataset for Plane Geometry Problems. ICPR (2022), 1763–1769.
  • He et al. (2022) Hangfeng He, Hongming Zhang, and Dan Roth. 2022. Rethinking with retrieval: Faithful large language model inference. arXiv (2022).
  • He-Yueya et al. (2023) Joy He-Yueya, Gabriel Poesia, Rose E Wang, and Noah D Goodman. 2023. Solving math word problems by combining language models with symbolic solvers. arXiv (2023).
  • Hendrycks et al. (2021a) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt. 2021a. Measuring Massive Multitask Language Understanding. ICLR (2021).
  • Hendrycks et al. (2021b) Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt. 2021b. Measuring Mathematical Problem Solving With the MATH Dataset. In NeurIPS Datasets and Benchmarks Track (Round 2).
  • Hossain et al. (2019) MD Zakir Hossain, Ferdous Sohel, Mohd Fairuz Shiratuddin, and Hamid Laga. 2019. A comprehensive survey of deep learning for image captioning. ACM Computing Surveys (CsUR) 51, 6 (2019), 1–36.
  • Hosseini et al. (2024) Arian Hosseini, Xingdi Yuan, Nikolay Malkin, Aaron Courville, Alessandro Sordoni, and Rishabh Agarwal. 2024. V-star: Training verifiers for self-taught reasoners. arXiv (2024).
  • Hosseini et al. (2014) Mohammad Javad Hosseini, Hannaneh Hajishirzi, Oren Etzioni, and Nate Kushman. 2014. Learning to Solve Arithmetic Word Problems with Verb Categorization. In EMNLP. 523–533.
  • Hsieh et al. (2023) Cheng-Yu Hsieh, Si-An Chen, Chun-Liang Li, Yasuhisa Fujii, Alexander Ratner, Chen-Yu Lee, Ranjay Krishna, and Tomas Pfister. 2023. Tool Documentation Enables Zero-Shot Tool-Usage with Large Language Models. arXiv:2308.00675
  • Huang et al. (2019) Daniel Huang, Prafulla Dhariwal, Dawn Song, and Ilya Sutskever. 2019. GamePad: A Learning Environment for Theorem Proving. In ICLR.
  • Huang et al. (2016) Danqing Huang, Shuming Shi, Chin-Yew Lin, Jian Yin, and Wei-Ying Ma. 2016. How well do computers solve math word problems? large-scale dataset construction and evaluation. In ACL. 887–896.
  • Huang et al. (2024) Xijie Huang, Li Lyna Zhang, Kwang-Ting Cheng, Fan Yang, and Mao Yang. 2024. Fewer is More: Boosting Math Reasoning with Reinforced Context Pruning. In EMNLP. 13674–13695.
  • Hurst et al. (2024) Aaron Hurst, Adam Lerer, Adam P Goucher, Adam Perelman, Aditya Ramesh, Aidan Clark, AJ Ostrow, Akila Welihinda, Alan Hayes, Alec Radford, et al. 2024. Gpt-4o system card. arXiv (2024).
  • Imani et al. (2023) Shima Imani, Liang Du, and Harsh Shrivastava. 2023. Mathprompter: Mathematical reasoning using large language models. arXiv (2023).
  • Irving et al. (2016) Geoffrey Irving, Christian Szegedy, Alexander A Alemi, Niklas Eén, François Chollet, and Josef Urban. 2016. Deepmath-deep sequence models for premise selection. NeurIPS 29 (2016).
  • Jelassi et al. (2023) Samy Jelassi, Stéphane d’Ascoli, Carles Domingo-Enrich, Yuhuai Wu, Yuanzhi Li, and François Charton. 2023. Length generalization in arithmetic transformers. arXiv (2023).
  • Ji et al. (2023) Ziwei Ji, Nayeon Lee, Rita Frieske, Tiezheng Yu, Dan Su, Yan Xu, Etsuko Ishii, Ye Jin Bang, Andrea Madotto, and Pascale Fung. 2023. Survey of hallucination in natural language generation. Comput. Surveys 55, 12 (2023), 1–38.
  • Jiang et al. (2021) Albert Qiaochu Jiang, Wenda Li, Jesse Michael Han, and Yuhuai Wu. 2021. LISA: Language models of ISAbelle proofs. In AITP. 378–392.
  • Jiang et al. (2022) Albert Q. Jiang, Wenda Li, Szymon Tworkowski, Konrad Czechowski, Tomasz Odrzygóźdź, Piotr Miłoś, Yuhuai Wu, and Mateja Jamnik. 2022. Thor: Wielding Hammers to Integrate Language Models and Automated Theorem Provers. arXiv:2205.10893
  • Jiang et al. (2023b) Albert Qiaochu Jiang, Sean Welleck, Jin Peng Zhou, Timothee Lacroix, Jiacheng Liu, Wenda Li, Mateja Jamnik, Guillaume Lample, and Yuhuai Wu. 2023b. Draft, Sketch, and Prove: Guiding Formal Theorem Provers with Informal Proofs. In ICLR.
  • Jiang et al. (2024) Dongwei Jiang, Marcio Fonseca, and Shay B Cohen. 2024. LeanReasoner: Boosting Complex Logical Reasoning with Lean. arXiv (2024).
  • Jiang et al. (2023a) Song Jiang, Zahra Shakeri, Aaron Chan, Maziar Sanjabi, Hamed Firooz, Yinglong Xia, Bugra Akyildiz, Yizhou Sun, Jinchao Li, Qifan Wang, et al. 2023a. Resprompt: Residual connection prompting advances multi-step reasoning in large language models. arXiv (2023).
  • Jie et al. (2022) Zhanming Jie, Jierui Li, and Wei Lu. 2022. Learning to Reason Deductively: Math Word Problem Solving as Complex Relation Extraction. arXiv:2203.10316
  • Kaliszyk et al. (2017) Cezary Kaliszyk, François Chollet, and Christian Szegedy. 2017. HolStep: A Machine Learning Dataset for Higher-order Logic Theorem Proving. In ICLR.
  • Kalyan et al. (2021) Ashwin Kalyan, Abhinav Kumar, Arjun Chandrasekaran, Ashish Sabharwal, and Peter Clark. 2021. How Much Coffee Was Consumed During EMNLP 2019? Fermi Problems: A New Reasoning Challenge for AI. In EMNLP. 7318–7328.
  • Kaplan et al. (2020) Jared Kaplan, Sam McCandlish, Tom Henighan, Tom B Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei. 2020. Scaling laws for neural language models. arXiv (2020).
  • Kazemi et al. (2022) Mehran Kazemi, Najoung Kim, Deepti Bhatia, Xin Xu, and Deepak Ramachandran. 2022. Lambada: Backward chaining for automated reasoning in natural language. arXiv (2022).
  • Kenton and Toutanova (2019) Jacob Devlin Ming-Wei Chang Kenton and Lee Kristina Toutanova. 2019. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. In NAACL. 4171–4186.
  • Khalifa et al. (2023) Muhammad Khalifa, Lajanugen Logeswaran, Moontae Lee, Honglak Lee, and Lu Wang. 2023. Discriminator-Guided Multi-step Reasoning with Language Models. arXiv (2023).
  • Kim et al. (2020) Bugeun Kim, Kyung Seo Ki, Donggeon Lee, and Gahgene Gweon. 2020. Point to the Expression: Solving Algebraic Word Problems using the Expression-Pointer Transformer Model. In EMNLP. Online, 3768–3779.
  • Kim et al. (2022) Bugeun Kim, Kyung Seo Ki, Sangkyu Rhim, and Gahgene Gweon. 2022. EPT-X: An Expression-Pointer Transformer model that generates eXplanations for numbers. In ACL. 4442–4458.
  • Kojima et al. ([n. d.]) Takeshi Kojima, Shixiang Shane Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa. [n. d.]. Large Language Models are Zero-Shot Reasoners. ([n. d.]).
  • Koncel-Kedziorski et al. (2015) Rik Koncel-Kedziorski, Hannaneh Hajishirzi, Ashish Sabharwal, Oren Etzioni, and Siena Dumas Ang. 2015. Parsing algebraic word problems into equations. TACL 3 (2015), 585–597.
  • Koncel-Kedziorski et al. (2016) Rik Koncel-Kedziorski, Subhro Roy, Aida Amini, Nate Kushman, and Hannaneh Hajishirzi. 2016. MAWPS: A math word problem repository. In NAACL. 1152–1157.
  • Kumar et al. (2024) Aviral Kumar, Vincent Zhuang, Rishabh Agarwal, Yi Su, John D Co-Reyes, Avi Singh, Kate Baumli, Shariq Iqbal, Colton Bishop, Rebecca Roelofs, et al. 2024. Training language models to self-correct via reinforcement learning. arXiv (2024).
  • Kushman et al. (2014) Nate Kushman, Yoav Artzi, Luke Zettlemoyer, and Regina Barzilay. 2014. Learning to Automatically Solve Algebra Word Problems. In ACL. 271–281.
  • Lai et al. (2024) Xin Lai, Zhuotao Tian, Yukang Chen, Senqiao Yang, Xiangru Peng, and Jiaya Jia. 2024. Step-dpo: Step-wise preference optimization for long-chain reasoning of llms. arXiv (2024).
  • Lample et al. ([n. d.]) Guillaume Lample, Marie-Anne Lachaux, Thibaut Lavril, Xavier Martinet, Amaury Hayat, Gabriel Ebner, Aurélien Rodriguez, and Timothée Lacroix. [n. d.]. HyperTree Proof Search for Neural Theorem Proving. ([n. d.]).
  • Lample et al. (2022) Guillaume Lample, Timothee Lacroix, Marie-Anne Lachaux, Aurelien Rodriguez, Amaury Hayat, Thibaut Lavril, Gabriel Ebner, and Xavier Martinet. 2022. Hypertree proof search for neural theorem proving. NeurIPS 35 (2022), 26337–26349.
  • Lei et al. (2023) Bin Lei, Chunhua Liao, Caiwen Ding, et al. 2023. Boosting logical reasoning in large language models through a new framework: The graph of thought. arXiv (2023).
  • Lewis et al. (2020) Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed, Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020. BART: Denoising Sequence-to-Sequence Pre-training for Natural Language Generation, Translation, and Comprehension. In ACL. 7871–7880.
  • Lewkowycz et al. (2022) Aitor Lewkowycz, Anders Andreassen, David Dohan, Ethan Dyer, Henryk Michalewski, Vinay Ramasesh, Ambrose Slone, Cem Anil, Imanol Schlag, Theo Gutman-Solo, et al. 2022. Solving quantitative reasoning problems with language models. NeurIPS 35 (2022), 3843–3857.
  • Li et al. (2023a) Guohao Li, Hasan Abed Al Kader Hammoud, Hani Itani, Dmitrii Khizbullin, and Bernard Ghanem. 2023a. CAMEL: Communicative Agents for” Mind” Exploration of Large Language Model Society. In NeurIPS.
  • Li et al. (2021) Wenda Li, Lei Yu, Yuhuai Wu, and Lawrence C Paulson. 2021. IsarStep: a Benchmark for High-level Mathematical Reasoning. In ICLR.
  • Li et al. (2023b) Yifei Li, Zeqi Lin, Shizhuo Zhang, Qiang Fu, Bei Chen, Jian-Guang Lou, and Weizhu Chen. 2023b. Making language models better reasoners with step-aware verifier. In ACL. 5315–5333.
  • Li et al. (2022) Zhongli Li, Wenxuan Zhang, Chao Yan, Qingyu Zhou, Chao Li, Hongzhi Liu, and Yunbo Cao. 2022. Seeking Patterns, Not Just Memorizing Procedures: Contrastive Learning for Solving Math Word Problems.
  • Liang et al. (2022) Zhenwen Liang, Jipeng Zhang, Lei Wang, Wei Qin, Yunshi Lan, Jie Shao, and Xiangliang Zhang. 2022. MWP-BERT: Numeracy-augmented Pre-Training for Math Word Problem Solving.
  • Lightman et al. (2024) Hunter Lightman, Vineet Kosaraju, Yuri Burda, Harrison Edwards, Bowen Baker, Teddy Lee, Jan Leike, John Schulman, Ilya Sutskever, and Karl Cobbe. 2024. Let’s Verify Step by Step. In ICLR.
  • Ling et al. (2017) Wang Ling, Dani Yogatama, Chris Dyer, and Phil Blunsom. 2017. Program Induction by Rationale Generation: Learning to Solve and Explain Algebraic Word Problems. In ACL. 158–167.
  • Ling et al. (2023) Zhan Ling, Yunhao Fang, Xuanlin Li, Zhiao Huang, Mingu Lee, Roland Memisevic, and Hao Su. 2023. Deductive Verification of Chain-of-Thought Reasoning. arXiv (2023).
  • Liu et al. (2023a) Bo Liu, Yuqian Jiang, Xiaohan Zhang, Qiang Liu, Shiqi Zhang, Joydeep Biswas, and Peter Stone. 2023a. Llm+ p: Empowering large language models with optimal planning proficiency. arXiv (2023).
  • Liu et al. (2022) Jiachang Liu, Dinghan Shen, Yizhe Zhang, Bill Dolan, Lawrence Carin, and Weizhu Chen. 2022. What Makes Good In-Context Examples for GPT-3?. In DeeLIO. 100–114.
  • Liu and Low (2023) Tiedong Liu and Bryan Kian Hsiang Low. 2023. Goat: Fine-tuned llama outperforms gpt-4 on arithmetic tasks. arXiv (2023).
  • Liu et al. (2024) Wentao Liu, Qianjun Pan, Yi Zhang, Zhuo Liu, Ji Wu, Jie Zhou, Aimin Zhou, Qin Chen, Bo Jiang, and Liang He. 2024. Cmm-math: A chinese multimodal math dataset to evaluate and enhance the mathematics reasoning of large multimodal models. arXiv (2024).
  • Liu et al. (2019) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019. RoBERTa: A Robustly Optimized BERT Pretraining Approach.
  • Liu et al. (2023b) Yixin Liu, Avi Singh, C Daniel Freeman, John D Co-Reyes, and Peter J Liu. 2023b. Improving large language model fine-tuning for solving math problems. arXiv (2023).
  • Long (2023) Jieyi Long. 2023. Large Language Model Guided Tree-of-Thought. arXiv (2023).
  • Lu et al. (2023a) Pan Lu, Hritik Bansal, Tony Xia, Jiacheng Liu, Chunyuan Li, Hannaneh Hajishirzi, Hao Cheng, Kai-Wei Chang, Michel Galley, and Jianfeng Gao. 2023a. Mathvista: Evaluating mathematical reasoning of foundation models in visual contexts. arXiv (2023).
  • Lu et al. (2021a) Pan Lu, Ran Gong, Shibiao Jiang, Liang Qiu, Siyuan Huang, Xiaodan Liang, and Song-chun Zhu. 2021a. Inter-GPS: Interpretable Geometry Problem Solving with Formal Language and Symbolic Reasoning. In ACL. 6774–6786.
  • Lu et al. (2023b) Pan Lu, Baolin Peng, Hao Cheng, Michel Galley, Kai-Wei Chang, Ying Nian Wu, Song-Chun Zhu, and Jianfeng Gao. 2023b. Chameleon: Plug-and-Play Compositional Reasoning with Large Language Models. arXiv:2304.09842
  • Lu et al. (2023c) Pan Lu, Liang Qiu, Kai-Wei Chang, Ying Nian Wu, Song-Chun Zhu, Tanmay Rajpurohit, Peter Clark, and Ashwin Kalyan. 2023c. Dynamic Prompt Learning via Policy Gradient for Semi-structured Mathematical Reasoning. In ICLR.
  • Lu et al. (2021b) Pan Lu, Liang Qiu, Jiaqi Chen, Tony Xia, Yizhou Zhao, Wei Zhang, Zhou Yu, Xiaodan Liang, and Song-Chun Zhu. 2021b. IconQA: A New Benchmark for Abstract Diagram Understanding and Visual Language Reasoning. In NeurIPS Datasets and Benchmarks Track.
  • Lu et al. (2023d) Pan Lu, Liang Qiu, Wenhao Yu, Sean Welleck, and Kai-Wei Chang. 2023d. A Survey of Deep Learning for Mathematical Reasoning. In ACL, Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). 14605–14631.
  • Luo et al. (2023) Haipeng Luo, Qingfeng Sun, Can Xu, Pu Zhao, Jianguang Lou, Chongyang Tao, Xiubo Geng, Qingwei Lin, Shifeng Chen, and Dongmei Zhang. 2023. Wizardmath: Empowering mathematical reasoning for large language models via reinforced evol-instruct. arXiv (2023).
  • Luo et al. (2024) Liangchen Luo, Yinxiao Liu, Rosanne Liu, Samrat Phatale, Harsh Lara, Yunxuan Li, Lei Shu, Yun Zhu, Lei Meng, Jiao Sun, et al. 2024. Improve Mathematical Reasoning in Language Models by Automated Process Supervision. arXiv (2024).
  • Luong et al. (2024) Trung Quoc Luong, Xinbo Zhang, Zhanming Jie, Peng Sun, Xiaoran Jin, and Hang Li. 2024. Reft: Reasoning with reinforced fine-tuning. arXiv (2024).
  • Ma et al. (2023) Qianli Ma, Haotian Zhou, Tingkai Liu, Jianbo Yuan, Pengfei Liu, Yang You, and Hongxia Yang. 2023. Let’s reward step by step: Step-Level reward model as the Navigators for Reasoning. arXiv (2023).
  • Macina et al. (2023) Jakub Macina, Nico Daheim, Sankalan Pal Chowdhury, Tanmay Sinha, Manu Kapur, Iryna Gurevych, and Mrinmaya Sachan. 2023. Mathdial: A dialogue tutoring dataset with rich pedagogical properties grounded in math reasoning problems. arXiv (2023).
  • Madaan et al. (2023) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, et al. 2023. Self-refine: Iterative refinement with self-feedback. arXiv (2023).
  • Matzakos et al. (2023) Nikolaos Matzakos, Spyridon Doukakis, and Maria Moundridou. 2023. Learning Mathematics with Large Language Models: A Comparative Study with Computer Algebra Systems and Other Tools. International Journal of Emerging Technologies in Learning (Online) 18, 20 (2023), 51.
  • Maynez et al. (2020) Joshua Maynez, Shashi Narayan, Bernd Bohnet, and Ryan McDonald. 2020. On Faithfulness and Factuality in Abstractive Summarization. In ACL. 1906–1919.
  • Megill and Wheeler (2019) Norman Megill and David A Wheeler. 2019. Metamath: a computer language for mathematical proofs. Lulu. com.
  • Miao et al. (2023) Ning Miao, Yee Whye Teh, and Tom Rainforth. 2023. Selfcheck: Using llms to zero-shot check their own step-by-step reasoning. arXiv (2023).
  • Miao et al. (2020) Shen-Yun Miao, Chao-Chun Liang, and Keh-Yih Su. 2020. A Diverse Corpus for Evaluating and Developing English Math Word Problem Solvers. In ACL. 975–984.
  • Min et al. (2023) Bonan Min, Hayley Ross, Elior Sulem, Amir Pouran Ben Veyseh, Thien Huu Nguyen, Oscar Sainz, Eneko Agirre, Ilana Heintz, and Dan Roth. 2023. Recent Advances in Natural Language Processing via Large Pre-Trained Language Models: A Survey. ACM Comput. Surv. 56, 2 (2023).
  • Min et al. (2022) Sewon Min, Mike Lewis, Luke Zettlemoyer, and Hannaneh Hajishirzi. 2022. MetaICL: Learning to Learn In Context. In NAACL. 2791–2809.
  • Mishra et al. (2022a) Swaroop Mishra, Matthew Finlayson, Pan Lu, Leonard Tang, Sean Welleck, Chitta Baral, Tanmay Rajpurohit, Oyvind Tafjord, Ashish Sabharwal, Peter Clark, et al. 2022a. LILA: A Unified Benchmark for Mathematical Reasoning. In EMNLP. 5807–5832.
  • Mishra et al. (2022b) Swaroop Mishra, Arindam Mitra, Neeraj Varshney, Bhavdeep Sachdeva, Peter Clark, Chitta Baral, and Ashwin Kalyan. 2022b. NUMGLUE: A Suite of Fundamental yet Challenging Mathematical Reasoning Tasks. In ACL. 3505–3523.
  • Mitra and Baral (2016) Arindam Mitra and Chitta Baral. 2016. Learning to use formulas to solve simple arithmetic problems. In ACL. 2144–2153.
  • Mo and Xin (2023) Shentong Mo and Miao Xin. 2023. Tree of Uncertain Thoughts Reasoning for Large Language Models. arXiv (2023).
  • Muffo et al. ([n. d.]) Matteo Muffo, Aldo Cocco, and Enrico Bertino. [n. d.]. Evaluating Transformer Language Models on Arithmetic Operations Using Number Decomposition. ([n. d.]).
  • Naik et al. (2023) Ranjita Naik, Varun Chandrasekaran, Mert Yuksekgonul, Hamid Palangi, and Besmira Nushi. 2023. Diversity of Thought Improves Reasoning Abilities of Large Language Models. arXiv (2023).
  • NAKAMOTO et al. (2023) Ryosuke NAKAMOTO, Brendan Flanagan, Taisei Yamauchi, Dai Yilling, Kyosuke Takami, and Horoaki Ogata. 2023. Enhancing Automated Scoring of Math Self-Explanation Quality using LLM-Generated Datasets: A Semi-Supervised Approach. (2023).
  • Nelson (1980) Leonard Nelson. 1980. The socratic method. Thinking: The Journal of Philosophy for Children 2, 2, 34–38.
  • Newell et al. (1957) A. Newell, J. C. Shaw, and H. A. Simon. 1957. Empirical Explorations of the Logic Theory Machine: A Case Study in Heuristic. 218–230.
  • Nogueira et al. (2021) Rodrigo Nogueira, Zhiying Jiang, and Jimmy Lin. 2021. Investigating the limitations of transformers with simple arithmetic tasks. arXiv (2021).
  • Noorbakhsh et al. (2021) Kimia Noorbakhsh, Modar Sulaiman, Mahdi Sharifi, Kallol Roy, and Pooyan Jamshidi. 2021. Pretrained Language Models are Symbolic Mathematics Solvers too! arXiv (2021).
  • Nye et al. (2021) Maxwell Nye, Anders Johan Andreassen, Guy Gur-Ari, Henryk Michalewski, Jacob Austin, David Bieber, David Dohan, Aitor Lewkowycz, Maarten Bosma, David Luan, Charles Sutton, and Augustus Odena. 2021. Show Your Work: Scratchpads for Intermediate Computation with Language Models.
  • Olausson et al. (2023) Theo X Olausson, Alex Gu, Benjamin Lipkin, Cedegao E Zhang, Armando Solar-Lezama, Joshua B Tenenbaum, and Roger Levy. 2023. LINC: A neurosymbolic approach for logical reasoning by combining language models with first-order logic provers. arXiv (2023).
  • OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. arXiv:2303.08774
  • OpenAI (2024) OpenAI. 2024. OpenAI O1 System Card. https://openai.com/index/openai-o1-system-card/
  • Paranjape et al. (2023) Bhargavi Paranjape, Scott Lundberg, Sameer Singh, Hannaneh Hajishirzi, Luke Zettlemoyer, and Marco Tulio Ribeiro. 2023. ART: Automatic multi-step reasoning and tool-use for large language models. arXiv:2303.09014
  • Parisi et al. (2022) Aaron Parisi, Yao Zhao, and Noah Fiedel. 2022. TALM: Tool Augmented Language Models. arXiv:2205.12255
  • Paster et al. (2023) Keiran Paster, Marco Dos Santos, Zhangir Azerbayev, and Jimmy Ba. 2023. OpenWebMath: An Open Dataset of High-Quality Mathematical Web Text. arXiv:2310.06786
  • Patel et al. (2021) Arkil Patel, Satwik Bhattamishra, and Navin Goyal. 2021. Are NLP Models really able to Solve Simple Math Word Problems?. In NAACL. 2080–2094.
  • Peng et al. (2021) Shuai Peng, Ke Yuan, Liangcai Gao, and Zhi Tang. 2021. MathBERT: A Pre-Trained Model for Mathematical Formula Understanding.
  • Pitis et al. (2023) Silviu Pitis, Michael R Zhang, Andrew Wang, and Jimmy Ba. 2023. Boosted prompt ensembles for large language models. arXiv (2023).
  • Polu and Sutskever (2020a) Stanislas Polu and Ilya Sutskever. 2020a. Generative Language Modeling for Automated Theorem Proving. arXiv:2009.03393
  • Polu and Sutskever (2020b) Stanislas Polu and Ilya Sutskever. 2020b. Generative Language Modeling for Automated Theorem Proving. CoRR abs/2009.03393 (2020).
  • Qi et al. (2023) Jingyuan Qi, Zhiyang Xu, Ying Shen, Minqian Liu, dingnan jin, Qifan Wang, and Lifu Huang. 2023. The Art of SOCRATIC QUESTIONING: Recursive Thinking with Large Language Models. In EMNLP.
  • Qian et al. (2023) Cheng Qian, Chi Han, Yi R Fung, Yujia Qin, Zhiyuan Liu, and Heng Ji. 2023. Creator: Tool creation for disentangling abstract and concrete reasoning of large language models. arXiv (2023).
  • Qiao et al. (2024) Runqi Qiao, Qiuna Tan, Guanting Dong, Minhui Wu, Chong Sun, Xiaoshuai Song, Zhuoma GongQue, Shanglin Lei, Zhe Wei, Miaoxuan Zhang, Runfeng Qiao, Yifan Zhang, Xiao Zong, Yida Xu, Muxi Diao, Zhimin Bao, Chen Li, and Honggang Zhang. 2024. We-Math: Does Your Large Multimodal Model Achieve Human-like Mathematical Reasoning? arXiv:2407.01284
  • Qiao et al. (2023) Shuofei Qiao, Yixin Ou, Ningyu Zhang, Xiang Chen, Yunzhi Yao, Shumin Deng, Chuanqi Tan, Fei Huang, and Huajun Chen. 2023. Reasoning with Language Model Prompting: A Survey. In ACL. 5368–5393.
  • Qin et al. (2020) Jinghui Qin, Lihui Lin, Xiaodan Liang, Rumin Zhang, and Liang Lin. 2020. Semantically-Aligned Universal Tree-Structured Solver for Math Word Problems. In EMNLP. 3780–3789.
  • Qiu et al. (2020) Xipeng Qiu, Tianxiang Sun, Yige Xu, Yunfan Shao, Ning Dai, and Xuanjing Huang. 2020. Pre-trained models for natural language processing: A survey. Science China Technological Sciences 63, 10 (2020), 1872–1897.
  • Radford et al. (2018) Alec Radford, Karthik Narasimhan, Tim Salimans, and Ilya Sutskever. 2018. Improving Language Understanding by Generative Pre-Training. (June 2018), 12.
  • Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, and Ilya Sutskever. 2019. Language Models are Unsupervised Multitask Learners. (Feb. 2019), 24.
  • Rafailov et al. (2024) Rafael Rafailov, Archit Sharma, Eric Mitchell, Christopher D Manning, Stefano Ermon, and Chelsea Finn. 2024. Direct preference optimization: Your language model is secretly a reward model. NeurIPS 36 (2024).
  • Rawte et al. (2023) Vipula Rawte, Amit Sheth, and Amitava Das. 2023. A survey of hallucination in large foundation models. arXiv (2023).
  • Raffel et al. ([n. d.]) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J Liu. [n. d.]. Exploring the Limits of Transfer Learning with a Unified Text-to-Text Transformer. ([n. d.]).
  • Roy and Roth (2015) Subhro Roy and Dan Roth. 2015. Solving General Arithmetic Word Problems. In EMNLP. 1743–1752.
  • Roy and Roth (2017) Subhro Roy and Dan Roth. 2017. Unit dependency graph and its application to arithmetic word problem solving. In AAAI, Vol. 31.
  • Roy and Roth (2018) Subhro Roy and Dan Roth. 2018. Mapping to declarative knowledge for word problem solving. TACL 6 (2018), 159–172.
  • Roy et al. (2015) Subhro Roy, Tim Vieira, and Dan Roth. 2015. Reasoning about quantities in natural language. TACL 3 (2015), 1–13.
  • Rudnicki (1992) Piotr Rudnicki. 1992. An overview of the Mizar project. In Proceedings of the 1992 Workshop on Types for Proofs and Programs. 311–330.
  • Saxton et al. (2019) David Saxton, Edward Grefenstette, Felix Hill, and Pushmeet Kohli. 2019. Analysing Mathematical Reasoning Abilities of Neural Models. In ICLR.
  • Scharpf et al. (2022) Philipp Scharpf, Moritz Schubotz, and Bela Gipp. 2022. Mining mathematical documents for question answering via unsupervised formula labeling. In ICDL. 1–11.
  • Schick et al. (2023) Timo Schick, Jane Dwivedi-Yu, Roberto Dessì, Roberta Raileanu, Maria Lomeli, Luke Zettlemoyer, Nicola Cancedda, and Thomas Scialom. 2023. Toolformer: Language Models Can Teach Themselves to Use Tools. arXiv:2302.04761
  • Schulman et al. (2017) John Schulman, Filip Wolski, Prafulla Dhariwal, Alec Radford, and Oleg Klimov. 2017. Proximal policy optimization algorithms. arXiv (2017).
  • Sel et al. (2023) Bilgehan Sel, Ahmad Al-Tawaha, Vanshaj Khattar, Lu Wang, Ruoxi Jia, and Ming Jin. 2023. Algorithm of thoughts: Enhancing exploration of ideas in large language models. arXiv (2023).
  • Seo et al. (2015) Minjoon Seo, Hannaneh Hajishirzi, Ali Farhadi, Oren Etzioni, and Clint Malcolm. 2015. Solving Geometry Problems: Combining Text and Diagram Interpretation. In EMNLP. 1466–1476.
  • Shen et al. (2021) Jianhao Shen, Yichun Yin, Lin Li, Lifeng Shang, Xin Jiang, Ming Zhang, and Qun Liu. 2021. Generate & Rank: A Multi-task Framework for Math Word Problems.
  • Shi et al. (2022) Freda Shi, Mirac Suzgun, Markus Freitag, Xuezhi Wang, Suraj Srivats, Soroush Vosoughi, Hyung Won Chung, Yi Tay, Sebastian Ruder, Denny Zhou, et al. 2022. Language models are multilingual chain-of-thought reasoners. In ICLR.
  • Shi et al. (2015) Shuming Shi, Yuehui Wang, Chin-Yew Lin, Xiaojiang Liu, and Yong Rui. 2015. Automatically solving number word problems by semantic parsing and reasoning. In EMNLP. 1132–1142.
  • Shi et al. (2024) Wenhao Shi, Zhiqiang Hu, Yi Bin, Junhua Liu, Yang Yang, See-Kiong Ng, Lidong Bing, and Roy Ka-Wei Lee. 2024. Math-LLaVA: Bootstrapping Mathematical Reasoning for Multimodal Large Language Models.
  • Shinn et al. (2023) Noah Shinn, Federico Cassano, Ashwin Gopinath, Karthik R Narasimhan, and Shunyu Yao. 2023. Reflexion: Language agents with verbal reinforcement learning. In NeurIPS.
  • Shridhar et al. (2023) Kumar Shridhar, Harsh Jhamtani, Hao Fang, Benjamin Van Durme, Jason Eisner, and Patrick Xia. 2023. SCREWS: A Modular Framework for Reasoning with Revisions. arXiv (2023).
  • Shridhar et al. (2022) Kumar Shridhar, Jakub Macina, Mennatallah El-Assady, Tanmay Sinha, Manu Kapur, and Mrinmaya Sachan. 2022. Automatic Generation of Socratic Subquestions for Teaching Math Word Problems. arXiv (2022).
  • Shum et al. (2023) Kashun Shum, Shizhe Diao, and Tong Zhang. 2023. Automatic Prompt Augmentation and Selection with Chain-of-Thought from Labeled Data. In EMNLP. 12113–12139.
  • Slind and Norrish (2008) Konrad Slind and Michael Norrish. 2008. A brief overview of HOL4. In TPHOLs. 28–32.
  • Song et al. (2024) Peiyang Song, Kaiyu Yang, and Anima Anandkumar. 2024. Towards large language models as copilots for theorem proving in lean. arXiv (2024).
  • Spithourakis et al. (2016) Georgios Spithourakis, Isabelle Augenstein, and Sebastian Riedel. 2016. Numerically Grounded Language Models for Semantic Error Correction. In EMNLP. 987–992.
  • Spithourakis and Riedel (2018) GP Spithourakis and S Riedel. 2018. Numeracy for language models: Evaluating and improving their ability to predict numbers. In ACL, Vol. 56. 2104–2115.
  • Sun et al. (2023) Haotian Sun, Yuchen Zhuang, Lingkai Kong, Bo Dai, and Chao Zhang. 2023. AdaPlanner: Adaptive Planning from Feedback with Language Models. arXiv (2023).
  • Taylor et al. (2022) Ross Taylor, Marcin Kardas, Guillem Cucurull, Thomas Scialom, Anthony Hartshorn, Elvis Saravia, Andrew Poulton, Viktor Kerkez, and Robert Stojnic. 2022. Galactica: A Large Language Model for Science.
  • Team et al. (2023) Gemini Team, Rohan Anil, Sebastian Borgeaud, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, Katie Millican, et al. 2023. Gemini: a family of highly capable multimodal models. arXiv (2023).
  • Team (2024) Qwen Team. 2024. QwQ: Reflect Deeply on the Boundaries of the Unknown. https://qwenlm.github.io/blog/qwq-32b-preview/
  • Thoppilan et al. (2022) Romal Thoppilan, Daniel De Freitas, Jamie Hall, Noam Shazeer, Apoorv Kulshreshtha, Heng-Tze Cheng, Alicia Jin, Taylor Bos, Leslie Baker, Yu Du, et al. 2022. Lamda: Language models for dialog applications. arXiv (2022).
  • Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023a. Llama: Open and efficient foundation language models. arXiv (2023).
  • Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023b. Llama 2: Open foundation and fine-tuned chat models. arXiv (2023).
  • Trung et al. (2024) Luong Trung, Xinbo Zhang, Zhanming Jie, Peng Sun, Xiaoran Jin, and Hang Li. 2024. Reft: Reasoning with reinforced fine-tuning. In ACL. 7601–7614.
  • Upadhyay and Chang (2015) Shyam Upadhyay and Ming-Wei Chang. 2015. Draw: A challenging and diverse algebra word problem set. Technical Report. Citeseer.
  • Upadhyay and Chang (2017) Shyam Upadhyay and Ming-Wei Chang. 2017. Annotating Derivations: A New Evaluation Strategy and Dataset for Algebra Word Problems. In EACL. 494–504.
  • Upadhyay et al. (2023) Shriyash Upadhyay, Etan Ginsberg, and Chris Callison-Burch. 2023. Improving Mathematics Tutoring With A Code Scratchpad. In BEA. 20–28.
  • Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Łukasz Kaiser, and Illia Polosukhin. 2017. Attention is all you need. NeurIPS 30 (2017).
  • Wallace et al. (2019) Eric Wallace, Yizhong Wang, Sujian Li, Sameer Singh, and Matt Gardner. 2019. Do NLP Models Know Numbers? Probing Numeracy in Embeddings. In EMNLP. 5307–5315.
  • Wang and Komatsuzaki (2021) Ben Wang and Aran Komatsuzaki. 2021. GPT-J-6B: A 6 billion parameter autoregressive language model.
  • Wang et al. (2021) Cunxiang Wang, Boyuan Zheng, Yuchen Niu, and Yue Zhang. 2021. Exploring Generalization Ability of Pretrained Language Models on Arithmetic and Logical Reasoning. In NLPCC. 758–769.
  • Wang et al. (2024b) Ke Wang, Junting Pan, Weikang Shi, Zimu Lu, Mingjie Zhan, and Hongsheng Li. 2024b. Measuring multimodal mathematical reasoning with math-vision dataset. arXiv (2024).
  • Wang et al. (2024a) Peng Wang, Shuai Bai, Sinan Tan, Shijie Wang, Zhihao Fan, Jinze Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, Yang Fan, Kai Dang, Mengfei Du, Xuancheng Ren, Rui Men, Dayiheng Liu, Chang Zhou, Jingren Zhou, and Junyang Lin. 2024a. Qwen2-VL: Enhancing Vision-Language Model’s Perception of the World at Any Resolution. arXiv (2024).
  • Wang et al. (2023a) Peiyi Wang, Lei Li, Liang Chen, Feifan Song, Binghuai Lin, Yunbo Cao, Tianyu Liu, and Zhifang Sui. 2023a. Making large language models better reasoners with alignment. arXiv (2023).
  • Wang et al. (2023b) Weihan Wang, Qingsong Lv, Wenmeng Yu, Wenyi Hong, Ji Qi, Yan Wang, Junhui Ji, Zhuoyi Yang, Lei Zhao, Xixuan Song, Jiazheng Xu, Bin Xu, Juanzi Li, Yuxiao Dong, Ming Ding, and Jie Tang. 2023b. CogVLM: Visual Expert for Pretrained Language Models. arXiv:2311.03079
  • Wang et al. (2023c) Xingyao Wang, Zihan Wang, Jiateng Liu, Yangyi Chen, Lifan Yuan, Hao Peng, and Heng Ji. 2023c. Mint: Evaluating llms in multi-turn interaction with tools and language feedback. arXiv (2023).
  • Wang et al. (2023d) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc V Le, Ed H Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou. 2023d. Self-Consistency Improves Chain of Thought Reasoning in Language Models. In ICLR.
  • wang and Fu (2024) Yifan wang and Yun Fu. 2024. Understanding, Abstracting and Checking: Evoking Complicated Multimodal Reasoning in LMMs.
  • Wang et al. (2017) Yan Wang, Xiaojiang Liu, and Shuming Shi. 2017. Deep neural solver for math word problems. In EMNLP. 845–854.
  • Wang et al. (2023e) Zengzhi Wang, Rui Xia, and Pengfei Liu. 2023e. Generative AI for Math: Part I–MathPile: A Billion-Token-Scale Pretraining Corpus for Math. arXiv (2023).
  • Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. NeurIPS 35 (2022), 24824–24837.
  • Welleck et al. (2021) Sean Welleck, Jiacheng Liu, Ronan Le Bras, Hannaneh Hajishirzi, Yejin Choi, and Kyunghyun Cho. 2021. NaturalProofs: Mathematical Theorem Proving in Natural Language. In NeurIPS Datasets and Benchmarks Track (Round 1).
  • Welleck et al. (2022) Sean Welleck, Jiacheng Liu, Ximing Lu, Hannaneh Hajishirzi, and Yejin Choi. 2022. Naturalprover: Grounded mathematical proof generation with language models. NeurIPS 35 (2022), 4913–4927.
  • Welleck and Saha (2023) Sean Welleck and Rahul Saha. 2023. llmstep: LLM proofstep suggestions in Lean. In NeurIPS.
  • Wenzel et al. (2008) Makarius Wenzel, Lawrence C Paulson, and Tobias Nipkow. 2008. The isabelle framework. In TPHOLs. 33–38.
  • Wu et al. (2021) Yuhuai Wu, Albert Jiang, Jimmy Ba, and Roger Baker Grosse. 2021. INT: An Inequality Benchmark for Evaluating Generalization in Theorem Proving. In ICLR.
  • Wu et al. (2022) Yuhuai Wu, Albert Qiaochu Jiang, Wenda Li, Markus N. Rabe, Charles Staats, Mateja Jamnik, and Christian Szegedy. 2022. Autoformalization with Large Language Models. In NeurIPS.
  • Wu et al. ([n. d.]) Yuhuai Wu, Markus Rabe, and Wenda Li. [n. d.]. LIME: Learning Inductive Bias for Primitives of Mathematical Reasoning. ([n. d.]).
  • Xia et al. (2024) Shijie Xia, Xuefeng Li, Yixin Liu, Tongshuang Wu, and Pengfei Liu. 2024. Evaluating Mathematical Reasoning Beyond Accuracy. arXiv (2024).
  • Xiang et al. (2024) Kun Xiang, Zhili Liu, Zihao Jiang, Yunshuang Nie, Runhui Huang, Haoxiang Fan, Hanhui Li, Weiran Huang, Yihan Zeng, Jianhua Han, et al. 2024. AtomThink: A Slow Thinking Framework for Multimodal Mathematical Reasoning. arXiv (2024).
  • Xu et al. (2023) Can Xu, Qingfeng Sun, Kai Zheng, Xiubo Geng, Pu Zhao, Jiazhan Feng, Chongyang Tao, and Daxin Jiang. 2023. Wizardlm: Empowering large language models to follow complex instructions. arXiv (2023).
  • Xu et al. (2024b) Guowei Xu, Peng Jin, Li Hao, Yibing Song, Lichao Sun, and Li Yuan. 2024b. LLaVA-o1: Let Vision Language Models Reason Step-by-Step. arXiv (2024).
  • Xu et al. (2024a) Jundong Xu, Hao Fei, Liangming Pan, Qian Liu, Mong-Li Lee, and Wynne Hsu. 2024a. Faithful Logical Reasoning via Symbolic Chain-of-Thought. arXiv (2024).
  • Yan et al. (2024) Yibo Yan, Jiamin Su, Jianxiang He, Fangteng Fu, Xu Zheng, Yuanhuiyi Lyu, Kun Wang, Shen Wang, Qingsong Wen, and Xuming Hu. 2024. A Survey of Mathematical Reasoning in the Era of Multimodal Large Language Model: Benchmark, Method & Challenges. arXiv (2024).
  • Yang et al. (2024) An Yang, Beichen Zhang, Binyuan Hui, Bofei Gao, Bowen Yu, Chengpeng Li, Dayiheng Liu, Jianhong Tu, Jingren Zhou, Junyang Lin, et al. 2024. Qwen2. 5-math technical report: Toward mathematical expert model via self-improvement. arXiv (2024).
  • Yang and Deng (2019) Kaiyu Yang and Jia Deng. 2019. Learning to prove theorems via interacting with proof assistants. In ICML. 6984–6994.
  • Yang et al. (2023) Zhen Yang, Ming Ding, Qingsong Lv, Zhihuan Jiang, Zehai He, Yuyi Guo, Jinfeng Bai, and Jie Tang. 2023. GPT Can Solve Mathematical Problems Without a Calculator. arXiv (2023).
  • Yang et al. (2022) Zhicheng Yang, Jinghui Qin, Jiaqi Chen, Liang Lin, and Xiaodan Liang. 2022. Logicsolver: Towards interpretable math word problem solving with logical prompt-enhanced learning. arXiv (2022).
  • Yao et al. (2023b) Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Thomas L Griffiths, Yuan Cao, and Karthik Narasimhan. 2023b. Tree of thoughts: Deliberate problem solving with large language models. arXiv (2023).
  • Yao et al. (2023a) Yao Yao, Zuchao Li, and Hai Zhao. 2023a. Beyond Chain-of-Thought, Effective Graph-of-Thought Reasoning in Large Language Models. arXiv (2023).
  • Yin et al. (2024) Shuo Yin, Weihao You, Zhilong Ji, Guoqiang Zhong, and Jinfeng Bai. 2024. MuMath-Code: Combining Tool-Use Large Language Models with Multi-perspective Data Augmentation for Mathematical Reasoning. arXiv (2024).
  • Ying et al. (2024) Huaiyuan Ying, Shuo Zhang, Linyang Li, Zhejian Zhou, Yunfan Shao, Zhaoye Fei, Yichuan Ma, Jiawei Hong, Kuikun Liu, Ziyi Wang, Yudong Wang, Zijian Wu, Shuaibin Li, Fengzhe Zhou, Hongwei Liu, Songyang Zhang, Wenwei Zhang, Hang Yan, Xipeng Qiu, Jiayu Wang, Kai Chen, and Dahua Lin. 2024. InternLM-Math: Open Math Large Language Models Toward Verifiable Reasoning.
  • Yoran et al. (2023) Ori Yoran, Tomer Wolfson, Ben Bogin, Uri Katz, Daniel Deutch, and Jonathan Berant. 2023. Answering questions by meta-reasoning over multiple chains of thought. arXiv (2023).
  • Yu et al. (2023a) Junchi Yu, Ran He, and Rex Ying. 2023a. Thought Propagation: An Analogical Approach to Complex Reasoning with Large Language Models. arXiv (2023).
  • Yu et al. (2023b) Longhui Yu, Weisen Jiang, Han Shi, Jincheng Yu, Zhengying Liu, Yu Zhang, James T Kwok, Zhenguo Li, Adrian Weller, and Weiyang Liu. 2023b. Metamath: Bootstrap your own mathematical questions for large language models. arXiv (2023).
  • Yuan et al. (2024) Lifan Yuan, Ganqu Cui, Hanbin Wang, Ning Ding, Xingyao Wang, Jia Deng, Boji Shan, Huimin Chen, Ruobing Xie, Yankai Lin, et al. 2024. Advancing llm reasoning generalists with preference trees. arXiv (2024).
  • Yuan et al. (2023a) Zheng Yuan, Hongyi Yuan, Chengpeng Li, Guanting Dong, Chuanqi Tan, and Chang Zhou. 2023a. Scaling relationship on learning mathematical reasoning with large language models. arXiv (2023).
  • Yuan et al. (2023b) Zheng Yuan, Hongyi Yuan, Chuanqi Tan, Wei Wang, and Songfang Huang. 2023b. How well do Large Language Models perform in Arithmetic tasks? arXiv (2023).
  • Yue et al. (2024) Xiang Yue, Yuansheng Ni, Kai Zhang, Tianyu Zheng, Ruoqi Liu, Ge Zhang, Samuel Stevens, Dongfu Jiang, Weiming Ren, Yuxuan Sun, Cong Wei, Botao Yu, Ruibin Yuan, Renliang Sun, Ming Yin, Boyuan Zheng, Zhenzhu Yang, Yibo Liu, Wenhao Huang, Huan Sun, Yu Su, and Wenhu Chen. 2024. MMMU: A Massive Multi-discipline Multimodal Understanding and Reasoning Benchmark for Expert AGI. arXiv:2311.16502
  • Yue et al. (2023) Xiang Yue, Xingwei Qu, Ge Zhang, Yao Fu, Wenhao Huang, Huan Sun, Yu Su, and Wenhu Chen. 2023. Mammoth: Building math generalist models through hybrid instruction tuning. arXiv (2023).
  • Zelikman et al. (2024) Eric Zelikman, Georges Harik, Yijia Shao, Varuna Jayasiri, Nick Haber, and Noah D Goodman. 2024. Quiet-star: Language models can teach themselves to think before speaking. arXiv (2024).
  • Zelikman et al. (2022) Eric Zelikman, Yuhuai Wu, Jesse Mu, and Noah Goodman. 2022. Star: Bootstrapping reasoning with reasoning. NeurIPS 35 (2022), 15476–15488.
  • Zeng et al. (2022) Andy Zeng, Maria Attarian, Brian Ichter, Krzysztof Choromanski, Adrian Wong, Stefan Welker, Federico Tombari, Aveek Purohit, Michael Ryoo, Vikas Sindhwani, et al. 2022. Socratic models: Composing zero-shot multimodal reasoning with language. arXiv (2022).
  • Zhang et al. (2024b) Di Zhang, Xiaoshui Huang, Dongzhan Zhou, Yuqiang Li, and Wanli Ouyang. 2024b. Accessing gpt-4 level mathematical olympiad solutions via monte carlo tree self-refine with llama-3 8b. arXiv (2024).
  • Zhang et al. (2023b) Mengxue Zhang, Zichao Wang, Zhichao Yang, Weiqi Feng, and Andrew Lan. 2023b. Interpretable math word problem solution generation via step-by-step planning. arXiv (2023).
  • Zhang et al. (2022) Ming-Liang Zhang, Fei Yin, Yi-Han Hao, and Cheng-Lin Liu. 2022. Plane Geometry Diagram Parsing. In IJCAI. 1636–1643.
  • Zhang et al. (2023d) Ming-Liang Zhang, Fei Yin, and Cheng-Lin Liu. 2023d. A multi-modal neural geometric solver with textual clauses parsed from diagram. In IJCAI. 3374–3382.
  • Zhang et al. (2025) Renrui Zhang, Dongzhi Jiang, Yichi Zhang, Haokun Lin, Ziyu Guo, Pengshuo Qiu, Aojun Zhou, Pan Lu, Kai-Wei Chang, Yu Qiao, et al. 2025. Mathverse: Does your multi-modal llm truly see the diagrams in visual math problems?. In ECCV. 169–186.
  • Zhang et al. (2023a) Shengyu Zhang, Linfeng Dong, Xiaoya Li, Sen Zhang, Xiaofei Sun, Shuhe Wang, Jiwei Li, Runyi Hu, Tianwei Zhang, Fei Wu, et al. 2023a. Instruction tuning for large language models: A survey. arXiv (2023).
  • Zhang et al. (2020a) Xikun Zhang, Deepak Ramachandran, Ian Tenney, Yanai Elazar, and Dan Roth. 2020a. Do language embeddings capture scales?. In EMNLP. 4889–4896.
  • Zhang et al. (2020b) Xikun Zhang, Deepak Ramachandran, Ian Tenney, Yanai Elazar, and Dan Roth. 2020b. Do Language Embeddings capture Scales?. In BlackboxNLP Workshop on Analyzing and Interpreting Neural Networks for NLP. 292–299.
  • Zhang et al. (2023c) Yifan Zhang, Jingqin Yang, Yang Yuan, and Andrew Chi-Chih Yao. 2023c. Cumulative reasoning with large language models. arXiv (2023).
  • Zhang et al. (2024a) Zhihan Zhang, Tao Ge, Zhenwen Liang, Wenhao Yu, Dian Yu, Mengzhao Jia, Dong Yu, and Meng Jiang. 2024a. Learn Beyond The Answer: Training Language Models with Reflection for Mathematical Reasoning. arXiv (2024).
  • Zhang et al. (2023e) Zhuosheng Zhang, Aston Zhang, Mu Li, and Alex Smola. 2023e. Automatic Chain of Thought Prompting in Large Language Models. In ICLR.
  • Zhao et al. (2024b) Jun Zhao, Jingqi Tong, Yurong Mou, Ming Zhang, Qi Zhang, and Xuan-Jing Huang. 2024b. Exploring the Compositional Deficiency of Large Language Models in Mathematical Reasoning Through Trap Problems. In EMNLP. 16361–16376.
  • Zhao et al. (2023a) Ruochen Zhao, Xingxuan Li, Shafiq Joty, Chengwei Qin, and Lidong Bing. 2023a. Verify-and-edit: A knowledge-enhanced chain-of-thought framework. arXiv (2023).
  • Zhao et al. (2023b) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. 2023b. A survey of large language models. arXiv (2023).
  • Zhao et al. (2022) Yilun Zhao, Yunxiang Li, Chenying Li, and Rui Zhang. 2022. MultiHiertt: Numerical Reasoning over Multi Hierarchical Tabular and Textual Data. In ACL. 6588–6600.
  • Zhao et al. (2024c) Yu Zhao, Huifeng Yin, Bo Zeng, Hao Wang, Tianqi Shi, Chenyang Lyu, Longyue Wang, Weihua Luo, and Kaifu Zhang. 2024c. Marco-o1: Towards open reasoning models for open-ended solutions. arXiv (2024).
  • Zhao et al. (2024a) Zilong Zhao, Yao Rong, Dongyang Guo, Emek Gözlüklü, Emir Gülboy, and Enkelejda Kasneci. 2024a. Stepwise Self-Consistent Mathematical Reasoning with Large Language Models. arXiv (2024).
  • Zheng et al. (2023) Kunhao Zheng, Jesse Michael Han, and Stanislas Polu. 2023. miniF2F: a cross-system benchmark for formal Olympiad-level mathematics. In ICLR.
  • Zhong et al. (2023) Wanjun Zhong, Ruixiang Cui, Yiduo Guo, Yaobo Liang, Shuai Lu, Yanlin Wang, Amin Saied, Weizhu Chen, and Nan Duan. 2023. AGIEval: A Human-Centric Benchmark for Evaluating Foundation Models. arXiv:2304.06364 [cs.CL]
  • Zhou et al. (2023d) Aojun Zhou, Ke Wang, Zimu Lu, Weikang Shi, Sichun Luo, Zipeng Qin, Shaoqing Lu, Anya Jia, Linqi Song, Mingjie Zhan, et al. 2023d. Solving challenging math word problems using gpt-4 code interpreter with code-based self-verification. arXiv (2023).
  • Zhou et al. (2023e) Andy Zhou, Kai Yan, Michal Shlapentokh-Rothman, Haohan Wang, and Yu-Xiong Wang. 2023e. Language agent tree search unifies reasoning acting and planning in language models. arXiv (2023).
  • Zhou et al. (2022) Hattie Zhou, Azade Nova, Hugo Larochelle, Aaron Courville, Behnam Neyshabur, and Hanie Sedghi. 2022. Teaching Algorithmic Reasoning via In-Context Learning.
  • Zhou et al. (2015) Lipu Zhou, Shuaixiang Dai, and Liwei Chen. 2015. Learn to solve algebra word problems using quadratic programming. In EMNLP. 817–822.
  • Zhou et al. (2023a) Wangchunshu Zhou, Yuchen Eleanor Jiang, Ethan Wilcox, Ryan Cotterell, and Mrinmaya Sachan. 2023a. Controlled Text Generation with Natural Language Instructions. In ICML, Vol. 202. 42602–42613.
  • Zhou et al. (2023b) Zhehua Zhou, Jiayang Song, Kunpeng Yao, Zhan Shu, and Lei Ma. 2023b. ISR-LLM: Iterative Self-Refined Large Language Model for Long-Horizon Sequential Task Planning. arXiv (2023).
  • Zhou et al. (2023c) Zihao Zhou, Qiufeng Wang, Mingyu Jin, Jie Yao, Jianan Ye, Wei Liu, Wei Wang, Xiaowei Huang, and Kaizhu Huang. 2023c. MathAttack: Attacking Large Language Models Towards Math Solving Ability. arXiv:2309.01686 [cs.CL]
  • Zhu et al. (2021) Fengbin Zhu, Wenqiang Lei, Youcheng Huang, Chao Wang, Shuo Zhang, Jiancheng Lv, Fuli Feng, and Tat-Seng Chua. 2021. TAT-QA: A Question Answering Benchmark on a Hybrid of Tabular and Textual Content in Finance. In ACL. 3277–3287.
  • Zhu et al. (2022) Xinyu Zhu, Junjie Wang, Lin Zhang, Yuxiang Zhang, Ruyi Gan, Jiaxing Zhang, and Yujiu Yang. 2022. Solving math word problems via cooperative reasoning induced language models. arXiv (2022).